### v2

# ToDo (#1): #2

Abstract: This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source. Volumes I and II are site-independent and describe the cyclotron driver providing a 10 mA proton beam, and the medium energy beam transport line and target, respectively. Volume III describes the installation at the Yemilab underground laboratory in South Korea. The IsoDAR driver and target will produce a mole of electron-antineutrinos over the course of five years. Paired with a kton-scale liquid scintillator detector, it will enable an impressive particle physics program including searches for new symmetries, new interactions and new particles. Here in Volume I, we describe the driver, which includes the ion source, low energy beam transport, and cyclotron. The latter features radiofrequency quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of vortex motion.
[…]
for the IsoDAR collaboration
${}^{1}$on Beam Applications, 3 chemin du cyclotron Louvain-la-Neuve, BELGIUM
${}^{2}$assachusetts Institute of Technology, 77 Massachusetts Ave. Cambridge, MA, USA
Corresponding Authors: Joshua Spitz (spitzj@umich.edu) and Daniel Winklehner (winklehn@mit.edu)
V0.9
Chapter 1 Introduction
Figure 1.1: Schematic of the IsoDAR experiment installed at the Yemilab facility in South Korea. From right to left: The compact cyclotron with RFQ injector, the medium energy beam transport, the target, and the LSC neutrino detector. From [1].
IsoDAR will be the first high-intensity, proton-driven source to be located alongside an underground multi-kiloton-scale scintillator detector. The first version of IsoDAR will be constructed at the Yemilab facility [2] within the Handuk mine, located in South Korea where the “Liquid Scintillator Counter” (LSC) is planned to be installed [3]. The novel source, consisting of a cyclotron, beam transport line, and a target/sleeve is shown in Fig. 1.1. In five calendar years (four live years), the source will produce 1.67 million $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ (inverse beta decay or ”IBD”) events and 7000 $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$ (elastic scattering or ES) events in the LSC that can be used for searches for new physics. More than a mole of neutrons and monoenergetic photons are also produced in the target, and these may act as progenitors for new particles that enter the LSC and leave signatures from interactions and decays. IsoDAR can address a broad range of new physics manifested as new symmetries, new interactions and new particles that is inaccessible to traditional experiments. For further discussion see Refs. [4, 5].
This document is Volume 1 in a three-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering solutions necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D. The contents of the PDR will cover the Technical Facility: the cyclotron, the beam transport and the target/sleeve. This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [6]. This step of producing a PDR substantially reduces risk related to IsoDAR approval: the remaining detailed engineering that will follow will build on these results and be straightforward to estimate in scope and cost.
Volume 1 describes the IsoDAR proton driver. This consists of a “front end” that delivers ions to a cyclotron for acceleration and extraction and the cyclotron itself. This document presupposes a basic understanding of cyclotrons; for an introduction to cyclotrons and a cost/benefit analysis of this choice of machine for IsoDAR, see the discussion in the CDR.
The 60 MeV/amu IsoDAR Proton Driver is exceptional in that it is designed and engineered to produce an order of magnitude higher intensity proton beam at extraction than commercial machines of similar energy. Achieving this high power necessitated three key breakthroughs. First, the cyclotron accelerates $\mathrm{H}_{2}^{+}$, mitigating “space charge effects” (electromagnetic interactions between beam particles). Second, the beam is axially injected through an RFQ, acting as a high-efficiency buncher. Third, the design harnesses vortex motion, a complex effect previously observed at PSI Injector II [7, 8], inducing stable spiraling motion in the high-current accelerated beam, significantly reducing beam growth. For more concerning the accelerator physics that shapes the IsoDAR design, see Ref. [9, 10].
These three physics requirements for reaching high current drive many of the engineering choices presented in this PDR. The other design driver is the need to construct the cyclotron underground. The cavern for the IsoDAR Proton Driver has already been constructed at Yemilab, and so the dimensions and constraints are understood, leading to fully realistic engineering plans.
This PDR is structured as follows:
Chapter 1:
(This chapter) proceeds with an overview of the physics searches that will be possible with IsoDAR, followed by an overview of the requirements for the proton driver.
Chapter 2:
Describes engineering related to the IsoDAR front end.
Chapter 3:
Describes engineering related to the general IsoDAR cyclotron design.
Chapter 4:
Describes specifics of the design related to installation underground.
1.1 The IsoDAR Physics Case
Figure 1.2: Energy dependence of fluxes. Left: $\bar{\nu}_{e}$, unit normalized; Middle: photons per proton on target (POT) from excited nuclei. Right: kinetic energy of neutrons at production in target and sleeve, normalized per proton.
IsoDAR offers a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the proposed LSC at Yemilab [3]. The exceptional reach is owed to the high flux rates of $\bar{\nu}_{e}$, photons, and neutrons from the $\approx 2\cdot 10^{24}$ protons on target over 5 years (cf. Fig. 1.2), and placing the source very close to the detector. Here we briefly discuss the highlights of the IsoDAR BSM physics program enabled by a 2.26 kton detector like the LSC in three categories: “New Particles” - experimentally and theoretically motivated; “New Symmetries” - CP-violation and the QCD Axion; And “New Interactions” - Precision Electroweak Physics. See Refs. [4, 11, 5] for further details.
1.1.1 Fluxes in IsoDAR
We consider three fluxes of particles produced in the IsoDAR target for BSM physics searches: the $\bar{\nu}_{e}$ flux, the neutron flux, and the $\gamma$ flux from the excitation of nuclei as neutrons are moderated. These fluxes are shown in Fig. 1.2.
The $\bar{\nu}_{e}$ flux stems from a single beta-decay of ${}^{8}$i, which is produced by neutron capture on ${}^{7}$i and is well-understood. It peaks at 6 MeV, which puts it above backgrounds from the environment. These are clear advantages over reactor-neutrino fluxes. Interactions in the detector at these neutrino energies are exclusively inverse-beta-decay (IBD) and elastic scattering (ES). Cross sections for these reactions are known with uncertainties at $<1$% at these energies.
Neutrons in the IsoDAR target are primarily produced by spallation interactions of the beam protons with the beryllium target hemispheres. Additional secondary neutrons stem from interactions of the primary neutrons with beryllium in the target and sleeve. We model the processes in the target with Geant4 using an accurate representation of the latest target and sleeve design (see Refs. [12]). We show the resulting neutron energy spectrum in Fig 1.2 (right). Most neutrons capture on the ${}^{7}$i in the sleeve. To ensure $<5$ neutrons per year enter the LSC veto region, the target is further shielded using boron-loaded concrete and steel [13].
The third flux we consider is that of photons as seen in Fig.1.2 (middle). Here, one can observe a line structure that stems from the transitions of excited nuclei and and underlying smooth continuum due to bremsstrahlung. Again, the spectrum was generated using Geant4 using the full IsoDAR Target geometry [5].
1.1.2 New Particles: Experimentally and Theoretically Motivated
Figure 1.3: An example of a potential 3+2 signal in IsoDAR, orange–true signature, black–IsoDAR reconstruction with uncertainty shown.
[…]
Figure 1.4: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with ${}^{8}$i-produced $\bar{\nu}_{e}$ background (gray). Right: $X$ mass and $X$-to-photon production ratio 90% CL sensitivity [4].
Another example of IsoDAR’s capability to look for new particles is the “$X$,” which is a low mass mediator that would couple to the photons produced in the target [4]. The $X$ would be a light new particle that decays to $\bar{\nu}\nu$ and would leave a distinct signature above background in the IBD spectrum. Fig.1.4 (left) displays resulting signatures for various masses (see color bar) assuming 0.5% $X$-to-photon production. The IBD background from $\bar{\nu}_{e}$ stemming from ${}^{8}$i decays (gray) constitutes the primary background and has a distinct shape. The sensitivity is depicted in Fig.1.4 (right), where the blue region is excluded in the case of the SM. No existing limits are shown because this search is first-of-its-kind. A 17 MeV $X$ particle is consistent with the Atomki anomaly.
1.1.3 New Symmetries: Strong CP Violation and the QCD Axion
Figure 1.5: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [5]. Gold dots–axion allowed; Teal–excluded regions; Gray–parameters with astrophysical implications.
The neutron beam/bottle experiment lifetime discrepancy [17, 18, 19, 20] motivates neutron–dark sector couplings. Through $n\rightarrow n^{\prime}\rightarrow n$ transitions, we may see neutrons disappearing from the target and reappear in the LSC. The underground environment provides a low background and the LSC’s size further adds to the sensitivity [11]. Alternatively, if the $n^{\prime}$ are Majorana, then $n\rightarrow n^{\prime}\rightarrow\bar{n}$ may occur, depositing $\sim$2 GeV of energy in the LSC upon annihilation. This is a $B$ violating effect with implications for the early universe [21].
The photons and electrons in the target may also couple to axions and IsoDAR can probe uncharted territory with astrophysical and cosmological implications [5]. In Fig. 1.5 we show QCD axions coupling to electrons.
1.1.4 New Interactions: Tests via Precision Electroweak Physics
IsoDAR will record about 7000 elastic scatter events over its 4-year accumulated up-time. This constitutes the world’s largest sample at low momentum transfer. LEP and LHC data predict rate and energy dependency very precisely within the standard model [22]. With the normalization determined by the $\approx 1.7$M IBD events in the LSC, we will be able to measure $\sin^{2}\theta_{W}$ with a precision of 1.9% (see Fig.1.6 left).
As a neutrino-nucleon experiment, the observed deviation of NuTeV may stem from unaccounted-for nuclear physics effects; however, numerous possibilities have been investigated, but none fully explain the deviation. Alternatively, the deviation might signal new physics in the neutrino sector. If so, this BSM signature will also manifest in the IsoDAR ES measurement. Fig.1.6, right, presents two expectations for IsoDAR, where the red assumes background reduction from Cherenkov ring identification in LSC, while the green indicates the expectation without this additional information. Statistics limit the outcome, and further running can proceed with IsoDAR and DUNE as it comes online if a NuTeV-like deviation is observed.
Figure 1.6: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation. Right: expectation for IsoDAR in NSI parameter space. See Ref. [4].
The accurately predicted energy dependence of the ES sample enables a precise phenomenological search for non-standard neutrino interactions (NSIs). This powerful search for lepton-lepton couplings complements the lepton-quark coupling searches by the active, global coherent neutrino scattering program. The SM right- and left-handed couplings, $g_{R}$ and $g_{L}$, are altered by NSI corrections for electron flavor antineutrino interactions with electrons (hence ”ee”) to $\tilde{g}_{R}=g_{R}+\varepsilon_{ee}^{eR}$ and $\tilde{g}_{L}=g_{L}+\varepsilon_{ee}^{eL}$. The SM expectation for IsoDAR@Yemilab in the parameter space of these deviations is seen in Fig. 1.6 (right), where red assumes background reduction using Cherenkov light, and green assumes only scintillation light reconstruction. This can be compared to the world result [23] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments and, hence, is rotated with respect to IsoDAR, which has a pure $\bar{\nu}_{e}$ flux.
1.2 Requirements for the Proton Driver
As shown in Fig. 1.1, the Proton Driver consists of a 60 MeV/amu cyclotron accelerating and extracting 5 electrical milliamperes of $\mathrm{H}_{2}^{+}$, a transport line that strips the $\mathrm{H}_{2}^{+}$ ions close to the extraction point, analyzes the stripped beam to monitor the health of the stripping foil, and transports the 10 milliamperes of protons to the target. The target consists of three nested hemispherical beryllium shells, the outermost being 20 cm in diameter, cooled with heavy water flowing around the shell structure. The proton beam is shaped, with quadrupoles and wobbler magnets to produce a distribution that allows for optimal heat transfer to the cooling water. Neutrons produced in the target system are moderated and flood into a roughly-spherical sleeve, about 1 meter in diameter, containing a mixture of beryllium chips and highly-enriched ${}^{7}$i. The beta-decay of the ${}^{8}$i resulting from the capture of neutrons produces the neutrinos of interest in the experiment.
The principal requirements for the accelerator system are to reliably produce a 10 mA beam of protons in the most efficient manner possible with the lowest possible loss of particles in the acceleration process. Experience from the cyclotron system at PSI [24] are that keeping the total power of beam loss inside the cyclotron vault to less than 200 watts still allows access for component maintenance. Considering that the total beam power is 600 kilowatts, one sees that this is a very stringent requirement.
As mentioned above, Volume 1 of this PDR addresses the physics and engineering aspects of the 60 MeV/amu cyclotron. The final chapter addresses the complex issue of transport and assembly of the large and heavy cyclotron components in the underground environment at Yemilab.
Subsequent volumes will address the details of the beam transport system and the target, along with shielding requirements for the environment of the target including the prevention of fast neutrons from reaching the fiducial volume of the liquid scintillator detector and the underground installation procedure, power and cooling, HVAC, and other site-specific topics.
Chapter 2 Design of the Front End
2.1 The Ion Source
The $\mathrm{H}_{2}^{+}$ beam to be injected into the cyclotron and accelerated to 60 MeV/amu will be produced in a filament-driven multicusp ion source which we will refer to as MIST-1. The technical details of this ion source are reported in Refs. [25, 26, 27, 28, 29] and are summarized here.
2.1.1 Design
[…]
2.1.1.1 Gas Input
Hydrogen gas is introduced via an inlet in the backplate of the source as indicated in Fig. 2.2. The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 SCCM full range. This system interfaces with the PC control system which is documented in detail in Ref. [29].
2.1.1.2 Ionization by the Filament
After the hydrogen gas is introduced into the chamber, it is ionized to form a plasma using electrons released from a tungsten filament by thermionic emission. The filament is mounted on DI water-cooled power feedthroughs in the back plate (cf. Fig. 2.2). The filament is heated by an (8.4V, 300 A max) power supply.
In addition to this power supply which is used to heat the filament, a second power supply is connected to the filament and to the source chamber to maintain a potential difference of 80-140V between them in order to facilitate electron discharge. The source body, back plate, and front plate are all mutually electrically insulated from each other for added flexibility in the electric field shape. A schematic of the wiring of the ion source is illustrated in Fig. 2.2
[…]
Figure 2.1: A 3D CAD rendering of the MIST-1 ion source.
Figure 2.1: A 3D CAD rendering of the MIST-1 ion source.
Figure 2.2: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Thick blue lines are data cables. Black lines are power cables.
2.1.1.3 Confinement & Extraction of the Plasma
[…]
The electrons are then shaped and accelerated out via an extraction system (cf. Fig. 2.2). The extraction system comprises five electrodes which shape and accelerate the ion beam. These include a plasma electrode (+15 kV), a puller electrode (-3 to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV). The RFQ entrance flange is held at ground potential. We simulated beam formation and shaping in the extraction system with the IBSimu package [30], which performs 3D simulations including space charge. Due to the computational resources needed to perform full 3D simulations in IBSimu, we adopted a hybrid IBSimu (for the plasma modeling)/WARP (for the beam transport) [31] model for the final design. We show example trajectories in Fig. 2.3.
Figure 2.3: Cross section of the extraction system showing the ion beam moving through the extraction system prior to entering the RFQ. Simulation of the extraction system performed by IBSimu and WARP software.
Table 2.1: A summary of the parameters for the MIST-1 ion source.
Parameter Nominal Value
Plasma chamber length 6.5 cm
Plasma chamber diameter 15 cm
[…]
Permanent magnet strength 1.05 T on surface
Front plate magnets 12 bars (star shape)
Radial magnets 12 bars
Back plate magnets Four bars in three rows
Front plate cooling Embedded steel tube
Back plate cooling Embedded copper pipe
Chamber cooling Water jacket
Water flow (total) $\approx$2 L/min
Filament feedthrough cooling Water cooled
Filament material Water mixed with Cu and Ni
Filament diameter $\approx$0.8 mm
Discharge voltage Max. 150 V
Discharge current Max. 24 A
Filament heating voltage Max. 8 V
Filament heating current Max. 100 A
2.1.2 Performance
2.1.2.1 Goals & Measurement Methods
In this section we summarize the latest measured results as well as discussing the results of several optimizations that have been thus far performed as well as several that remain to be performed. The ion source target parameters that will be required for IsoDAR to achieve the planned decisive results for a sterile neutrino search include the following benchmarks:
  • High current: $\approx$10 mA of H+${}_{2}^{+}$
  • High purity: H+${}_{2}^{+}$ fraction $>$ 80%
[…]
  • Ability to run continuously in DC mode
Characterization of the ion source is done using a short diagnostic beam line. This test beam line comprises two Faraday cups, a dipole magnet with separation slits, and a pair of Allison emittance scanners. The total beam current is measured using the first Faraday cup located posterior to the extraction system. The dipole magnet with separation slips separates the ions by mass and passes them to the second Faraday cup which, in conjunction with the Allison scanners, measures each species horizontal and vertical emittances.
2.1.2.2 Optimization of Ion Source Parameters
The ion source is designed such that several parameters can be varied in order to optimize the magnitude, purity, and emittance tightness of the current from the ion source.
[…]
  • Filament discharge voltage
  • Filament discharge current
  • Filament size, shape, & position (Optimization not yet studied)
  • Permenant magnet type (Optimization not yet studied)
[…]
Filament discharge voltage optimization. Two series of filament discharge voltage optimizations have been performed differing in how discharge current is managed. In the first, discharge current is permitted to freely find a steady state using the same method as used in the flow optimization described above. In the second, the discharge current was fixed at 4 A.
Filament discharge current optimization. One series of current optimizations was explored in which the flow rate and discharge were kept constant while discharge current was varied from 2 to 8 A. No strong dependence on current was observed.
2.1.2.3 Latest Optimized Results
To date, only an incomplete optimization of the ion source has been performed. As such, parameters which simultaneously meet all the target goals have not yet been fully identified. The ion source’s peak performance to date varies, therefore, depending on which of the performance metrics is optimized. Optimization has been performed for maximum total current density, maximum H+${}_{2}^{+}$ purity, and a hybrid option optimizing for the maximum H+${}_{2}^{+}$ current density with the constraint that H+${}_{2}^{+}$ be dominant.
[…]
With current optimizations and a 4mm aperture, the ion source is capable of delivering approximately 1.1 mA of H+${}_{2}^{+}$ with 76% purity. With an 8 mm aperture this corresponds 4.4 mA. This is a world leading result and is only about a factor of 2 short of target.
The main bottleneck for increased beam current was the extraction system, which was designed for initial testing and commissioning. We have recently replaced it with the version presented in this chapter. Our high fidelity simulations suggest that the new extraction system will be able to handle the nominal beam current. Measurements are planned for summer 2024.
2.2 The Low Energy Beam Transport
The low energy beam transport line (LEBT) serves the purpose of connecting the ion source to the RFQ. The LEBT steers and shapes the beam, guiding it into the RFQ. The LEBT also provides a location at which to take beam diagnostics. The requirements for the LEBT are:
  1. 1.
    Transport beam with minimal losses,
  2. 2.
    Provide space for diagnostics,
  3. 3.
    Match the beam to the RFQ,
  4. 4.
    Provide safety mechanisms for personnel and machine protection.
Details of the LEBT design and performance were given in Refs. [32, 33, 34] here we summarize and update these.
2.2.1 Design Overview
At this stage, ions emerge from the ion source. The beam now passes through a gap surrounded by a series of six electrodes arranged axially within a six-way cross (see Fig. 2.2). These six electrodes can have their voltage optimized in order to best match the parameters for acceptance into the RFQ. A conical flange is connected to the end of the six-way cross. The conical shape minimizes beam loss on exit.
Following the six-way cross lies a diagnostic segment. A gate valve separates the six-way cross from this segment. The diagnostic section comprises a second cross, which can be used for additional vacuum attachments, diagnostic readouts, and a chopping electrode. The diagnostic section also includes a solenoid and two magnetic steerers which can be used for additional beam tuning. A flange in place toward the end allows space for an ACCT which characterizes the beam immediately prior to entering the RFQ. The diagnostic segment is illustrated in Fig. 2.3.
2.2.2 Beam Steering & Chopping
The diagnostic segment also serves to steer the beam into the RFQ or terminate the beam as needed. Even small, sub-millimeter misalignment can prevent efficient injection of the beam into the RFQ. In order to ensure that the beam enters the RFQ at both the correct position and angle, two magnetic steerers and a solenoid are placed around the beamline. These are controlled independently by an external power source to produce tunable dipole fields which optimize beam entry into the RFQ. The optimal fields were calculated using COMSOL and Warp. These are discussed further in the following section.
In order to control the amount of power delivered from the ion source, a beam chopper is built into the cross of the diagnostic segment to allow for a duty factor in an otherwise DC beam. The beam chopper is built into the cross in the diagnostic section. The chopper comprises a copper ring surrounded by a steel beampipe which in turn is surrounded by a copper water cooling jacket. The copper ring can be put at a high potential relative to the cross, which is grounded. If the potential is sufficiently high the beam can be diverted such that it terminates. See Fig. 2.5. This method permits control over the amount of power delivered to the RFQ without disrupting the plasma itself.
A 12 mA beam of 15 keV particles fully terminating on the beam stop corresponds to $\sim$180 W power delivered. We need to ensure that the dump is capable of dissipating this amount of power. The system was modeled assuming a water cooling line at room temperature. To be conservative, the simulation was run using four times the actual anticipated beam power, i.e. 720 W. The beam dump never rose above 360 K in our simulation, significantly below the melting point of copper. The COMSOL simulation is shown in Fig. 2.5
Figure 2.4: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid. If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring.
Figure 2.4: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid. If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring.
Figure 2.5: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior.
2.2.3 Beam Dynamics
Beam simulations are performed in order to characterize the performance of the LEBT with the ultimate goal of ensuring that beam’s Twiss parameters matches the desired input parameters for the RFQ.
Beam simulations for the LEBT have been performed using IBSimu and Warp. IBSimu provides a more accurate plasma model, but is computationally intensive and so was used only for the region by the ion source where the plasma is extracted and the density is thus high, requiring this accuracy. IBSimu is used for the first several mm after the extraction aperture where density remains high. After this, and for the remainder of the downstream region of the LEBT, the distributions from IBSimu are handed off to Warp which provides accurate results for the lower density regions of the beam with less computational burden.
For the IBSimu simulations a 12 mA beam composed of 80% H+${}_{2}^{+}$, 10% ${}^{+}$, and 10% H+${}_{3}^{+}$ was used. This provides ample current for IsoDAR assuming 50% global transmission efficiency while using current from the ion source that has already been empirically demonstrated. As illustrated in Fig. 2.6 and in Table 2.2 the LEBT is able to match the required input parameters for the RFQ. Moreover, as illustrated in Fig. 2.7 this match is achieved with emittance that surpasses that used for baseline RFQ simulations.
Figure 2.6: A temperature plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT.
Figure 2.7: Phase space from the LEBT simulation (left) vs phase space for ideal beam input to the RFQ (right).
Table 2.2: A comparison of LEBT Twiss parameters and baseline RFQ parameters.
Parameter LEBT Output Baseline
[…]
$\alpha$ 1.3 1.0
$\beta$ 0.13 mm/mrad 0.17 mm/mrad
2.3 The Radiofrequency Quadrupole
Figure 2.8: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle.
The $\mathrm{H}_{2}^{+}$ beam coming from the LEBT is injected into a Radio Frequency Quadrupole (RFQ) that is partially embedded in the cyclotron yoke. A schematic of the RFQ and cyclotron can be seen in Fig. 3.1 where only about a third of the RFQ is visible above the top surface of the cyclotron. We show a cut view of the RFQ in Fig. 2.8. Of note are the small diameter of the RFQ (27.6 cm) while operating at a low resonant frequency of 32.8 MHz. This is achieved by utilizing the so-called “split-coaxial” mode where the vanes are attached longitudinally at the entrance and exit flanges instead of radially at the mantle. The main operating parameters of the RFQ are listed in Tab. 2.3. The RFQ accelerates only moderately from 7.5 keV/amu to 35 keV/amu and acts mainly as a very efficient buncher. This leads to a conservatively estimated power consumption of $<6$ kW.
Table 2.3: A summary of the design parameters for the RFQ.
Parameter Nominal Value
Diameter 27.6 cm
Frequency 32.8 MHz
Tuning range ±$\pm 270$ kHz
Q simulated 2800
[…]
Input Energy 7.5 keV/amu
Output Energy 35 keV/amu
Duty factor 100% (cw)
Power $<6$ kW
Cooling DI Water channels in tank and vanes
2.3.1 Technical Design
The technical design was presented in Refs. [35, 36, 37] and the optimization of the RF loop coupler in Ref. [38]. The thermal properties and cooling considerations were presented in Ref. [39]. An optimization scheme using machine learning was presented in Ref. [40].
Figure 2.9: Modes of the mechanical deformation of the RFQ vanes. The frequencies are 18 Hz and 25 Hz.
In terms of mechanical stability, our analyses show two quarter wave modes which are associated with the choice of a split-coaxial resonator. The addition of bridges between opposite vanes (see Fig. 2.9) significantly reduces detuning due to the Lorentz force with mechanical frequencies around 18 Hz and 25 Hz.
2.3.2 Beam Dynamics
Beam dynamics studies were performed of the full 3D particle distribution resulting from the LEBT simulations going through the RFQ, including the effects of the entrance gap and asymmetry (due to the split coaxial design) and the exit gap and asymmetry. Fine tuning of the phase of the exit gap (i.e. tuning the length of a drift preceding the exit gap), leads to final output beam parameters listed in Tab. 2.4.
Table 2.4: RFQ output beam parameters.
Parameter Nominal Value
$\varepsilon_{x}$ (1-RMS, norm.) 0.54 mm-mrad
$\varepsilon_{y}$ (1-RMS, norm.) 0.52 mm-mrad
$\varepsilon_{z}$ (1-RMS, norm.) 0.114 MeV-deg
Phase width (1-RMS) ${}^{\circ}$
Chapter 3 Design of the Cyclotron
Figure 3.1: Isometric view of the cyclotron
In this section we describe the details of the cyclotron design at the level of a preliminary design report.
3.1 Magnet Yoke
The magnet design presented here will necessarily be modified for a final design, as the specific details of the extraction orbit and all required extraction elements have not been finalized.
The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron, which is slightly larger, at 7 meters outer diameter.
Figure 3.2: C400 on the milling machine, C230 in front of it.
3.1.1 Design and fabrication of the cyclotron magnet
The cyclotron is divided in the same fashion as the C400, into smaller parts facilitating transport to the site. The magnet is split radially and will be assembled on site. The vacuum chamber is split in 3 (self-centering) parts along its height, the vacuum tightness is ensured by 2 axial O-rings at each interface except for the interface between the upper and lower chambers were 2 radial O-rings are used. An exploded of the upper and lower parts is shown in Figs. 3.4 and 3.4. The materials used for the magnet parts is cyclotron steel (the exact composition shall be determined after a detailed physics study), the vacuum chambers will be made of stainless steel.
Figure 3.3: Cyclotron upper steel exploded.
Figure 3.3: Cyclotron upper steel exploded.
Figure 3.4: Cyclotron lower steel exploded.
An overview of a cut in the cyclotron is shown in Fig. 3.5. The external and vacuum disks are connected to each other as well as to the flux return using specific bolts (such as the Superbolt from Nord-Lock). The poles are split in several parts and an easily removable part should be added in the next iteration to ease the mapping and possibly improve the magnetic performance. A detailed view of the cyclotron components is shown in Fig. 3.6.
Figure 3.5: Cyclotron cut view
Figure 3.6: Cyclotron detailed cut view
Table 3.1 gives some more information regarding the yoke parts.
Name Mass (Metric Ton) Dimensions (mm)
Upper and lower external disk 30 Ø4700*250
Upper and lower internal disk 36 Ø4620*463
Upper and lower flux return (1 and 2) 36 3100*1950*1015
Upper and lower median plane flux return (1 and 2) 29 3100*1950*705
Upper and lower pole 7 2013*1290*600
Table 3.1: Yoke part properties
3.2 Interface with the RFQ
Figure 3.7: RFQ overview
The RFQ will be outside of the cyclotron’s vacuum, therefore a vacuum sleeve will be used to ensure the integrity of the cyclotron’s vacuum. This sleeve will be fixed on top of the upper external disk and the vacuum sealing will be done radially on the upper internal disk (as seen in figure number 3.9 ). The O-rings could have a cross section diameter of 8mm and an inner diameter of 305mm.
In order to allow some adjustement, the RFQ sleeve allows a rotation of 5mrad around the lower O-ring contact patch (discussed with the manufacturer of the RFQ).
Figure 3.8: RFQ zoom
The RFQ sleeve is as thin as possible to avoid any loss in magnetic field in the central region (hence the O-rings in the Yoke parts).
Figure 3.9: RFQ cut section 1
Fig. 3.10 shows that the vacuum sealing will be done at the bottom of the RFQ using a thick O-ring (diameter 10 mm) which will be compressed by the RFQ’s weight ($\approx 3000$ N). The dimensions of the vacuum sleeve are listed in Tab. 3.2.
The adjustment feet shown in Fig. 3.8 are fixed to the cyclotron yoke and allow the following alignment modifications through four sets of four set screws one and vertical adjustment nut:
  • Rotation: ±5plus-or-minus5$\pm$ 5± ${}^{\circ}$
  • Tilt: ±$\pm 5$ mrad
  • Lateral movement: ±$\pm 1$ mm
In addition to this fine-tuned alignment, a coarse alignment step is envisioned during the initial assembly and commissioning procedure.
Figure 3.10: RFQ cut section 2
Table 3.2: Vacuum Sleeve Dimensions.
Name Value
Height 1030.5mm
Outer diameter 304.55mm
Inner diameter 296.55mm
Weight 18kg
3.3 Central Region
Figure 3.11: CAD rendering of the spiral inflector and particle trajectories for early tests at Best Cyclotron Systems Inc. [41]. The positive and negative electrostatic deflectors are shown in blue and green respectively. The cyclotron magnetic field in this image is directed vertically upwards. Particles enter the spiral inflector via the rectangular grounded collimator (solid gray) and are guided into the cyclotron mid-plane by means of the cyclotron magnetic field and the electrostatic potential between the electrodes. The copper housing (transparent gray) isolates the spiral inflector from the RF fields driving the cyclotron.
The injection system design is a crucial component to achieve the highest beam current. Our design balances injecting the $\mathrm{H}_{2}^{+}$ molecule at high energy against avoiding the use of complex high-voltage platforms. Higher energy injection mitigates space charge and to reduces geometrical (non-normalized) emittance. An energy of 35 keV/amu was selected which determined the output energy of the RFQ. As in commercial cyclotrons, an axial injection system based on a Spiral Inflector (SI) will be used to bend the beam from the axial direction to the median plane. This concept is visualized in Figure 3.11 using trajectories calculated with OPAL [42]. The exit point of the spiral inflector is also the starting point of acceleration in the Central Region (CR) of the cyclotron. The central region and the spiral inflector have to be designed carefully and together because they have strong interplay. This procedure has already been performed and tested for the injection tests during the summers of 2013 and 2014 at the Best Cyclotron Systems Inc. (BCSI) laboratory in Vancouver. In the following subsection, we will describe the design process of the central region for the BCS test stand cyclotron followed by a preliminary design for the full IsoDAR machine.
3.3.0.1 Spiral Inflector Design
The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Figure 3.11). The electrostatic potential between these electrodes is able to bend the beam ${}^{\circ}$ from the axial line to the median plane of the cyclotron. The helical trajectory of the beam is determined both by the shape of the electrodes (electric field) and the magnetic field of the cyclotron as the beam is bent into the median plane. The applied voltage depends on the velocity of the beam as well as the rigidity of the ions.
Figure 3.12: Photograph of the spiral inflector used during the tests at Best Cyclotron Systems.
The spiral inflector tested at BCSI (shown in the photograph in Figure 3.12) was designed to mimic the conditions required for the IsoDAR inflector, its primary defining characteristic being the 15 mm gap between the electrodes. Compared to other similar designs, this is rather large because it has to take into account the larger beam size due to space-charge effects and the higher rigidity of $\mathrm{H}_{2}^{+}$. Due to the large distance between the electrodes, the SI occupies a volume in which the magnetic field components’ variation is not negligible, as it is assumed in the analytical treatment of the spiral inflector. This effect has to be carefully taken into account in order to shape correctly the spiral inflector electrodes and to avoid the introduction of a high energy-spread during the beam transport through the device. The spiral inflector has a copper housing that surrounds the electrodes to minimize the interaction between the electric fields generated by the dees and the electrostatic fields between the electrodes. To minimize beam striking the uncooled electrodes, a water-cooled, grounded, rectangular collimator shields the spiral inflector entrance.
The preliminary design of the spiral inflector called for a nominal voltage of ±$\pm 11$ kV and a tilt angle of ${}^{\circ}$ for optimal beam injection at 60 keV. The total height of the device was approximately 80 mm. These parameters were found by using a MATLAB code based on an analytic theory for spiral inflectors. The electrode shape was calculated using VectorFields OPERA. This design was later modified to account for some effects due to fringe fields on the electrodes. One of these modifications was to reduce the overall length of each electrode by 5 mm.
First turn acceleration is achieved by the dee tips extending into recesses in the spiral inflector housing. The shapes of the dee tips and the spiral inflector housing have been designed to guide the particles from the spiral inflector exit to the acceleration region while providing the necessary energy gain and beam focusing.
[…]
The main differences between the BCSI test stand design and the IsoDAR design are:
  • The IsoDAR cyclotron will be equipped with 4 accelerating RF cavities, while the test cyclotron used only 2 cavities.
  • The harmonic mode (and the frequency) will be decreased from 6th harmonic (49.2 MHz) to 4th harmonic (32.8 MHz).
  • The injection energy will be increased from 30 keV/amu to 35 keV/amu.
  • The spiral inflector gap will be increased from 15 mm to 16.5 mm to accommodate a larger beam.
3.4 Vacuum Chamber
The vacuum enclosure is composed of 5 parts : the upper and lower internal disk and 3 vacuum chambers. The vacuum sealing is done using 2 metallic O-rings per interface (from Omniseal [https://www.omniseal-solutions.com/] for example) except for the radial interface where Polymer O-rings could be used. The interface for the accessories will be done using a sleeve design and radial O-ring, described in the figure hereunder.
Figure 3.13: Cyclotron vacuum chamber details
Name Mass (Metric Ton) Dimensions (mm)
Upper and lower internal disk 36 Ø4620*463
Upper vacuum chamber 2.3 Ø4452*483
Lower vacuum chamber - median plane 2.3 Ø4452*483
Lower vacuum chamber - internal disk 1.8 Ø4386*380
Table 3.3: Vacuum part properties
3.5 Coil configuration
[…]
Figure 3.14: Left: schematic view of a cross-section of the main coil, with six pancakes per half coil. Right: possible configuration of the winding, using a square conductor with a round cooling channel in its center.
[…]
Working in pancakes gives the possibility of manufacturing the pancakes outside the cave and stacking them inside. The increase in height from 200 to 240 mm is the equivalent of adding an extra pancake.
3.6 Cyclotron RF System
Table 3.4 provides a detailed description of the RF system. The system for IsoDAR cyclotron is divided into subsystems as follows:
  • The RF cavities allowing a fast acceleration to the final particle energy and keeping a large inter-turn separation. The cavities are triangular shaped and are located in the magnet valleys. There are 4 accelerating electrodes (Dees).
  • The RF tuning system that will compensate for thermal drift of cavities and maintain a stable resonating frequency.
  • The RF amplifiers that provide the necessary power to create the required RF electric field inside the cavity and the power to accelerate the beam.
  • The RF couplers that inject the RF power from the amplifiers into the cavities
  • The Low Level RF system (LLRF) that controls and regulates the RF amplitude on the Dee and drives the tuning system.
Table 3.4: Details of the RF design.
RF System Component Design Value
Resonance Frequency 32.8 MHz
Dee voltage in the central region $70$ kV
Dee voltage at extraction radius $230$ kV
Dee radial extension 2 m
Dee angular extension ${}^{\circ}$
Cavity height 1800 mm
Number of Dee stems 4 per Dee (2 up, 2 down)
Number of Dee’s 4
Acceleration harmonic 4th
Power dissipated per cavity 116 kW
Cavity Q-factor 9620
We consider these in the Sections below.
3.7 RF Cavities (Dees, Stems and Liners)
Figure 3.15: Early RF Dee computation – color scheme represent current density
The RF cavities of the cyclotron have been calculated using CST microwave studio [43] in order to define the shape of the Dee and the Dee stems. They have then been optimized also with CST from this early design. The optimization work is intended to reach the right resonance frequency, the Dee voltage Law, together with the minimal power dissipation and hence the highest quality factor (Q-factor). The field maps allow also the calculation of the couplers and voltage pick-up but this last calculation is yet to be performed.
The Dees and cavity walls are all made of OFHC copper that combines good electrical conductivity and good solder ability. The Dees are fabricated out of a 20 mm thick copper plate and are made rigid by a supporting arm. The four round stems support the Dee and provide the right inductance in order to get the required resonance frequency. Their size and location allow reaching the required Dee voltage profile versus radius. Each Dee will be equipped with vacuum pumping panels to efficiently pump the median plane. One Dee is equipped with one electrostatic deflector for extraction. This fact has not been taken into account in detail in the RF simulations to date, but some room for housing the deflector has been provisioned at the back of the Dee.
3.7.1 Integration of preliminary central region in RF Cavities
In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA into the RF cavity modelling, several technical choices were made to ensure the effective implementation of this innovative configuration:
  • The central region 3D model was limited to 200 mm from centre.
  • Both the Dee and counter dee original central region was removed by trimming the cavities at 800 mm from centre.
  • Sections were cut straight, and a linear transition (loft) technique was used to achieve the connections between the trimmed components.
These lofted transition leads to an acceptable linear vertical gap opening from 26 mm at a radius of 200 mm, to 50 mm at a radius of 800 mm. In addition, adequate clearance provisions were incorporated, specifically to fit the inflector and optics coil. Symmetry with regards median plane symmetry was maintained. Power bridges between dees in the central regions were removed since we recommend to power the 4 cavities independently.
Figure 3.16: dee at central region close view
Figure 3.17: dee lofted transition top view
Figure 3.18: vertical gap evolution close view
3.7.2 RFQ Integration in RF Cavities
Some clearance provision has been set for the integration of the RFQ, ass seen on figure 3.19
  • Diameter 280 mm
  • Down to 35 cm of median plane
While the RFQ will only be in the upper location, we have decided to keep median plane symmetry, for computation, manufacturing, and performances purposes. We have observed that the RFQ clearance has a low impact on cavity frequency and performances, and the output of the RFQ could be lowered more without impacting the cavity behavior.
Figure 3.19: Clearance for RFQ
3.7.3 RFQ getter pumps in RF Cavities
Mastering the quality of the high vacuum is key to the success of high-power beam cyclotrons. Incorporating getter pumps in the dees, in the case of IsoDAR, would allow to enhance vacuum quality. By integrating three housing for SAES UHV1400 getter pumps per dee, we aim to ensure highly efficient and reliable vacuum performance within the Cyclotron system. The use of six of those getter pumps per cavity facilitates the effective removal of residual gases, since achieving and maintaining ultra-high vacuum conditions essential for optimal particle acceleration. Furthermore, our preliminary assessments suggest that it may be feasible to scale up the configuration to include up to two sets of five getter pumps per cavity. The very low electrical field of the locations of the housings will allow the getter pumps to function safely.
Figure 3.20: Provisional getter housings
This early integration of the housings of the getters in the dees is mandatory to assess their impact on the cavity frequency, and on surfaces currents. However, to finalize this integration, comprehensive vacuum computations will be needed. These computations will enable us to determine the ideal amount of getter pumps, ensuring a properly engineered vacuum system that underpins the overall success and efficiency of the Cyclotron.
3.7.4 RF Cavities design optimization
3.7.4.1 Performance requirement target
We have optimized the cavity geometry to ensure the following requirements:
  • RF frequency of accelerating mode: 32.8 MHz
  • Reference accelerating voltages: 70 kV @ 110 mm, 230 kV @ 2000 mm
The initial target value for the outer radius voltage was initially set at 250 kV, but according to our expertise, such high voltage will be very difficult to stand in cyclotron accelerator gaps of this size. It has been decided to reduce the target to 230 kV which remains a challenging value.
3.7.4.2 Parametric optimization of the design
We have intensively used parameters while building the CST model, to allow fast design exploration. We have retained four parameters to play with, for optimizing the cavity behavior. The parameters, shown at figure 3.21, are defined as a deviation from the early cavity design values.
  • A: 1st pillar radial position
  • B: 1st pillar diameter
  • C: Deflector wall distance
  • D: Deflector height
Figure 3.21: design optimization parameters
CST studio suite offers an efficient optimization tools set. After some initial exploration of the values of the above-mentioned design parameters, we have optimized them using the CST optimizer.
Table 3.5: RF cavities optimization goals definition and achieved values.
Goal description Target Final Value
Ratio between voltage at 110 mm and 2000mm 0.304348 0.303991
Mode1 frequency divided by 32.8MHz 1 0.999871
We selected the trust region algorithm and have defined scaled goals that would ensure convergence toward the required frequency and voltages. The trust region algorithm converged in a few tens of iterations. The definition of the optimization goals, and their target and reached values are contained in table 3.5. The table 3.6 shows the optimal values reached by the optimizer.
Table 3.6: RF cavities optimized design parameters.
Parameter Value [mm] Description
Deflector offset (D) 5.67 Offset of upper and lower face of deflector
Increase_diameter (B) -3.81 Change of diameter of the pillar
Move_pillar (A) -124.1 Radial displacement of the pillar
Back_face_offset (C) 4 Outward move of the back of the deflector
3.7.5 RF cavity computation results
The RF cavity computation employing the eigen mode solver on the optimized geometry yielded the following performance and characteristics of the cyclotron system:
  • Frequency: The RF cavity operates at a frequency of 32.796 MHz, ensuring compatibility with the designed acceleration, and the RFQ integration,
  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.303991, indicating that computed voltages at center would be 69.9 kV for a chosen voltage of 230 kV at the outer radius (2m),
  • Quality Factor (Q): 9617,
  • Stored Energy (for max RF of 230 kV at 2 m): 3.8 Joules
  • CST Computed Losses: The computed losses per cavity using CST simulations are 81.3 kW,
  • Estimated Power Dissipation: By experience, senior RF designers at IBA state that the CST eigenmode simulations underestimate the actual losses by up to 30%. If we increase the CST computed losses (81.3 kW) by this percentage, the actual estimated power dissipation per cavity becomes 116.2 kW.
  • Estimated Power Requirement: To effectively drive the Cyclotron, each cavity will require an estimated power input of at least 266.2 kW, which include power dissipation and beam loading.
  • Maximal Surface Current: The maximal surface current reaches 24324.6 A/m in some reduced locations. Which leads to a maximum surface power of 88.4 W/cm². This maximal surface power is localized on the dee edges. Careful attention will be required for thermal management in these regions. For the rest of the surfaces, the thermal dissipation is well below 10 W/cm². See figure 3.26 and 3.27 While matching the requirement, these computation results provide valuable information for next steps of the Cyclotron’s RF cavity design including consideration in managing thermal effects.
Some illustrations of the electrical and magnetic fields computed by CST are given for reference at figures 3.22, 3.23, 3.24 and 3.25.
Figure 3.22: E field Vertical cross section
Figure 3.23: E-field Horizontal cross section
Figure 3.24: H-field horizontal cross section
Figure 3.25: H-field vertical cross section
Figure 3.26: Surface currents
3.7.5.1 Detailed power analysis
We will now detail the post treatment of the CST results that lead to the results presented in Sec. 3.7.5. The CST eigenmode solver results are computed for a given reference energy in the system (about 1J). They must then be scaled accordingly, to reach the expected voltage at a reference location. In our case, the reference location, is the orbit at 2m, and the desired voltage is 230 kV.
[…]
The ratio of the desired voltage to the reference voltage at 2m is 1.947. We can scale the currents and voltages using this factor. We can also scale the energies, power and losses using this factor squared (formulae 3.1, 3.2 and 3.3). We thus have:
$E_{stored}=E_{stored,ref}\frac{V_{target}^{2}}{V_{ref}^{2}}$ (3.1)
$P_{loss}=P_{loss,ref}{V_{target}^{2}\over V_{ref}^{2}}$ (3.2)
$I{{}_{s}}=I_{s,ref}{V_{target}\over V_{ref}}$ (3.3)
Table 3.7: Optimized cavities, functional characteristics
CST reference Scaled value
Stored energy [J] 1.001 3.796
Voltage at on the 2m orbit kV] 118.12 230
Voltage at center [kV] 35.91 69.9
Maximal voltage at outer edge of the dee [kV] 126.55 246.4
Computed losses [kW] 21.452 81.3
Maximal surface current [A/m] 12492.1 24324.6
[…]
$R_{s}={1\over\sigma.\delta}$ (3.4)
$\delta=\sqrt{2\over{\sigma.\mu.\omega}}$ (3.5)
$P_{s}=R_{s}I_{s}^{2}$ (3.6)
Table 3.8: Cavity properties for computing surface dissipation
Property Value
Copper magnetic permeability $\mu$ $0.99991\times 4\pi\times 10^{-7}$ H/m
Copper conductivity $\sigma$ $5.8\times 10^{7}$ S/m
Angular Frequency $\omega$ $2\pi\times 32.796\times 10^{6}$ rad/s
Skin Depth $\delta$ $11.53$ µm
3.7.5.2 Radial voltage dependency
The voltages at different radii have been calculated by integrating the electrical field along circular arcs crossing the acceleration gaps, and reference curves crossing central gap and the dee outer edge gap. The transit time factor has not been considered. The computed voltages have then been scaled to match a target voltage of 230 kV at 2 m. The resulting absolute and normalized voltage radial dependency are given in figure 3.28 and table 3.9
Figure 3.28: Dee voltage radial dependency
Table 3.9: RF voltage radial dependency.
Radius [mm] Voltage [kV]
center 69.9
110 69.918
200 68.04
300 69.149
400 69.187
500 73.054
600 78.677
700 84.162
800 91.511
900 102.353
1000 111.984
1100 123.881
1200 138.013
1300 151.364
1400 162.731
1500 177.622
1600 189.519
1700 203.636
1800 212.224
1900 222.408
2000 230
Outer radius 246.4
3.7.6 RF Tuning System
The tuning system is a movable plate or loop that corrects the resonance frequency drift mainly caused by RF heating. The system here is based on four separated tuning plates (one per cavity) and located in the median plane. The plate is accurately displaced by an electrical motor via a ball screw. Similar systems are already used many commercial cyclotrons.
3.7.7 RF Amplifiers
[…]
The traditional way to produce such a high RF power is still today the use of vacuum tubes (triodes or tetrodes) and this is especially true in the accelerator world where the load constraints are more severe than simple antennas. On the other hand solid state technology made a significant step forward in the past years in terms of power density and robustness. Even though the last generation of LDMOS allows to make solid state amplifiers at nearly competitive price with respect to vacuum tubes and that are also very reliable, we do believe vacuum tubes still present reliability and maintainability advantages for powering the isoDAR cyclotron, given the harsh environment where those amplifiers will be located. Tubes amplifiers will also accommodate easily to the rapid variations of load due to beam temporal pattern.
3.7.8 RF Lines and Couplers
The RF power is fed to the cavities through EIA 6 1/8” solid copper coax line. There are basically two ways to couple RF to the cavities:
  • Capacitively in the median plane
  • Inductively in the bottom (or top) of the cavity liner.
The choice between these two alternatives is mainly driven by the available space, the field patterns and the cooling concerns. The usual experience in this field is that both systems can work and couplers up to 800 kW at 107 MHz have been successfully built and operated.
Four couplers are needed for the complete cyclotron. The 300 kW power will be injected via one coupler per cavity.
Eventhough capacitive coupling might be less prone to multipacting by the fact that the coupling antenna could be easily DC biased, our expertise in high power coupler design advocates for an inductive solution. The design of inductive coupler being easier to watercool, and there is a wide area where H-field is high enough for efficient coupling behind the second pillar of the cavity.
We have initiated a preliminary design computation of such a inductance coupler, feeding the RF power from a EIA 6”1/8 port. The suggested design is a loop made out of a 20x60 mm rectangular section bar, offering room for water-cooling. See figure 3.29. The adaptation tuning will be performed by rotation of the loop around the vertical axis. The design has been optimized with a default angle of 45° to the cavity axis, in order to allow a maximal tuning range.
The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 362 mm, wich provides a good matching (figure 3.30) and a $S_{1,1}$ scattering parameter inferior to 0.01.
Figure 3.29: RF loop coupler design
Figure 3.30: Coupler port impedance
3.7.9 LLRF Electronics
The LLRF (Low Level RF) controls the RF amplitude and phase of the signal sent to the amplifiers. It also controls the tuning mechanism in order to maintain constant resonant frequencies in the cavities. The amplitude and phase control is based on feed-back signals coming from the cavities and amplifiers. In order to avoid the multipactor effect the LLRF manages also a pulsing mode that provides for an easy start-up and reduces the power dissipated in the amplifiers. One LLRF system per cavity is planned.
An additional electronic element is needed to keep a stable phase difference between the different cavities.
3.8 Cyclotron magnetic performance
In this section we describe the latest modifications applied to the cyclotron regarding its magnetic properties. We will compare the results with the model from the Conceptual Design Report (CDR) from [6] to ensure its magnetic performance remain within specifications.
With the latest design of the vacuum chamber, see section 3.4, the pole radius was slightly reduced. To keep the magnetic field isochronous up to extraction at 120 MeV, the outer radius was therefore somewhat modified, see left image in Fig. 3.31. First, the curvature was modified to follow the curvature of the last orbit. This enhances the total flux in the region where needed. Second, the top layer of the pole at its outer radius will be made of permendur. This alloy has better magnetic properties at high field, and enhances again the magnetic flux. Finally, the pole height is increased by a slanted ridge at its extremity. This reduction of the pole gap should have no effect on the transmission and can be tuned to improve the dB/dr for the last orbits.
Figure 3.31: Left: Detailed view of the outer radius of the pole; only half of one pole is shown. In red is the coil, in light green is the pole made of SNVA6342 steel. The turquoise part is a piece 30 mm in height and 200 mm in width made of permendur. In blue we see the pole edge made of SNVA6342 steel, which can be shimmed to isochronize the magnetic field. Right: BH curves for the two steel types used: Permendur and SNVA6342. As reference, the BH curve used for the cyclotron in the CDR is also shown.
The steel used for the cyclotron yoke and poles is the SNVA6342, also used for example in the C400 cyclotron (see Ref. [44]). Its BH-curve is shown in fig 3.31. Here, we also see the better performance of the permendur alloy.
3.8.1 Isochronization
The edge of the pole can be shimmed to isochronize the magnetic field map. In Fig. 3.32 the latest results are given. The left image shows the shimming applied, the right image shows the offset in magnetic field, as required for a given RF-frequency. For comparison, the offset (expressed as dB in Gauss) for the model of the CDR is also shown.
Figure 3.32: Left: Pole edge of isochronized model. Right: the dB [Gauss] as a function of radius [meters], for the CDR model (in black) and the newest model (in red). Note that a positive dB means there is too little field at specific radius, a negative dB implies there is too much field. The RF-frequency for the CDR model is at 32.77 MHz, for the newest model it is 32.82 MHz.
The plot shows that the magnetic field is stable and strong enough up to a radius of 2 meters, which is the equivalent of 121 MeV. In Fig 3.33 the phase slip can be seen to stay within the ±$\pm 10^{\circ}$. The horizontal and vertical tunes are also given in the same figure. Here we see that there is a crossing of the tunes at high radius, invoking resonances which could impact the extraction.
Figure 3.33: Left: phase slip as a function of radius. Right: vertical and horizontal tunes as a function of energy (dashed and solid lines respectively). Shown are the results for the CDR model (in black) and the newest model (in red).
As a conclusion, we can state that a configuration is possible with an isochronous magnetic field for up to 121 MeV. The transition to permendur at high radius, together with an abrupt slanted pole face, requires shimming of the pole edge. In a further study, these last modifications can be made more smooth, as not to have too much of a fluctuation in the dB for the last orbits and thus tunes that do not cross anymore, avoiding possible resonances.
Chapter 4 Installation in the Yemilab Setting
The most likely deployment scenario is at the Yemilab underground laboratory, 1 km below the surface adjacent to the Handuk iron mine in the Republic of Korea. Fig. 4.1 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 meters at a 15% slope from a surface access point into the mine, or by an elevator conveyance in a new 600 meter vertical shaft. Unfortunately, the elevator is suitable only for personnel access, all heavy and bulky equipment must be brought down through the ramp. Fig. 4.2 details the area where the IsoDAR experiment would be installed, with the target only a few meters from the planned 3 kiloton LSC (Liquid Scintillator Counter). Site-specific questions will be addressed in Volume III of this PDR and the following section aims at a general description of deployment strategies in an underground setting, pertaining only to the cyclotron. In this document we use Yemilab as the example, noting that most of the issues discussed will arise in an equivalent manner at other installation sites.
Figure 4.1: (a) Schematic of Handuk/Yemilab area, (b) Low resolution laser scan of Handuk mine and Yemilab access ramps, green area denotes high resolution scans of IsoDAR deployment area, shown in more detail in the next figure.
Figure 4.2: Layout of Yemilab, showing (in green) the area where IsoDAR will be installed. The cyclotron will be assembled in the dedicated cyclotron vault as shown.
4.1 Transport of Components
All of the components of the experiment must be brought to the site from the port-of-entry into the Republic of Korea. The preferred method is to use trucks along the highways. Prior experience of IBA with rail transport has been that it is difficult to provide adequate shock protection, and heavy pieces can be damaged in transit. As stated above, transport from the surface to the experimental site must be accomplished by way of the mine ramp depicted in Fig. 4.1. Nominally, its cross section is 5 x 5 meters, however we are told that pinch points and sharp bends may occur and might provide challenging areas to transport large, heavy loads through. We have received assurance that large mining trucks, 3.5 meters wide, can traverse the entire ramp, as this is the way ore was transported to the surface for many years.
While high-resolution laser scans of the IsoDAR and LSC areas have been performed, (cf. Figures 4.8 and 4.9), the resolution of the laser scans of the access ramp areas are not of sufficient resolution to provide enough detailed information about the location or size of the possible constrictions. Improved scans must be performed as part of final installation planning, and will be included in volume III which will be site-specific to Yemilab.
As seen in the previous chapter, the iron pieces that make up the cyclotron magnet, although heavy, are not in themselves large, so the pinch points are not likely to be a problem for bringing the magnet steel to the installation site. The plan is to use an SPMT vehicle (Self-Propelled Modular Transporter) shown in Fig. 4.3 for bringing the magnet and vacuum liner pieces from the surface. These electrically-powered vehicles are flexible in their size, the sets of 4 tires are independently steerable, so the vehicle is extremely maneuverable, and can easily accommodate the heaviest piece of magnet steel [45].
Figure 4.3: A typical SPMT vehicle, extremely versatile and capable of extremely heavy loads over irregular terrain
The only challenge will be the magnet coils. Each of the two coil packages measures 5 meters across and 0.3 meters high, the individual coils inside measure 4.960 x 0.252 meters. Figs. 4.5 and 4.5 gives an approximate view of a coil package and the clearances in the access ramp to the Laboratory (marked as the ”Entrance Tunnel” in Fig. 4.1 (a). This particular ramp is well-constructed and engineered, and has a measured width of 5 meters. A special mounting fixture capable of holding the coil packages at an angle should be able to bring these coils through this particular passageway. There are constrictions even in this ramp: cable trays and service boxes, however these can be moved if necessary. Uncertainty remains as to whether this type of conveyance device for the coils can navigate the winding ramp from the surface.
In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections. This has already been studied in Chapter 3 of the CDR document [46, 2]. The heavy winding machine and fixtures would need to be transported, but could be broken down into parts that will fit through the ramps, then assembled on site. The biggest challenge would be the ventilation scheme for the potting process. There is precedence for sectioned coils. For the large TRIUMF cyclotron [47], coils are separated into six different segments as the outer diameter of the coils is almost 20 meters, making the transport of fully assembled coils essentially impossible over conventional roadways. The inconvenience for a segmented coil, is that the ends of every turn of every section must be spliced, substantially increasing the cost and decreasing the efficiency of operation, as the coil must have as few turns as possible with very high current flowing through each turn.
Figure 4.4: Isometric view of the coil package, on a specially-designed cart, transporting the coil through the access ramp to the Yemilab campus.
Figure 4.4: Isometric view of the coil package, on a specially-designed cart, transporting the coil through the access ramp to the Yemilab campus.
Figure 4.5: This ramp is engineered to a 5-meter width, and maximum height of 5 meters. Unfortunately we do not have sufficient information about the winding access ramp from the surface to the base of the shaft to assess at this time whether transport of the coils will be possible with this conveyance device.
4.2 Assembly underground
The cyclotron will be assembled in a hall measuring 12 meters wide, 17 meters deep and 10 meters high. This hall was excavated to our specifications at the 90 degree bend in the access ramp to the LSC, as seen in Fig. 4.2. A flat area, roughly 10 m x 5 m at the corner will be the staging area for components transported from the surface. A bridge crane riding on tracks will be used to rig parts to the cyclotron assembly point. Conversations with a crane manufacturer, Industrias Electromecánicas GH, S.A. [48] gave us assurances that a 40-ton capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig. 4.6. It could be configured to fit the dimensions of the cavern, and broken into modular parts that could be transported to the site.
Figure 4.6: Examples of GH gantry cranes, capacities: foreground 20 tons, background 260 tons
Figure 4.7: Cyclotron and crane on cavern floor.
Figure 4.7: Cyclotron and crane on cavern floor.
Figure 4.8: Cyclotron and crane inside the cavern.
An overview of the cyclotron inside the cavern with the rigging equipment is shown in Figures 4.8 - 4.9. It should be noted that the gantry crane will be dismantled once the heavy parts are rigged. This would increase the clearance above the cyclotron for the additional equipment. In addition, the large crane can be reconfigured and reassembled in the target hall, and used for assembly of the MEBT, target and shielding, and for subsequent maintenance of the target systems.
Figure 4.9: Cycloton open and crane with cavern mesh - cut-view.
Chapter 5 Conclusion
5.1 Summary
In this version 0.9 of the document, we presented the preliminary design of the IsoDAR compact cyclotron with RFQ injection, including ion source and low energy beam transport.
The presented design study shows the feasibility of producing a high current (10 mA of protons through stripping of a 5 mA $\mathrm{H}_{2}^{+}$ beam) in a compact cyclotron. It further shows a convincing strategy to segment the cyclotron yoke for transportation into the underground environment to be assembled in situ at the final location.
RF calculations show manageable power and cooling, the required radial voltage distribution in the accelerating gaps, and adequate room for vacuum getter pumps.
In a future update to this PDR document, we plan to show a full start-to-end beam dynamics study, based on the presented individual components’ beam simulations. We will also add a risk assessment, and additional details on the extraction system, based on the cited publications and ongoing design work. Finally, we will include a vacuum calculation based on MolFlow.
5.2 Acknowledgements
This work was primarily supported by NSF grant PHY-2012897. Preliminary work leading up to this was supported by NSF grants PHY-1912764 and PHY-1626069. Moon was supported by DOE grant DE-SC0024138. Winklehner was supported by the Heising-Simons Foundation.

### v3

@@
 ToDo (#1): #2
-Abstract: This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source.
-Volumes I and II are site-independent and describe the cyclotron driver providing a 10 mA proton beam, and the medium energy beam transport line and target, respectively.
-Volume III describes the installation at the Yemilab underground laboratory in South Korea.
+Abstract: This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source in two volumes which are mostly site-independent and describe the cyclotron driver, providing a 10 mA proton beam, and the medium energy beam transport line and target, respectively.
 The IsoDAR driver and target will produce a mole of electron-antineutrinos over the course of five years.
@@
 Here in Volume I, we describe the driver, which includes the ion source, low energy beam transport, and cyclotron.
-The latter features radiofrequency quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of vortex motion.
-[…]
+The latter features radiofrequency quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of so-called vortex motion.
+Section Editors (in alphabetical order): M.
+Abs1, J.R Alonso2, J.M.
+Conrad2, S.J.B.
+Engebretson2, E.
+Forton1, A.
+Herrod1, D.
+Joassin1, J.
+Moon2, S.
+de Neuter1, E.
+Van der Kraaij1, G.
+Wéry1, D.
+Winklehner2, E.
+Winkler2
 for the IsoDAR collaboration
-${}^{1}$on Beam Applications, 3 chemin du cyclotron Louvain-la-Neuve, BELGIUM
-${}^{2}$assachusetts Institute of Technology, 77 Massachusetts Ave. Cambridge, MA, USA
+1Ion Beam Applications, 3 chemin du cyclotron Louvain-la-Neuve, BELGIUM
+2Massachusetts Institute of Technology, 77 Massachusetts Ave. Cambridge, MA, USA
 Corresponding Authors: Joshua Spitz (spitzj@umich.edu) and Daniel Winklehner (winklehn@mit.edu)
-V0.9
+V2.0
 Chapter 1 Introduction
@@
 From right to left: The compact cyclotron with RFQ injector, the medium energy beam transport, the target, and the LSC neutrino detector.
-From [1].
+From Ref.
+[calvo_analyzing_2023].
 IsoDAR will be the first high-intensity, proton-driven source to be located alongside an underground multi-kiloton-scale scintillator detector.
-The first version of IsoDAR will be constructed at the Yemilab facility [2] within the Handuk mine, located in South Korea where the “Liquid Scintillator Counter” (LSC) is planned to be installed [3].
+The first version of IsoDAR will be constructed at the Yemilab facility [alonso_isodaryemilab_2022] within the Handuk mine, located in South Korea where the “Liquid Scintillator Counter” (LSC) [^1] The LSC was renamed $\nu$Eye in March 2025, but for consistency throughout this document, we will refer to the detector as the LSC.
+is planned to be installed [seo_physics_2023].
 The novel source, consisting of a cyclotron, beam transport line, and a target/sleeve is shown in Fig.
@@
 In five calendar years (four live years), the source will produce 1.67 million $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ (inverse beta decay or ”IBD”) events and 7000 $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$ (elastic scattering or ES) events in the LSC that can be used for searches for new physics.
-More than a mole of neutrons and monoenergetic photons are also produced in the target, and these may act as progenitors for new particles that enter the LSC and leave signatures from interactions and decays.
+More than a mole of neutrons and mono-energetic photons are also produced in the target, and these may act as progenitors for new particles that enter the LSC and leave signatures from interactions and decays.
 IsoDAR can address a broad range of new physics manifested as new symmetries, new interactions and new particles that is inaccessible to traditional experiments.
 For further discussion see Refs.
-[4, 5].
-This document is Volume 1 in a three-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering solutions necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D.
+[alonso_neutrino_2022, waites_axionlike_2023].
+This document is Volume I in a two-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering solutions necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D.
 The contents of the PDR will cover the Technical Facility: the cyclotron, the beam transport and the target/sleeve.
-This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [6].
+This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [abs_isodarkamland_2015].
 This step of producing a PDR substantially reduces risk related to IsoDAR approval: the remaining detailed engineering that will follow will build on these results and be straightforward to estimate in scope and cost.
-Volume 1 describes the IsoDAR proton driver.
+Volume I describes the IsoDAR proton driver.
 This consists of a “front end” that delivers ions to a cyclotron for acceleration and extraction and the cyclotron itself.
 This document presupposes a basic understanding of cyclotrons; for an introduction to cyclotrons and a cost/benefit analysis of this choice of machine for IsoDAR, see the discussion in the CDR.
-The 60 MeV/amu IsoDAR Proton Driver is exceptional in that it is designed and engineered to produce an order of magnitude higher intensity proton beam at extraction than commercial machines of similar energy.
+The 60 MeV/amu IsoDAR Proton Driver is exceptional in that it is designed and engineered to produce an order of magnitude higher current proton beams at extraction than commercial machines of similar energy.
 Achieving this high power necessitated three key breakthroughs.
@@
 Second, the beam is axially injected through an RFQ, acting as a high-efficiency buncher.
-Third, the design harnesses vortex motion, a complex effect previously observed at PSI Injector II [7, 8], inducing stable spiraling motion in the high-current accelerated beam, significantly reducing beam growth.
+Third, the design harnesses vortex motion, a complex effect previously observed at PSI Injector II [stetson:vortex, baumgarten:vortex1], inducing stable spiraling motion in the high-current accelerated beam, significantly reducing beam growth.
 For more concerning the accelerator physics that shapes the IsoDAR design, see Ref.
-[9, 10].
+[winklehner_order--magnitude_2022, winklehner_high_2018].
 These three physics requirements for reaching high current drive many of the engineering choices presented in this PDR.
@@
 This PDR is structured as follows:
-Chapter 1:
-(This chapter) proceeds with an overview of the physics searches that will be possible with IsoDAR, followed by an overview of the requirements for the proton driver.
-Chapter 2:
-Describes engineering related to the IsoDAR front end.
-Chapter 3:
-Describes engineering related to the general IsoDAR cyclotron design.
-Chapter 4:
-Describes specifics of the design related to installation underground.
+  • Chapter 1: (This chapter) proceeds with an overview of the physics searches that will be possible with IsoDAR, followed by an overview of the requirements for the proton driver.
+  • Chapter 2: Describes engineering related to the IsoDAR front end.
+  • Chapter 3: Describes engineering related to the general IsoDAR cyclotron design.
+  • Chapter 4: Describes specifics of the design related to installation underground.
 1.1 The IsoDAR Physics Case
 Figure 1.2: Energy dependence of fluxes.
-Left: $\bar{\nu}_{e}$, unit normalized; Middle: photons per proton on target (POT) from excited nuclei.
+Left: $\bar{\nu}_{e}$, unit normalized; Middle: photons per proton on target (POT) from excited nuclei [alonso_neutrino_2022].
 Right: kinetic energy of neutrons at production in target and sleeve, normalized per proton.
-IsoDAR offers a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the proposed LSC at Yemilab [3].
+IsoDAR offers a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the proposed LSC at Yemilab [seo_physics_2023].
 The exceptional reach is owed to the high flux rates of $\bar{\nu}_{e}$, photons, and neutrons from the $\approx 2\cdot 10^{24}$ protons on target over 5 years (cf.
@@
 See Refs.
-[4, 11, 5] for further details.
+[alonso_neutrino_2022, Hostert:2022ntu, waites_axionlike_2023] for further details.
 1.1.1 Fluxes in IsoDAR
@@
 1.2.
-The $\bar{\nu}_{e}$ flux stems from a single beta-decay of ${}^{8}$i, which is produced by neutron capture on ${}^{7}$i and is well-understood.
-It peaks at 6 MeV, which puts it above backgrounds from the environment.
+The $\bar{\nu}_{e}$ flux stems from a single beta-decay of 8Li, which is produced by neutron capture on 7Li and is well-understood.
+It peaks at 6 MeV, which puts the majority of the events above backgrounds from the environment.
 These are clear advantages over reactor-neutrino fluxes.
@@
 We model the processes in the target with Geant4 using an accurate representation of the latest target and sleeve design (see Refs.
-[12]).
+[bungau_optimizing_2019]).
 We show the resulting neutron energy spectrum in Fig 1.2 (right).
-Most neutrons capture on the ${}^{7}$i in the sleeve.
-To ensure $<5$ neutrons per year enter the LSC veto region, the target is further shielded using boron-loaded concrete and steel [13].
+Most neutrons capture on the 7Li in the sleeve.
+To ensure $<5$ neutrons per year enter the LSC veto region, the target is further shielded using boron-loaded concrete and steel [bungau_shielding_2020].
 The third flux we consider is that of photons as seen in Fig.1.2 (middle).
 Here, one can observe a line structure that stems from the transitions of excited nuclei and and underlying smooth continuum due to bremsstrahlung.
-Again, the spectrum was generated using Geant4 using the full IsoDAR Target geometry [5].
+Again, the spectrum was generated using Geant4 using the full IsoDAR Target geometry [waites_axionlike_2023].
 1.1.2 New Particles: Experimentally and Theoretically Motivated
@@
 […]
-Figure 1.4: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with ${}^{8}$i-produced $\bar{\nu}_{e}$ background (gray).
-Right: $X$ mass and $X$-to-photon production ratio 90% CL sensitivity [4].
-Another example of IsoDAR’s capability to look for new particles is the “$X$,” which is a low mass mediator that would couple to the photons produced in the target [4].
+Figure 1.4: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with 8Li-produced $\bar{\nu}_{e}$ background (gray).
+Right: $X$ mass and $X$-to-photon production ratio 90% CL sensitivity.
+From Ref.
+[alonso_neutrino_2022].
+Another example of IsoDAR’s capability to look for new particles is the “$X$,” which is a low mass mediator that would couple to the photons produced in the target [alonso_neutrino_2022].
 The $X$ would be a light new particle that decays to $\bar{\nu}\nu$ and would leave a distinct signature above background in the IBD spectrum.
 Fig.1.4 (left) displays resulting signatures for various masses (see color bar) assuming 0.5% $X$-to-photon production.
-The IBD background from $\bar{\nu}_{e}$ stemming from ${}^{8}$i decays (gray) constitutes the primary background and has a distinct shape.
+The IBD background from $\bar{\nu}_{e}$ stemming from 8Li decays (gray) constitutes the primary background and has a distinct shape.
 The sensitivity is depicted in Fig.1.4 (right), where the blue region is excluded in the case of the SM.
@@
 1.1.3 New Symmetries: Strong CP Violation and the QCD Axion
-Figure 1.5: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [5].
+Figure 1.5: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [waites_axionlike_2023].
 Gold dots–axion allowed; Teal–excluded regions; Gray–parameters with astrophysical implications.
-The neutron beam/bottle experiment lifetime discrepancy [17, 18, 19, 20] motivates neutron–dark sector couplings.
+The neutron beam/bottle experiment lifetime discrepancy [Nico:2004ie, Yue:2013qrc, Serebrov:2017bzo, UCNt:2021pcg] motivates neutron–dark sector couplings.
 Through $n\rightarrow n^{\prime}\rightarrow n$ transitions, we may see neutrons disappearing from the target and reappear in the LSC.
-The underground environment provides a low background and the LSC’s size further adds to the sensitivity [11].
+The underground environment provides a low background and the LSC’s size further adds to the sensitivity [Hostert:2022ntu].
 Alternatively, if the $n^{\prime}$ are Majorana, then $n\rightarrow n^{\prime}\rightarrow\bar{n}$ may occur, depositing $\sim$2 GeV of energy in the LSC upon annihilation.
-This is a $B$ violating effect with implications for the early universe [21].
-The photons and electrons in the target may also couple to axions and IsoDAR can probe uncharted territory with astrophysical and cosmological implications [5].
+This is a $B$ violating effect with implications for the early universe [Berezhiani:2020vbe].
+The photons and electrons in the target may also couple to axions and IsoDAR can probe uncharted territory with astrophysical and cosmological implications [waites_axionlike_2023].
 In Fig.
@@
 This constitutes the world’s largest sample at low momentum transfer.
-LEP and LHC data predict rate and energy dependency very precisely within the standard model [22].
+LEP and LHC data predict rate and energy dependency very precisely within the standard model [ParticleDataGroup:2020ssz].
 With the normalization determined by the $\approx 1.7$M IBD events in the LSC, we will be able to measure $\sin^{2}\theta_{W}$ with a precision of 1.9% (see Fig.1.6 left).
@@
 Statistics limit the outcome, and further running can proceed with IsoDAR and DUNE as it comes online if a NuTeV-like deviation is observed.
-Figure 1.6: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation. Right: expectation for IsoDAR in NSI parameter space. See Ref. [4].
+Figure 1.6: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation.
+Right: expectation for IsoDAR in NSI parameter space.
+From Ref.
+[alonso_neutrino_2022].
 The accurately predicted energy dependence of the ES sample enables a precise phenomenological search for non-standard neutrino interactions (NSIs).
@@
 1.6 (right), where red assumes background reduction using Cherenkov light, and green assumes only scintillation light reconstruction.
-This can be compared to the world result [23] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments and, hence, is rotated with respect to IsoDAR, which has a pure $\bar{\nu}_{e}$ flux.
+This can be compared to the world result [guzzo] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments and, hence, is rotated with respect to IsoDAR, which has a pure $\bar{\nu}_{e}$ flux.
 1.2 Requirements for the Proton Driver
@@
 The proton beam is shaped, with quadrupoles and wobbler magnets to produce a distribution that allows for optimal heat transfer to the cooling water.
-Neutrons produced in the target system are moderated and flood into a roughly-spherical sleeve, about 1 meter in diameter, containing a mixture of beryllium chips and highly-enriched ${}^{7}$i.
-The beta-decay of the ${}^{8}$i resulting from the capture of neutrons produces the neutrinos of interest in the experiment.
+Neutrons produced in the target system are moderated and flood into a roughly-spherical sleeve, about 1 meter in diameter, containing a mixture of beryllium chips and highly-enriched 7Li.
+The beta-decay of the 8Li resulting from the capture of neutrons produces the neutrinos of interest in the experiment.
 The principal requirements for the accelerator system are to reliably produce a 10 mA beam of protons in the most efficient manner possible with the lowest possible loss of particles in the acceleration process.
-Experience from the cyclotron system at PSI [24] are that keeping the total power of beam loss inside the cyclotron vault to less than 200 watts still allows access for component maintenance.
-Considering that the total beam power is 600 kilowatts, one sees that this is a very stringent requirement.
-As mentioned above, Volume 1 of this PDR addresses the physics and engineering aspects of the 60 MeV/amu cyclotron.
-The final chapter addresses the complex issue of transport and assembly of the large and heavy cyclotron components in the underground environment at Yemilab.
-Subsequent volumes will address the details of the beam transport system and the target, along with shielding requirements for the environment of the target including the prevention of fast neutrons from reaching the fiducial volume of the liquid scintillator detector and the underground installation procedure, power and cooling, HVAC, and other site-specific topics.
+Particularly above the Coulomb threshold.
+The experience of the cyclotron system at PSI [Stammbach] is that keeping the total power of beam loss inside the cyclotron vault to less than 200 W still allows access for component maintenance.
+Considering that the total beam power is 600 kW, one sees that this is a very stringent requirement.
+As mentioned above, Volume I of this PDR addresses the physics and engineering aspects of the 60 MeV/amu cyclotron.
+The final chapter addresses the complex issue of transport and assembly of large and heavy cyclotron components in the underground environment at Yemilab.
+A forthcoming Volume II will address the details of the beam transport system and the target, along with shielding requirements for the environment of the target, including the prevention of fast neutrons from reaching the fiducial volume of the liquid scintillator detector.
+1.3 Overview of the Accelerator Layout
+Figure 1.7: Overview of the components of the HCHC-60 Cyclotron System. Only one RF cavity is shown to not clutter the image.
+To guide the reader through this PDR and, indeed, the path of an ion from genesis in the ion source to the full 60 MeV/amu and extraction, we provide an overview in Fig.
+1.7.
+The main components of the system, which will be described in much detail in the following chapters, are:
+  • The ion source, providing 10 mA of $\mathrm{H}_{2}^{+}$ at an emittance $<0.1\pi$-mm-mrad.
+  • The Low Energy Beam Transport (LEBT), shaping and guiding the beam to the RFQ and providing diagnostics, and chopping for commissioning and personnel- and machine protection.
+  • The Radiofrequency Quadrupole (RFQ), embedded in the cyclotron yoke, to bunch and pre-accelerate the beam to 35 keV/amu.
+  • The central region with electrostatic quadrupoles for transverse beam focusing, the spiral inflector to bring the beam from the vertical axis onto the median plane, and the specific RF electrode (aka “dees”) shapes to center the beam during the first turns.
+  • The main magnet and RF Cavities for acceleration to 60 MeV/amu.
+  • (not shown) The extraction system consisting of electrostatic and magnetic channels guiding the beam out of the cyclotron with the goal of extracting 5 mA of $\mathrm{H}_{2}^{+}$.
 Chapter 2 Design of the Front End
 2.1 The Ion Source
-The $\mathrm{H}_{2}^{+}$ beam to be injected into the cyclotron and accelerated to 60 MeV/amu will be produced in a filament-driven multicusp ion source which we will refer to as MIST-1.
-The technical details of this ion source are reported in Refs.
-[25, 26, 27, 28, 29] and are summarized here.
-2.1.1 Design
-[…]
-2.1.1.1 Gas Input
-Hydrogen gas is introduced via an inlet in the backplate of the source as indicated in Fig.
-2.2.
-The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 SCCM full range.
-This system interfaces with the PC control system which is documented in detail in Ref.
-[29].
-2.1.1.2 Ionization by the Filament
-After the hydrogen gas is introduced into the chamber, it is ionized to form a plasma using electrons released from a tungsten filament by thermionic emission.
-The filament is mounted on DI water-cooled power feedthroughs in the back plate (cf.
-Fig.
-2.2).
-The filament is heated by an (8.4V, 300 A max) power supply.
-In addition to this power supply which is used to heat the filament, a second power supply is connected to the filament and to the source chamber to maintain a potential difference of 80-140V between them in order to facilitate electron discharge.
-The source body, back plate, and front plate are all mutually electrically insulated from each other for added flexibility in the electric field shape.
-A schematic of the wiring of the ion source is illustrated in Fig.
-2.2
-[…]
-Figure 2.1: A 3D CAD rendering of the MIST-1 ion source.
-Figure 2.1: A 3D CAD rendering of the MIST-1 ion source.
-Figure 2.2: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Thick blue lines are data cables. Black lines are power cables.
-2.1.1.3 Confinement & Extraction of the Plasma
-[…]
-The electrons are then shaped and accelerated out via an extraction system (cf.
-Fig.
-2.2).
-The extraction system comprises five electrodes which shape and accelerate the ion beam.
-These include a plasma electrode (+15 kV), a puller electrode (-3 to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV).
-The RFQ entrance flange is held at ground potential.
-We simulated beam formation and shaping in the extraction system with the IBSimu package [30], which performs 3D simulations including space charge.
-Due to the computational resources needed to perform full 3D simulations in IBSimu, we adopted a hybrid IBSimu (for the plasma modeling)/WARP (for the beam transport) [31] model for the final design.
-We show example trajectories in Fig.
-2.3.
-Figure 2.3: Cross section of the extraction system showing the ion beam moving through the extraction system prior to entering the RFQ.
-Simulation of the extraction system performed by IBSimu and WARP software.
+The $\mathrm{H}_{2}^{+}$ beam to be injected into the cyclotron and accelerated to 60 MeV/amu will be produced by a filament-driven multicusp ion source.
+We have built a first prototype, the MIST-1 (Multicusp Ion Source at MIT - version 1), the technical details and first measured results of which are reported in Refs.
+[axani_high_2015, winklehner_first_2018, winklehner_high-current_2021-1, winklehner_new_2022, weigel_epics_2023].
+Here we summarize these results and present ongoing upgrade work (MIST-2).
+The upgrade to MIST-2 improves on the MIST-1 by implementing several lessons learned during the systematic measurements and runtime experiences described below.
+A publication on the MIST-2 design and performance is forthcoming.
+2.1.1 Operating Principle
+In a multicusp ion source [ehlers:multicusp1], a plasma is confined by a set of permanent magnets that are arranged around the plasma chamber with alternating polarity, thereby creating “cusps” in the magnetic field.
+On average, the electrons and ions produced in the center of the chamber see an increasing magnetic field as they travel outward, which eventually turns them back.
+One notable exception are the cusps, which are the natural loss points of this configuration.
+If the source is filament-driven [^1] Other options exist, e.g., RF-driven., electrons are emitted from a hot filament.
+A plasma discharge is initiated and sustained by applying a potential difference between the filament and (parts of) the source chamber.
+Thus, the filament acts as the cathode, and the source chamber acts as the anode of a discharge.
+Electron-impact ionization with neutral molecules and atoms in the source chamber produces secondary electrons and ions.
+Neutral gas is fed into the source to facilitate the generation of a high-density plasma.
+The type of gas determines the ion species produced in the source.
+Here we use H2 gas to form a plasma – and beam – comprising protons, $\mathrm{H}_{2}^{+}$, and H+${}_{3}^{+}$.
 Table 2.1: A summary of the parameters for the MIST-1 ion source.
@@
 Plasma chamber diameter 15 cm
-[…]
+Permanent magnet material Sm2Co17
 Permanent magnet strength 1.05 T on surface
@@
 Filament heating current Max. 100 A
+[…]
+2.1.2 Mechanical Design
+Figure 2.1: Cross section of MIST-1 (left) with extraction system and LEBT. Also shown is an example set of ion trajectories simulated with IBSimu and WARP. From Ref. [waites_high_2022].
+For visual reference as we discuss the details of the ion source, 3D CAD renderings of MIST-1 (with subsequent LEBT) and MIST-2 are provided in Figs.
+2.1 and 2.2, respectively.
+The main ion source parameters are also summarized in Table 2.1.
+The MIST-type ion sources comprise a plasma chamber surrounded by a “water-jacket,” a back plate, and a front plate.
+The water jacket lets us pump de-ionized (DI) water around the plasma chamber for cooling.
+The radial confinement magnets are mounted inside the water-jacket and are cooled directly.
+At 6.5 cm, the plasma chamber is short compared to other ion sources.
+This is a direct consequence of the short mean free path of $\mathrm{H}_{2}^{+}$ in the plasma/neutral gas of the ion source discussed above.
+The nominal configuration for the confinement magnets is: Four rows of alternating magnets on the back plate, twelve alternating radial magnets, and a star formation of twelve smaller magnets on the front plate, also alternating in polarity.
+This leaves a small volume of low magnetic flux around the extraction aperture, allowing the ions to easily drift out of the source.
+All magnets are Sm2Co17.
+For MIST-2 we have begun studying different configurations (no magnets, Nd2Fe14B magnets).
+See below for more details.
+The back plate holds four strips of confinement magnets, as well as the water-cooled filament feedthroughs, and the gas inlet (a KF NW-16 elbow).
+Temperature sensors (thermocouples) can be attached to the back plate to monitor the temperature of feedthroughs and magnets.
+The back plate itself is also water-cooled with cooling channels embedded in the plate.
+The filament is mounted to the high-current feedthroughs using molybdenum filament holders.
+The filament itself is made from tungsten and can have different thickness as well as shapes (this is a parameter currently being optimized).
+The front plate holds the front confinement magnets as well as the plasma aperture, a 75%W-25%Cu (aka “Elkonite”) plate that is embedded in the front plate.
+We have several plasma apertures with different-size openings ranging from 4 mm to 8 mm to test beam quality and total extractable current for various aperture diameters.
+As seen in Figs.
+2.1 and 2.2, following the plasma aperture is the “extraction system” which shapes and guides the beam as it is initially formed.
+The MIST-1 extraction system comprises five electrodes which shape and accelerate the ion beam.
+These include a plasma electrode (+15 kV), a puller electrode (-3 to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV).
+Figure 2.2: MIST-2 CAD rendering calling out the most important features.
+2.1.2.1 MIST-2 vs. MIST-1
+The changes we made in the MIST-2 design are mainly to improve cooling, assembly and maintenance, and diagnostics. They are:
+  • Replace all stainless steel components with OFHC copper to improve heat transfer
+  • Increase the cross-section areas of all water cooling channels
+  • New high-power feedthroughs that are detachable from the back plate
+  • An additional port on the back plate for a Langmuir probe
+  • A redesigned extraction system.
+The new pentode extraction system comprises a water-cooled puller, an intermediate electrode, a negative electrode and a grounded final electrode.
+The more compact extraction system with negative potential followed by ground will help keep space charge compensation high in the LEBT and allows us to use a Faraday cup and emittance scanners for beam diagnostics directly after the source.
+These changes will improve long-term running stability, reduce maintenance downtime, and allow for a better understanding of the plasma and beam conditions.
+Our high-fidelity simulations using IBSimu [kalvasIBSIMUThreedimensionalSimulation2010a] and WARP [vayNovelMethodsParticleInCell2012] suggest that the beam dynamics downstream of the six-way cross will not change between the MIST-1 and MIST-2 configurations.
+All particle simulations and designs still hold for LEBT, RFQ, and Cyclotron.
+2.1.3 Electrical Design
+Hydrogen gas is introduced via an inlet in the backplate of the source as indicated in Fig.
+2.2.
+The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 SCCM full range.
+This system interfaces with the EPICS PC control system which is documented in detail in Ref.
+[weigel_epics_2023].
+The filament is heated by an TDK Lambda (8.4 V, 300 A max) power supply.
+In addition to this power supply which is used to heat the filament, a second power supply (Matsusada REK) is connected to the filament and to the source chamber to maintain a potential difference of 80 - 140 V between them in order to facilitate electron discharge of up to 24 A max (typical during our tests was 5-10 A).
+The source body, back plate, and front plate are all mutually electrically insulated from each other for added flexibility in the electric field shape.
+Power is transferred to the platform via two isolation transformers of 2 kVA each.
+A schematic of the wiring of the ion source is illustrated in Fig.
+2.3.
+The ion source (platform) is lifted to high voltage using a Matsusada AU 15 mA, 20 kV power supply, all lens elements in the extraction system are powered by Matsusada AU 7.5 mA, 20 kV power supplies of either negative or positive polarity.
+The devices on the HV platform (control computer, filament discharge, filament heating power supply, etc.) are inside a Faraday cage on platform potential to reduce electromagnetic interference (EMF).
+Surrounding the platform and ion source is another Faraday cage at ground potential for personnel protection and EMF reduction.
+Figure 2.3: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Green lines are data connections. Black lines are power cables.
+2.1.4 Beam Dynamics
+We simulate beam formation and shaping in the extraction system with the IBSimu package [kalvasIBSIMUThreedimensionalSimulation2010a], which performs cylindrically symmetric 2D (aka “RZ”) and 3D simulations including space charge through an iterative process.
+IBSimu also contains a 1D plasma sheath model which is a good approximation for the plasma-beam interface at the source aperture.
+Our extraction simulations of the MIST-1 ion source followed by a WARP [vayNovelMethodsParticleInCell2012] simulation of the LEBT yielded excellent agreement with measurements [winklehner_first_2018, winklehner_new_2022].
+Example trajectories of the new MIST-1 extraction system and RFQ matching LEBT are shown in Fig.
+2.1.
+An extraction system simulation of the MIST-2 system is shown in Fig.
+2.4.
+The simulated normalized RMS emittance for the new MIST-2 extraction system using an 8 mm diameter plasma aperture is 0.077 $\pi$-mm-mrad.
+Simulated normalized RMS emittances for the MIST-1 extraction system were between 0.05 and 0.1 $\pi$-mm-mrad depending on source aperture and extracted current and agreed well with measurements [winklehner_first_2018].
+Figure 2.4: IBSimu Simulation of MIST-2 extraction. Left: trajectories, Right: x-x’ phase space. Both MIST-1 and MIST-2 exhibit a very low emittance (see text).
 2.1.5 Performance
@@
   • High purity: H+${}_{2}^{+}$ fraction $>$ 80%
-[…]
+  • Low emittance: $<$ 0.1 $\pi$-mm-mrad (RMS, normalized)
   • Ability to run continuously in DC mode
@@
 The total beam current is measured using the first Faraday cup located posterior to the extraction system.
-The dipole magnet with separation slips separates the ions by mass and passes them to the second Faraday cup which, in conjunction with the Allison scanners, measures each species horizontal and vertical emittances.
+The dipole magnet separates the ions by mass and passes them to the second Faraday cup for mass analysis.
+Allison-type electrostatic emittance scanners measure each species’ horizontal and vertical emittances.
 2.1.5.2 Optimization of Ion Source Parameters
 The ion source is designed such that several parameters can be varied in order to optimize the magnitude, purity, and emittance tightness of the current from the ion source.
-[…]
+  • H2 input flow rate
   • Filament discharge voltage
   • Filament discharge current
-  • Filament size, shape, & position (Optimization not yet studied)
-  • Permenant magnet type (Optimization not yet studied)
-[…]
-Filament discharge voltage optimization.
-Two series of filament discharge voltage optimizations have been performed differing in how discharge current is managed.
+  • Filament size, shape, & position (Optimization not yet fully studied)
+  • Permanent magnet type
+H2 input flow optimization.
+Three series of optimizations on the input gas flow rate have been performed.
+The three series differ in how the filament discharge voltage is managed.
+In the first series, the PID loop which stabilizes discharge was disabled allowing the discharge current to rise to the source steady-state value and then stabilized at that value by reengaging the PID.
+In the second series, the discharge current was kept fixed at 4 A.
+In the third series, the discharge current was reduced to keep the total extracted current fixed at 1.25 mA.
+Under each of these three conditions, the gas input rate was varied from 0.25 - 1.25 SCCM.
+This study demonstrates a general trend of greater H+${}_{2}^{+}$ production at higher discharge currents and, as theoretically expected, a higher fraction of H+${}_{3}^{+}$ with higher flow rates.
+Discharge voltage optimization.
+Two series of discharge voltage optimizations have been performed differing in how discharge current is managed.
 In the first, discharge current is permitted to freely find a steady state using the same method as used in the flow optimization described above.
 In the second, the discharge current was fixed at 4 A.
-Filament discharge current optimization.
+Discharge current optimization.
 One series of current optimizations was explored in which the flow rate and discharge were kept constant while discharge current was varied from 2 to 8 A.
@@
 Optimization has been performed for maximum total current density, maximum H+${}_{2}^{+}$ purity, and a hybrid option optimizing for the maximum H+${}_{2}^{+}$ current density with the constraint that H+${}_{2}^{+}$ be dominant.
-[…]
-[…]
-[…]
+Highest total current density: 41 mA/cm2 with 31% H+${}_{2}^{+}$ purity.
+Highest H+${}_{2}^{+}$ purity: 76% purity with 11.4 mA/cm2 total current density.
+Highest H+${}_{2}^{+}$ current density with H+${}_{2}^{+}$ dominant: 10.4 mA/cm2 with 62% purity [^2] Note that this is the current density of $\mathrm{H}_{2}^{+}$ alone..
 With current optimizations and a 4mm aperture, the ion source is capable of delivering approximately 1.1 mA of H+${}_{2}^{+}$ with 76% purity.
-With an 8 mm aperture this corresponds 4.4 mA.
-This is a world leading result and is only about a factor of 2 short of target.
-The main bottleneck for increased beam current was the extraction system, which was designed for initial testing and commissioning.
+With an 8 mm aperture this corresponds to 4.4 mA.
+This is a world-leading result and is only about a factor of 2 short of the target.
+The main bottleneck for an increased beam current was the extraction system, which was designed for initial testing and commissioning.
 We have recently replaced it with the version presented in this chapter.
-Our high fidelity simulations suggest that the new extraction system will be able to handle the nominal beam current.
+Our high-fidelity simulations suggest that the new extraction system will be able to handle the nominal beam current.
 Measurements are planned for summer 2024.
+2.1.5.4 Magnet Study
+The choice of permanent magnet material and strength is one of the optimization parameters of this type of ion source.
+Here, plasma density, confinement time, and magnet thermal properties are important parameters to consider.
+Originally, we selected Sm2Co17 magnets for their higher Curie temperature.
+However, long-term running tests suggest that the magnet temperature is well controlled, which would allow the use of Nd2Fe14B magnets, if they yielded a better beam.
+We aim to maximize the $\mathrm{H}_{2}^{+}$ fraction of the MIST-1 ion source.
+Within the plasma, atomic processes include ionization, dissociation, and recombination, leading to protons, $\mathrm{H}_{2}^{+}$, and H+${}_{3}^{+}$ in the beam.
+Magnets with a higher field strength yield higher plasma density and longer confinement of the ions in the plasma, which can alter the rates of these processes.
+Thus, we are interested in how different magnet types alter the ion fraction and total current.
+We investigated the current and ion fraction of the beam with no magnets, Sm2Co17 magnets (grade 28), and Nd2Fe14B (grade N50H).
+The configuration without magnets provides no confinement, and the Nd2Fe14B magnets provide the strongest confinement.
+We performed our measurements using the diagnostic beam line presented in Ref.
+[winklehner_first_2018], containing a dipole magnet and a Faraday cup with 8 mm entrance slit for mass analysis.
+We investigated each magnet type at a range of H2 mass flows (0.25 SCCM to 1.5 SCCM) and discharge voltages (60 V to 140 V).
+In our preliminary analysis, we observed a trade-off between the total beam current and the $\mathrm{H}_{2}^{+}$ fraction: with lower confinement, the $\mathrm{H}_{2}^{+}$ fraction was high, but the current was low.
+This result was expected; lower confinement and lower plasma density leads to less time and fewer reactants for the ions to undergo processes that result in $\mathrm{H}_{2}^{+}$ dissociation or the pick-up of another proton (producing H+${}_{3}^{+}$).
+However, the lower plasma density also leads to a lower total beam current.
+For MIST-2, the baseline design is now to use Nd2Fe14B magnets because of the higher current they provide. A paper on the detailed analysis of the magnet study is forthcoming.
+2.1.5.5 Results with the MIST-2 back plate
+Figure 2.5: A typical spectrum with the new MIST-2 back plate.
+The total beam current was 3.43 mA, the $\mathrm{H}_{2}^{+}$ fraction $\approx 60$%.
+For a discussion about the small contaminant peak, cf.
+text.
+We already performed the first test of the MIST-2 components described in Section 2.1.2.1 by installing the new back plate (which is backward compatible) on MIST-1.
+We were able to run the source up to a stable beam current of 3.43 mA.
+We did this with the source set to 7.5 kV (the nominal source voltage of 15 kV requires a full beam current of 10 mA to form the appropriate plasma meniscus), puller set to -5 kV, lens 1 at 1.5 kV, lens 2 at -3 kV, lens 3 at 0.5 kV and the filament heating set to 150 A.
+Due to the improved cooling and filament feedthroughs and holders, the source ran very stably; as we expected.
+A sample spectrum from this series of studies is shown in Fig.
+2.5.
+Here, a small cluster of beam contaminants can still be seen, which we attribute to insufficient baking and out-gassing of source components, including Viton O-rings.
+Past experience has shown that vacuum baking the O-rings before installation (done here), overall good vacuum conditions, and baking the source and beam line before operation (not done in this instance), decreases the contaminants to a negligible level after a few hours of operation [winklehner_first_2018].
+2.1.6 Risks and Mitigation
+Risk: Ion Source Long-Time Stability.
+As IsoDAR is designed to run continuously for 5 years, the long-time stability is a major concern for the ion source.
+Hands-on maintenance of the ion source is going to be limited, because of the close proximity of the cyclotron.
+Mitigation: Before the final technical design, we will develop alternative heating methods (external antenna and 2.45 GHz microwaves).
+MIST-1 and MIST-2 were being designed with an exchangeable backplate for quick filament replacement and easy further development of alternative heating methods.
+Another option is a dual-source operation using a switching magnet.
+Risk: Ion source maximum $\mathrm{H}_{2}^{+}$ current is too low. The current densities seen from the MIST-1 ion source are still a factor 2 below the requirements for IsoDAR.
+Mitigation: Our current upgrade path should lead us comfortably to a total extracted current of 12 mA, particularly considering that MIST-1 has demonstrated an exceptionally low emittance, which opens up the possibility of going to even larger extraction apertures (10 mm or 12 mm).
+However, we also have tested a 2.45 GHz flat-field ECR ion source (the VIS [alonso_isodar_2015]) in the past, from which we measured 15 mA of $\mathrm{H}_{2}^{+}$.
+The low ratio of 50% $\mathrm{H}_{2}^{+}$ (30 mA total current) deterred us, but as a mitigation, a VIS-type ion source paired with a dipole magnet is a viable alternative.
 2.2 The Low Energy Beam Transport
@@
     Provide safety mechanisms for personnel and machine protection.
-Details of the LEBT design and performance were given in Refs. [32, 33, 34] here we summarize and update these.
+Details of the LEBT design and performance were given in Refs. [waites_matching_2021, waites_low_2022, waites_high_2022] here we summarize and update these.
 2.2.1 Design Overview
@@
 The beam now passes through a gap surrounded by a series of six electrodes arranged axially within a six-way cross (see Fig.
-2.2).
-These six electrodes can have their voltage optimized in order to best match the parameters for acceptance into the RFQ.
+2.1).
+These six electrodes can have their voltage optimized in order to best match the parameters for acceptance into the rest of the LEBT.
 A conical flange is connected to the end of the six-way cross.
 The conical shape minimizes beam loss on exit.
+Figure 2.6: Left: Photo of Alpha Magnetics, Inc.
+steerers [2562NotreDame].
+Right: Result of LEBT steering simulations.
+It can be seen that several millimeters offset can be achieved without significant change in emittance.
+From Ref.
+[waites_high_2022].
 Following the six-way cross lies a diagnostic segment.
@@
 The diagnostic section also includes a solenoid and two magnetic steerers which can be used for additional beam tuning.
-A flange in place toward the end allows space for an ACCT which characterizes the beam immediately prior to entering the RFQ.
+A flange in place toward the end allows space for an AC Current Transformer (ACCT) [bergozACCT] which characterizes the beam immediately prior to entering the RFQ.
 The diagnostic segment is illustrated in Fig.
-2.3.
+2.1.
 2.2.2 Beam Steering & Chopping
@@
 Even small, sub-millimeter misalignment can prevent efficient injection of the beam into the RFQ.
-In order to ensure that the beam enters the RFQ at both the correct position and angle, two magnetic steerers and a solenoid are placed around the beamline.
+In order to ensure that the beam enters the RFQ at both the correct position and angle, two magnetic steerers and a solenoid magnet are placed around the beamline.
 These are controlled independently by an external power source to produce tunable dipole fields which optimize beam entry into the RFQ.
-The optimal fields were calculated using COMSOL and Warp.
-These are discussed further in the following section.
+The optimal fields were calculated using COMSOL and beam dynamics simulated in WARP.
+The effect of the steerers is demonstrated in Fig.
+2.6.
 In order to control the amount of power delivered from the ion source, a beam chopper is built into the cross of the diagnostic segment to allow for a duty factor in an otherwise DC beam.
 The beam chopper is built into the cross in the diagnostic section.
-The chopper comprises a copper ring surrounded by a steel beampipe which in turn is surrounded by a copper water cooling jacket.
-The copper ring can be put at a high potential relative to the cross, which is grounded.
-If the potential is sufficiently high the beam can be diverted such that it terminates.
+The chopper comprises a deflection plate, followed by a copper ring for particle dumping press-fitted into a steel beam pipe which in turn is surrounded by a copper water cooling jacket.
+The deflection plate can be put at a high potential relative to the cross, which is grounded, diverting the beam such that it terminates on the copper ring.
 See Fig.
-2.5.
-This method permits control over the amount of power delivered to the RFQ without disrupting the plasma itself.
+2.8.
+This method permits controlling the average beam power, without changing the charger per bunch resulting from the RFQ.
+This is required to maintain vortex motion in the cyclotron.
 A 12 mA beam of 15 keV particles fully terminating on the beam stop corresponds to $\sim$180 W power delivered.
 We need to ensure that the dump is capable of dissipating this amount of power.
-The system was modeled assuming a water cooling line at room temperature.
+The system was modeled by assuming a water cooling line at room temperature.
 To be conservative, the simulation was run using four times the actual anticipated beam power, i.e.
@@
 The COMSOL simulation is shown in Fig.
-2.5
+2.8
 Figure 2.7: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid.
 If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring.
-Figure 2.4: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid.
-If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring.
-Figure 2.5: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior.
+From Ref.
+[waites_high_2022].
+Figure 2.8: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior.From Ref. [waites_high_2022].
 2.2.3 Beam Dynamics
@@
 IBSimu provides a more accurate plasma model, but is computationally intensive and so was used only for the region by the ion source where the plasma is extracted and the density is thus high, requiring this accuracy.
-IBSimu is used for the first several mm after the extraction aperture where density remains high.
+IBSimu is used for the first several millimeters after the extraction aperture, where density remains high.
 After this, and for the remainder of the downstream region of the LEBT, the distributions from IBSimu are handed off to Warp which provides accurate results for the lower density regions of the beam with less computational burden.
-For the IBSimu simulations a 12 mA beam composed of 80% H+${}_{2}^{+}$, 10% ${}^{+}$, and 10% H+${}_{3}^{+}$ was used.
+Space charge is included in all calculations.
+However, as is well-known, in the absence of external electrical fields, slow secondary charged particles from the interaction of the beam with the residual gas can be accumulated inside the beam envelope, effectively neutralizing the beam charge.
+This process is called “space charge compensation” [winklehnerSpaceChargeCompensation2015].
+In all LEBT simulations, we estimate space charge compensation according to the model presented in Ref.
+[winklehnerSpaceChargeCompensation2015].
+It is usually above 80%, alleviating the detrimental effect of space charge in some parts of the LEBT.
+However, we also test our designs with no compensation and find that, while beam losses are slightly higher, we can still obtain the required beam inside the cyclotron, albeit requiring about 20% higher beam from the source.
+For the IBSimu simulations a 12 mA beam composed of 80% H+${}_{2}^{+}$, 10% H+, and 10% H+${}_{3}^{+}$ was used.
 This provides ample current for IsoDAR assuming 50% global transmission efficiency while using current from the ion source that has already been empirically demonstrated.
 As illustrated in Fig.
-2.6 and in Table 2.2 the LEBT is able to match the required input parameters for the RFQ.
+2.9 and in Table 2.2 the LEBT can match the required input parameters for the RFQ.
 Moreover, as illustrated in Fig.
 2.10 this match is achieved with emittance that surpasses that used for baseline RFQ simulations.
-Figure 2.6: A temperature plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT.
-Figure 2.7: Phase space from the LEBT simulation (left) vs phase space for ideal beam input to the RFQ (right).
+Figure 2.9: A temperature plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT. [waites_high_2022]
+Figure 2.10: Phase space from the LEBT simulation (left) vs phase space for ideal beam input to the RFQ (right). From Ref. [waites_high_2022].
 Table 2.2: A comparison of LEBT Twiss parameters and baseline RFQ parameters.
@@
 $\beta$ 0.13 mm/mrad 0.17 mm/mrad
+2.2.4 Risks and Mitigations
+Risk: Unexpected Power Loss.
+In the case of an unexpected loss of power to one of the power supplies, the beam dynamics could be very different from nominal and lead to damage downstream (e.g., the RFQ).
+Mitigation: Careful simulation of various failure modes showed that the LEBT is robust under such changes.
+The time constant involved with heating up of materials are such that the beam can be switched off before damage occurs.
+However, for the MIST-2 design, we are envisioning the addition of a Faraday cup directly downstream of the extraction system.
+Operating spring-loaded, it can be inserted quickly to intercept the beam early on.
+Risk: Lower than anticipated space charge compensation.
+Optimal beam transport was calculated for a high compensation factor.
+If this is not the case, transport efficiency and beam dynamics may deteriorate.
+Mitigation: The LEBT design was double-checked with no space charge compensation and care was taken that injection into the RFQ is nevertheless possible.
+Typically, higher losses, which are still manageable, are seen.
+The ion source must provide about 20% more beam current in this case.
 2.3 The Radiofrequency Quadrupole
-Figure 2.8: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle.
+Figure 2.11: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle. From Ref. [waites_high_2022].
 The $\mathrm{H}_{2}^{+}$ beam coming from the LEBT is injected into a Radio Frequency Quadrupole (RFQ) that is partially embedded in the cyclotron yoke.
@@
 We show a cut view of the RFQ in Fig.
-2.8.
+2.11.
 Of note are the small diameter of the RFQ (27.6 cm) while operating at a low resonant frequency of 32.8 MHz.
@@
 The technical design was presented in Refs.
-[35, 36, 37] and the optimization of the RF loop coupler in Ref.
-[38].
+[ratzingerTechnicalDesignReport2020, winklehner_high-current_2021, holtermann_technical_2021] and the optimization of the RF loop coupler in Ref.
+[sangroula_design_2021].
 The thermal properties and cooling considerations were presented in Ref.
-[39].
+[koser_thermal_2021].
 An optimization scheme using machine learning was presented in Ref.
-[40].
-Figure 2.9: Modes of the mechanical deformation of the RFQ vanes. The frequencies are 18 Hz and 25 Hz.
+[koser_input_2022].
+Figure 2.12: Modes of the mechanical deformation of the RFQ vanes. The frequencies are 18 Hz and 25 Hz. From Ref. [holtermann_technical_2021].
 In terms of mechanical stability, our analyses show two quarter wave modes which are associated with the choice of a split-coaxial resonator.
 The addition of bridges between opposite vanes (see Fig.
-2.9) significantly reduces detuning due to the Lorentz force with mechanical frequencies around 18 Hz and 25 Hz.
+2.12) significantly reduces detuning due to the Lorentz force with mechanical frequencies around 18 Hz and 25 Hz.
 2.3.2 Beam Dynamics
 Beam dynamics studies were performed of the full 3D particle distribution resulting from the LEBT simulations going through the RFQ, including the effects of the entrance gap and asymmetry (due to the split coaxial design) and the exit gap and asymmetry.
+This was performed by the subcontracted company Bevatech GmbH in Germany, who is currently also manufacturing the RFQ.
+The well-established tracking code TraceWin was used [tracewin].
 Fine tuning of the phase of the exit gap (i.e.
@@
 $\varepsilon_{z}$ (1-RMS, norm.) 0.114 MeV-deg
-Phase width (1-RMS) ${}^{\circ}$
+Phase width (1-RMS) 40
+2.3.3 Risks and Mitigation
+Risk: New RFQ Technology.
+Direct axial injection into a compact cyclotron has never been demonstrated.
+It is possible that there are limitations that have not been thought of, that will prove to be show-stoppers.
+Mitigation: A prototype machine is currently being built and will be thoroughly tested in the coming year.
+As a fallback, conventional LEBTs using a multi-harmonic buncher are well understood and reliable.
+However, they are much less efficient than RFQs.
+The focus in this case will shift to increasing ion source performance further to allow for higher beam losses during injection.
+Risk: RFQ Power Estimate The RFQ costs depend strongly on the RF power necessary to drive the rods/vanes, the length of the RFQ, the engineering complexity involved in machining the vanes/rods and designing the vacuum vessel.
+Mitigation: We have worked with RF engineers with intimate knowledge of RFQ design. The upcoming prototype will demonstrate the power requirements.
+Risk: Beam spread between RFQ exit and spiral inflector.
+The simulations showed that the beam spreads transversally and longitudinally after exiting the RFQ.
+Reality might prove more unforgiving than computer simulations and we may incur higher beam losses than anticipated.
+Mitigation: Our high-fidelity simulations suggest that the beam spread is manageable.
+The prototype will allow measuring this precisely.
+Additional focusing elements can be inserted between RFQ and spiral inflector if necessary.
+More aggressive collimation is possible as well, at the expense of requiring higher current from the source.
 Chapter 3 Design of the Cyclotron
-Figure 3.1: Isometric view of the cyclotron
-In this section we describe the details of the cyclotron design at the level of a preliminary design report.
+Figure 3.1: Isometric view of the cyclotron.
+In this section, we describe the details of the cyclotron at the level of a preliminary design report.
+This section includes details of the design of the cyclotron, the spiral inflector that transmits ions from the RFQ into the cyclotron, and the extraction system.
+This section also characterizes the latest in simulo performance of these systems.
 3.1 Magnet Yoke
-The magnet design presented here will necessarily be modified for a final design, as the specific details of the extraction orbit and all required extraction elements have not been finalized.
-The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron, which is slightly larger, at 7 meters outer diameter.
-Figure 3.2: C400 on the milling machine, C230 in front of it.
+The magnet design presented here is physically realistic, but is necessarily preliminary.
+We do not anticipate the broad design of the magnet to change in future.
+As we finalize the details of beam extraction small modifications to the magnets, especially at the fringes, may accordingly be modified.
+The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron (seen in Fig 3.2), which is slightly larger, at 7 meters outer diameter.
+Figure 3.2: C400 on the milling machine, C230 in front of it. Courtesy of IBA.
 3.1.1 Design and Fabrication of the Cyclotron Magnet
-The cyclotron is divided in the same fashion as the C400, into smaller parts facilitating transport to the site.
+As with the C400, the IsoDAR cyclotron is divided into smaller parts to facilitate transport to the site.
 The magnet is split radially and will be assembled on site.
-The vacuum chamber is split in 3 (self-centering) parts along its height, the vacuum tightness is ensured by 2 axial O-rings at each interface except for the interface between the upper and lower chambers were 2 radial O-rings are used.
-An exploded of the upper and lower parts is shown in Figs.
+The vacuum chamber is split into 3 self-centering parts along its height.
+The vacuum tightness is ensured by 2 axial O-rings at each interface except for the interface between the upper and lower chambers where 2 radial O-rings are used.
+An exploded visualization of the upper and lower parts is shown in Figs.
 3.4 and 3.4.
-The materials used for the magnet parts is cyclotron steel (the exact composition shall be determined after a detailed physics study), the vacuum chambers will be made of stainless steel.
-Figure 3.3: Cyclotron upper steel exploded.
+The material used for the magnet parts is cyclotron steel (the exact composition shall be determined after a detailed physics study), the vacuum chambers will be made of stainless steel.
 Figure 3.3: Cyclotron upper steel exploded.
 Figure 3.4: Cyclotron lower steel exploded.
-An overview of a cut in the cyclotron is shown in Fig.
+An cut-section view of the cyclotron is shown in Fig.
 3.5.
 The external and vacuum disks are connected to each other as well as to the flux return using specific bolts (such as the Superbolt from Nord-Lock).
-The poles are split in several parts and an easily removable part should be added in the next iteration to ease the mapping and possibly improve the magnetic performance.
+The poles are split into several parts and an easily removable part should be added in the next iteration to ease the mapping and possibly improve the magnetic performance.
 A detailed view of the cyclotron components is shown in Fig.
-3.6.
+3.6 and additional details about the yoke parts can be found in Table 3.1
 Figure 3.5: Cyclotron cut view
 Figure 3.6: Cyclotron detailed cut view
-Table 3.1 gives some more information regarding the yoke parts.
 Name Mass (Metric Ton) Dimensions (mm)
@@
 Table 3.1: Yoke part properties
-3.2 Interface with the RFQ
-Figure 3.7: RFQ overview
-The RFQ will be outside of the cyclotron’s vacuum, therefore a vacuum sleeve will be used to ensure the integrity of the cyclotron’s vacuum.
-This sleeve will be fixed on top of the upper external disk and the vacuum sealing will be done radially on the upper internal disk (as seen in figure number 3.9 ).
-The O-rings could have a cross section diameter of 8mm and an inner diameter of 305mm.
-In order to allow some adjustement, the RFQ sleeve allows a rotation of 5mrad around the lower O-ring contact patch (discussed with the manufacturer of the RFQ).
-Figure 3.8: RFQ zoom
-The RFQ sleeve is as thin as possible to avoid any loss in magnetic field in the central region (hence the O-rings in the Yoke parts).
-Figure 3.9: RFQ cut section 1
-Fig.
-3.10 shows that the vacuum sealing will be done at the bottom of the RFQ using a thick O-ring (diameter 10 mm) which will be compressed by the RFQ’s weight ($\approx 3000$ N).
-The dimensions of the vacuum sleeve are listed in Tab.
-3.2.
-The adjustment feet shown in Fig. 3.8 are fixed to the cyclotron yoke and allow the following alignment modifications through four sets of four set screws one and vertical adjustment nut:
-  • Rotation: ±5plus-or-minus5$\pm$ 5± ${}^{\circ}$
-  • Tilt: ±$\pm 5$ mrad
-  • Lateral movement: ±$\pm 1$ mm
-In addition to this fine-tuned alignment, a coarse alignment step is envisioned during the initial assembly and commissioning procedure.
-Figure 3.10: RFQ cut section 2
-Table 3.2: Vacuum Sleeve Dimensions.
-Name Value
-Height 1030.5mm
-Outer diameter 304.55mm
-Inner diameter 296.55mm
-Weight 18kg
-3.3 Central Region
-Figure 3.11: CAD rendering of the spiral inflector and particle trajectories for early tests at Best Cyclotron Systems Inc.
-[41].
-The positive and negative electrostatic deflectors are shown in blue and green respectively.
-The cyclotron magnetic field in this image is directed vertically upwards.
-Particles enter the spiral inflector via the rectangular grounded collimator (solid gray) and are guided into the cyclotron mid-plane by means of the cyclotron magnetic field and the electrostatic potential between the electrodes.
-The copper housing (transparent gray) isolates the spiral inflector from the RF fields driving the cyclotron.
-The injection system design is a crucial component to achieve the highest beam current.
-Our design balances injecting the $\mathrm{H}_{2}^{+}$ molecule at high energy against avoiding the use of complex high-voltage platforms.
-Higher energy injection mitigates space charge and to reduces geometrical (non-normalized) emittance.
-An energy of 35 keV/amu was selected which determined the output energy of the RFQ.
-As in commercial cyclotrons, an axial injection system based on a Spiral Inflector (SI) will be used to bend the beam from the axial direction to the median plane.
-This concept is visualized in Figure 3.11 using trajectories calculated with OPAL [42].
-The exit point of the spiral inflector is also the starting point of acceleration in the Central Region (CR) of the cyclotron.
-The central region and the spiral inflector have to be designed carefully and together because they have strong interplay.
-This procedure has already been performed and tested for the injection tests during the summers of 2013 and 2014 at the Best Cyclotron Systems Inc.
-(BCSI) laboratory in Vancouver.
-In the following subsection, we will describe the design process of the central region for the BCS test stand cyclotron followed by a preliminary design for the full IsoDAR machine.
-3.3.0.1 Spiral Inflector Design
-The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Figure 3.11).
-The electrostatic potential between these electrodes is able to bend the beam ${}^{\circ}$ from the axial line to the median plane of the cyclotron.
-The helical trajectory of the beam is determined both by the shape of the electrodes (electric field) and the magnetic field of the cyclotron as the beam is bent into the median plane.
-The applied voltage depends on the velocity of the beam as well as the rigidity of the ions.
-Figure 3.12: Photograph of the spiral inflector used during the tests at Best Cyclotron Systems.
-The spiral inflector tested at BCSI (shown in the photograph in Figure 3.12) was designed to mimic the conditions required for the IsoDAR inflector, its primary defining characteristic being the 15 mm gap between the electrodes.
-Compared to other similar designs, this is rather large because it has to take into account the larger beam size due to space-charge effects and the higher rigidity of $\mathrm{H}_{2}^{+}$.
-Due to the large distance between the electrodes, the SI occupies a volume in which the magnetic field components’ variation is not negligible, as it is assumed in the analytical treatment of the spiral inflector.
-This effect has to be carefully taken into account in order to shape correctly the spiral inflector electrodes and to avoid the introduction of a high energy-spread during the beam transport through the device.
-The spiral inflector has a copper housing that surrounds the electrodes to minimize the interaction between the electric fields generated by the dees and the electrostatic fields between the electrodes.
-To minimize beam striking the uncooled electrodes, a water-cooled, grounded, rectangular collimator shields the spiral inflector entrance.
-The preliminary design of the spiral inflector called for a nominal voltage of ±$\pm 11$ kV and a tilt angle of ${}^{\circ}$ for optimal beam injection at 60 keV.
-The total height of the device was approximately 80 mm.
-These parameters were found by using a MATLAB code based on an analytic theory for spiral inflectors.
-The electrode shape was calculated using VectorFields OPERA.
-This design was later modified to account for some effects due to fringe fields on the electrodes.
-One of these modifications was to reduce the overall length of each electrode by 5 mm.
-First turn acceleration is achieved by the dee tips extending into recesses in the spiral inflector housing.
-The shapes of the dee tips and the spiral inflector housing have been designed to guide the particles from the spiral inflector exit to the acceleration region while providing the necessary energy gain and beam focusing.
-[…]
-The main differences between the BCSI test stand design and the IsoDAR design are:
-  • The IsoDAR cyclotron will be equipped with 4 accelerating RF cavities, while the test cyclotron used only 2 cavities.
-  • The harmonic mode (and the frequency) will be decreased from 6th harmonic (49.2 MHz) to 4th harmonic (32.8 MHz).
-  • The injection energy will be increased from 30 keV/amu to 35 keV/amu.
-  • The spiral inflector gap will be increased from 15 mm to 16.5 mm to accommodate a larger beam.
-3.4 Vacuum Chamber
-The vacuum enclosure is composed of 5 parts : the upper and lower internal disk and 3 vacuum chambers.
-The vacuum sealing is done using 2 metallic O-rings per interface (from Omniseal [https://www.omniseal-solutions.com/] for example) except for the radial interface where Polymer O-rings could be used.
-The interface for the accessories will be done using a sleeve design and radial O-ring, described in the figure hereunder.
-Figure 3.13: Cyclotron vacuum chamber details
-Name Mass (Metric Ton) Dimensions (mm)
-Upper and lower internal disk 36 Ø4620*463
-Upper vacuum chamber 2.3 Ø4452*483
-Lower vacuum chamber - median plane 2.3 Ø4452*483
-Lower vacuum chamber - internal disk 1.8 Ø4386*380
-Table 3.3: Vacuum part properties
-3.5 Coil configuration
-[…]
-Figure 3.14: Left: schematic view of a cross-section of the main coil, with six pancakes per half coil.
-Right: possible configuration of the winding, using a square conductor with a round cooling channel in its center.
-[…]
-Working in pancakes gives the possibility of manufacturing the pancakes outside the cave and stacking them inside.
-The increase in height from 200 to 240 mm is the equivalent of adding an extra pancake.
-3.6 Cyclotron RF System
-Table 3.4 provides a detailed description of the RF system. The system for IsoDAR cyclotron is divided into subsystems as follows:
-  • The RF cavities allowing a fast acceleration to the final particle energy and keeping a large inter-turn separation.
-The cavities are triangular shaped and are located in the magnet valleys.
-There are 4 accelerating electrodes (Dees).
-  • The RF tuning system that will compensate for thermal drift of cavities and maintain a stable resonating frequency.
-  • The RF amplifiers that provide the necessary power to create the required RF electric field inside the cavity and the power to accelerate the beam.
-  • The RF couplers that inject the RF power from the amplifiers into the cavities
-  • The Low Level RF system (LLRF) that controls and regulates the RF amplitude on the Dee and drives the tuning system.
-Table 3.4: Details of the RF design.
-RF System Component Design Value
-Resonance Frequency 32.8 MHz
-Dee voltage in the central region $70$ kV
-Dee voltage at extraction radius $230$ kV
-Dee radial extension 2 m
-Dee angular extension ${}^{\circ}$
-Cavity height 1800 mm
-Number of Dee stems 4 per Dee (2 up, 2 down)
-Number of Dee’s 4
-Acceleration harmonic 4th
-Power dissipated per cavity 116 kW
-Cavity Q-factor 9620
-We consider these in the Sections below.
-3.7 RF Cavities (Dees, Stems and Liners)
-Figure 3.15: Early RF Dee computation – color scheme represent current density
-The RF cavities of the cyclotron have been calculated using CST microwave studio [43] in order to define the shape of the Dee and the Dee stems.
-They have then been optimized also with CST from this early design.
-The optimization work is intended to reach the right resonance frequency, the Dee voltage Law, together with the minimal power dissipation and hence the highest quality factor (Q-factor).
-The field maps allow also the calculation of the couplers and voltage pick-up but this last calculation is yet to be performed.
-The Dees and cavity walls are all made of OFHC copper that combines good electrical conductivity and good solder ability.
-The Dees are fabricated out of a 20 mm thick copper plate and are made rigid by a supporting arm.
-The four round stems support the Dee and provide the right inductance in order to get the required resonance frequency.
-Their size and location allow reaching the required Dee voltage profile versus radius.
-Each Dee will be equipped with vacuum pumping panels to efficiently pump the median plane.
-One Dee is equipped with one electrostatic deflector for extraction.
-This fact has not been taken into account in detail in the RF simulations to date, but some room for housing the deflector has been provisioned at the back of the Dee.
-3.7.1 Integration of preliminary central region in RF Cavities
-In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA into the RF cavity modelling, several technical choices were made to ensure the effective implementation of this innovative configuration:
-  • The central region 3D model was limited to 200 mm from centre.
-  • Both the Dee and counter dee original central region was removed by trimming the cavities at 800 mm from centre.
-  • Sections were cut straight, and a linear transition (loft) technique was used to achieve the connections between the trimmed components.
-These lofted transition leads to an acceptable linear vertical gap opening from 26 mm at a radius of 200 mm, to 50 mm at a radius of 800 mm.
-In addition, adequate clearance provisions were incorporated, specifically to fit the inflector and optics coil.
-Symmetry with regards median plane symmetry was maintained.
-Power bridges between dees in the central regions were removed since we recommend to power the 4 cavities independently.
-Figure 3.16: dee at central region close view
-Figure 3.17: dee lofted transition top view
-Figure 3.18: vertical gap evolution close view
-3.7.2 RFQ Integration in RF Cavities
-Some clearance provision has been set for the integration of the RFQ, ass seen on figure 3.19
-  • Diameter 280 mm
-  • Down to 35 cm of median plane
-While the RFQ will only be in the upper location, we have decided to keep median plane symmetry, for computation, manufacturing, and performances purposes.
-We have observed that the RFQ clearance has a low impact on cavity frequency and performances, and the output of the RFQ could be lowered more without impacting the cavity behavior.
-Figure 3.19: Clearance for RFQ
-3.7.3 RFQ getter pumps in RF Cavities
-Mastering the quality of the high vacuum is key to the success of high-power beam cyclotrons.
-Incorporating getter pumps in the dees, in the case of IsoDAR, would allow to enhance vacuum quality.
-By integrating three housing for SAES UHV1400 getter pumps per dee, we aim to ensure highly efficient and reliable vacuum performance within the Cyclotron system.
-The use of six of those getter pumps per cavity facilitates the effective removal of residual gases, since achieving and maintaining ultra-high vacuum conditions essential for optimal particle acceleration.
-Furthermore, our preliminary assessments suggest that it may be feasible to scale up the configuration to include up to two sets of five getter pumps per cavity.
-The very low electrical field of the locations of the housings will allow the getter pumps to function safely.
-Figure 3.20: Provisional getter housings
-This early integration of the housings of the getters in the dees is mandatory to assess their impact on the cavity frequency, and on surfaces currents.
-However, to finalize this integration, comprehensive vacuum computations will be needed.
-These computations will enable us to determine the ideal amount of getter pumps, ensuring a properly engineered vacuum system that underpins the overall success and efficiency of the Cyclotron.
-3.7.4 RF Cavities design optimization
-3.7.4.1 Performance requirement target
-We have optimized the cavity geometry to ensure the following requirements:
-  • RF frequency of accelerating mode: 32.8 MHz
-  • Reference accelerating voltages: 70 kV @ 110 mm, 230 kV @ 2000 mm
-The initial target value for the outer radius voltage was initially set at 250 kV, but according to our expertise, such high voltage will be very difficult to stand in cyclotron accelerator gaps of this size.
-It has been decided to reduce the target to 230 kV which remains a challenging value.
-3.7.4.2 Parametric optimization of the design
-We have intensively used parameters while building the CST model, to allow fast design exploration.
-We have retained four parameters to play with, for optimizing the cavity behavior.
-The parameters, shown at figure 3.21, are defined as a deviation from the early cavity design values.
-  • A: 1st pillar radial position
-  • B: 1st pillar diameter
-  • C: Deflector wall distance
-  • D: Deflector height
-Figure 3.21: design optimization parameters
-CST studio suite offers an efficient optimization tools set. After some initial exploration of the values of the above-mentioned design parameters, we have optimized them using the CST optimizer.
-Table 3.5: RF cavities optimization goals definition and achieved values.
-Goal description Target Final Value
-Ratio between voltage at 110 mm and 2000mm 0.304348 0.303991
-Mode1 frequency divided by 32.8MHz 1 0.999871
-We selected the trust region algorithm and have defined scaled goals that would ensure convergence toward the required frequency and voltages.
-The trust region algorithm converged in a few tens of iterations.
-The definition of the optimization goals, and their target and reached values are contained in table 3.5.
-The table 3.6 shows the optimal values reached by the optimizer.
-Table 3.6: RF cavities optimized design parameters.
-Parameter Value [mm] Description
-Deflector offset (D) 5.67 Offset of upper and lower face of deflector
-Increase_diameter (B) -3.81 Change of diameter of the pillar
-Move_pillar (A) -124.1 Radial displacement of the pillar
-Back_face_offset (C) 4 Outward move of the back of the deflector
-3.7.5 RF cavity computation results
-The RF cavity computation employing the eigen mode solver on the optimized geometry yielded the following performance and characteristics of the cyclotron system:
-  • Frequency: The RF cavity operates at a frequency of 32.796 MHz, ensuring compatibility with the designed acceleration, and the RFQ integration,
-  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.303991, indicating that computed voltages at center would be 69.9 kV for a chosen voltage of 230 kV at the outer radius (2m),
-  • Quality Factor (Q): 9617,
-  • Stored Energy (for max RF of 230 kV at 2 m): 3.8 Joules
-  • CST Computed Losses: The computed losses per cavity using CST simulations are 81.3 kW,
-  • Estimated Power Dissipation: By experience, senior RF designers at IBA state that the CST eigenmode simulations underestimate the actual losses by up to 30%.
-If we increase the CST computed losses (81.3 kW) by this percentage, the actual estimated power dissipation per cavity becomes 116.2 kW.
-  • Estimated Power Requirement: To effectively drive the Cyclotron, each cavity will require an estimated power input of at least 266.2 kW, which include power dissipation and beam loading.
-  • Maximal Surface Current: The maximal surface current reaches 24324.6 A/m in some reduced locations.
-Which leads to a maximum surface power of 88.4 W/cm².
-This maximal surface power is localized on the dee edges.
-Careful attention will be required for thermal management in these regions.
-For the rest of the surfaces, the thermal dissipation is well below 10 W/cm².
-See figure 3.26 and 3.27 While matching the requirement, these computation results provide valuable information for next steps of the Cyclotron’s RF cavity design including consideration in managing thermal effects.
-Some illustrations of the electrical and magnetic fields computed by CST are given for reference at figures 3.22, 3.23, 3.24 and 3.25.
-Figure 3.22: E field Vertical cross section
-Figure 3.23: E-field Horizontal cross section
-Figure 3.24: H-field horizontal cross section
-Figure 3.25: H-field vertical cross section
-Figure 3.26: Surface currents
-3.7.5.1 Detailed power analysis
-We will now detail the post treatment of the CST results that lead to the results presented in Sec.
-3.7.5.
-The CST eigenmode solver results are computed for a given reference energy in the system (about 1J).
-They must then be scaled accordingly, to reach the expected voltage at a reference location.
-In our case, the reference location, is the orbit at 2m, and the desired voltage is 230 kV.
-[…]
-The ratio of the desired voltage to the reference voltage at 2m is 1.947.
-We can scale the currents and voltages using this factor.
-We can also scale the energies, power and losses using this factor squared (formulae 3.1, 3.2 and 3.3).
-We thus have:
-$E_{stored}=E_{stored,ref}\frac{V_{target}^{2}}{V_{ref}^{2}}$ (3.1)
-$P_{loss}=P_{loss,ref}{V_{target}^{2}\over V_{ref}^{2}}$ (3.2)
-$I{{}_{s}}=I_{s,ref}{V_{target}\over V_{ref}}$ (3.3)
-Table 3.7: Optimized cavities, functional characteristics
-CST reference Scaled value
-Stored energy [J] 1.001 3.796
-Voltage at on the 2m orbit kV] 118.12 230
-Voltage at center [kV] 35.91 69.9
-Maximal voltage at outer edge of the dee [kV] 126.55 246.4
-Computed losses [kW] 21.452 81.3
-Maximal surface current [A/m] 12492.1 24324.6
-[…]
-$R_{s}={1\over\sigma.\delta}$ (3.4)
-$\delta=\sqrt{2\over{\sigma.\mu.\omega}}$ (3.5)
-$P_{s}=R_{s}I_{s}^{2}$ (3.6)
-Table 3.8: Cavity properties for computing surface dissipation
-Property Value
-Copper magnetic permeability $\mu$ $0.99991\times 4\pi\times 10^{-7}$ H/m
-Copper conductivity $\sigma$ $5.8\times 10^{7}$ S/m
-Angular Frequency $\omega$ $2\pi\times 32.796\times 10^{6}$ rad/s
-Skin Depth $\delta$ $11.53$ µm
-3.7.5.2 Radial voltage dependency
-The voltages at different radii have been calculated by integrating the electrical field along circular arcs crossing the acceleration gaps, and reference curves crossing central gap and the dee outer edge gap.
-The transit time factor has not been considered.
-The computed voltages have then been scaled to match a target voltage of 230 kV at 2 m.
-The resulting absolute and normalized voltage radial dependency are given in figure 3.28 and table 3.9
-Figure 3.28: Dee voltage radial dependency
-Table 3.9: RF voltage radial dependency.
-Radius [mm] Voltage [kV]
-center 69.9
-110 69.918
-200 68.04
-300 69.149
-400 69.187
-500 73.054
-600 78.677
-700 84.162
-800 91.511
-900 102.353
-1000 111.984
-1100 123.881
-1200 138.013
-1300 151.364
-1400 162.731
-1500 177.622
-1600 189.519
-1700 203.636
-1800 212.224
-1900 222.408
-2000 230
-Outer radius 246.4
-3.7.6 RF Tuning System
-The tuning system is a movable plate or loop that corrects the resonance frequency drift mainly caused by RF heating.
-The system here is based on four separated tuning plates (one per cavity) and located in the median plane.
-The plate is accurately displaced by an electrical motor via a ball screw.
-Similar systems are already used many commercial cyclotrons.
-3.7.7 RF Amplifiers
-[…]
-The traditional way to produce such a high RF power is still today the use of vacuum tubes (triodes or tetrodes) and this is especially true in the accelerator world where the load constraints are more severe than simple antennas.
-On the other hand solid state technology made a significant step forward in the past years in terms of power density and robustness.
-Even though the last generation of LDMOS allows to make solid state amplifiers at nearly competitive price with respect to vacuum tubes and that are also very reliable, we do believe vacuum tubes still present reliability and maintainability advantages for powering the isoDAR cyclotron, given the harsh environment where those amplifiers will be located.
-Tubes amplifiers will also accommodate easily to the rapid variations of load due to beam temporal pattern.
-3.7.8 RF Lines and Couplers
-The RF power is fed to the cavities through EIA 6 1/8” solid copper coax line. There are basically two ways to couple RF to the cavities:
-  • Capacitively in the median plane
-  • Inductively in the bottom (or top) of the cavity liner.
-The choice between these two alternatives is mainly driven by the available space, the field patterns and the cooling concerns.
-The usual experience in this field is that both systems can work and couplers up to 800 kW at 107 MHz have been successfully built and operated.
-Four couplers are needed for the complete cyclotron. The 300 kW power will be injected via one coupler per cavity.
-Eventhough capacitive coupling might be less prone to multipacting by the fact that the coupling antenna could be easily DC biased, our expertise in high power coupler design advocates for an inductive solution.
-The design of inductive coupler being easier to watercool, and there is a wide area where H-field is high enough for efficient coupling behind the second pillar of the cavity.
-We have initiated a preliminary design computation of such a inductance coupler, feeding the RF power from a EIA 6”1/8 port.
-The suggested design is a loop made out of a 20x60 mm rectangular section bar, offering room for water-cooling.
-See figure 3.29.
-The adaptation tuning will be performed by rotation of the loop around the vertical axis.
-The design has been optimized with a default angle of 45° to the cavity axis, in order to allow a maximal tuning range.
-The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 362 mm, wich provides a good matching (figure 3.30) and a $S_{1,1}$ scattering parameter inferior to 0.01.
-Figure 3.29: RF loop coupler design
-Figure 3.30: Coupler port impedance
-3.7.9 LLRF Electronics
-The LLRF (Low Level RF) controls the RF amplitude and phase of the signal sent to the amplifiers.
-It also controls the tuning mechanism in order to maintain constant resonant frequencies in the cavities.
-The amplitude and phase control is based on feed-back signals coming from the cavities and amplifiers.
-In order to avoid the multipactor effect the LLRF manages also a pulsing mode that provides for an easy start-up and reduces the power dissipated in the amplifiers.
-One LLRF system per cavity is planned.
-An additional electronic element is needed to keep a stable phase difference between the different cavities.
 3.2 Cyclotron magnetic performance
 In this section we describe the latest modifications applied to the cyclotron regarding its magnetic properties.
-We will compare the results with the model from the Conceptual Design Report (CDR) from [6] to ensure its magnetic performance remain within specifications.
+We will compare the results with the model from the Conceptual Design Report (CDR) from [abs_isodarkamland_2015] to ensure its magnetic performance remains within specifications.
 With the latest design of the vacuum chamber, see section 3.5, the pole radius was slightly reduced.
-To keep the magnetic field isochronous up to extraction at 120 MeV, the outer radius was therefore somewhat modified, see left image in Fig.
-3.31.
+To keep the magnetic field isochronous up to extraction at 120 MeV, the outer radius required small modifications, see left image in Fig.
+3.7.
 First, the curvature was modified to follow the curvature of the last orbit.
@@
 Second, the top layer of the pole at its outer radius will be made of permendur.
-This alloy has better magnetic properties at high field, and enhances again the magnetic flux.
+This alloy has better magnetic properties at high field strengths.
+Using permendur will enhance the magnetic flux.
 Finally, the pole height is increased by a slanted ridge at its extremity.
@@
 As reference, the BH curve used for the cyclotron in the CDR is also shown.
-The steel used for the cyclotron yoke and poles is the SNVA6342, also used for example in the C400 cyclotron (see Ref.
-[44]).
-Its BH-curve is shown in fig 3.31.
+The steel used for the cyclotron yoke and poles is the SNVA6342, also used, for example, in the C400 cyclotron (see Ref.
+[bib:c400_iron]).
+Its BH-curve is shown in Fig.3.7.
 Here, we also see the better performance of the permendur alloy.
@@
 In a further study, these last modifications can be made more smooth, as not to have too much of a fluctuation in the dB for the last orbits and thus tunes that do not cross anymore, avoiding possible resonances.
+3.2.2 Risks and Mitigation
+Risk: Manufacturing Imperfections.
+Because no manufacturing process is perfect, we must measure the local magnetic field developed by the accelerator magnet in an iterative process.
+During the mapping iterations, physicists compare measurements results with expected calculated (theoretical) magnetic field map values and propose local magnet iron shape adjustments, typically on pole tips.
+Difficulty of the process is a risk.
+Mitigation: For that reason, pole tips, or at least the edges of the pole tips (boundary valley/hill) must be dismountable to allow re-machining and multiple tuning iterations.
+This mapping process is repeated until the final expected map is obtained, allowing, theoretically, the appropriate beam acceleration and extraction.
+In the case of IsoDAR, we can either perform the entire process underground, or the machine must be dismounted after surface tests and reinstalled underground.
+This will require re-measuring the magnetic field after assembly on site, during a second (shorter) mapping campaign.
+Risk: Mapping Tool. Developing a mapping tool requires care as this tool accuracy will define performance accuracy of the measured magnet itself.
+Mitigation: Many precautions are mandatory concerning calibrations, mechanical stiffness, thermal stability and tracking as well as the use of perfect ”non-magnetic” materials (to avoid measures perturbations and unexpected distortions).
+3.3 Interface with the RFQ
+Figure 3.10: RFQ overview.
+The RFQ will sit on top of the cyclotron along its central axis, as shown in Fig 3.10.
+This places the RFQ outside of the cyclotron’s vacuum.
+Therefore, a vacuum sleeve will be used to ensure the integrity of the cyclotron’s vacuum.
+This sleeve will be fixed on top of the upper external disk and the vacuum sealing will be done radially on the upper internal disk (as seen in figure number 3.12 ).
+The O-rings willhave a cross sectional diameter of 8mm and an inner diameter of 305mm.
+In order to allow for adjustment, the RFQ sleeve allows a rotation of 5mrad around the lower O-ring contact patch (discussed with the manufacturer of the RFQ).
+The RFQ sleeve is as thin as possible to avoid any loss in magnetic field in the central region (hence the O-rings in the Yoke parts).
+Figure 3.11: RFQ zoom.
+Figure 3.12: RFQ cut section 1.
+As shown in Fig.
+3.13, vacuum sealing will occur at the base of the RFQ using a thick O-ring (diameter 10 mm) which will self-seal under the RFQ’s weight ($\approx 3000$ N).
+The dimensions of the vacuum sleeve are listed in Table 3.2.
+In order to allow fine-tuning the cyclotron alignment, a set of adjustment feet, shown in Fig.
+3.11, are fixed to the cyclotron yoke.
+By using four groups of set screws and one vertical adjustment nut, the following alignment modifications can be made.
+  • Rotation: ±$\pm 5$
+  • Tilt: ±$\pm 5$ mrad
+  • Lateral movement: ±$\pm 1$ mm
+In addition to this fine-tuned alignment, a coarse alignment step is envisioned during the initial assembly and commissioning procedure.
+Figure 3.13: RFQ cut section 2.
+Table 3.2: Vacuum Sleeve Dimensions.
+Name Value
+Height 1030.5mm
+Outer diameter 304.55mm
+Inner diameter 296.55mm
+Weight 18kg
+3.4 Central Region
+Figure 3.14: CAD rendering of the spiral inflector and particle trajectories as implemented previously for early tests at Best Cyclotron Systems, Inc.
+[alonso_isodar_2015].
+The positive and negative electrostatic deflectors are shown in blue and green respectively.
+The cyclotron magnetic field in this image is directed vertically upwards.
+Particles enter the spiral inflector via the rectangular grounded collimator (solid gray) and are guided into the cyclotron mid-plane by means of the cyclotron magnetic field and the electrostatic potential between the electrodes.
+The copper housing (transparent gray) isolates the spiral inflector from the RF fields driving the cyclotron.
+From Ref.
+[winklehner_realistic_2017].
+The injection system design is a crucial component that allows us to achieve a high beam current.
+Our design balances injecting the $\mathrm{H}_{2}^{+}$ molecule at high energy against avoiding the use of complex high-voltage platforms.
+Higher energy injection mitigates space charge and reduces geometrical (non-normalized) emittance.
+An energy of 35 keV/amu was selected.
+This in turn determines the target output energy of the RFQ.
+As in commercial cyclotrons, an axial injection system based on a Spiral Inflector (SI) will be used to bend the beam from the axial direction to the median plane.
+We have investigated several refinements to a traditional spiral inflector which will be discussed subsequently, but the concept, as implemented successfully in a previous iteration of this work and tested at Best Cyclotron Systems, Inc., is visualized in Figure 3.14 using trajectories calculated with OPAL [adelmann_opal_2019].
+The exit point of the spiral inflector is also the starting point of acceleration in the Central Region (CR) of the cyclotron.
+First turn acceleration is then achieved by the dee tips extending into recesses in the spiral inflector housing.
+The shapes of the dee tips and the spiral inflector housing have been designed to guide the particles from the spiral inflector exit to the acceleration region while providing the necessary energy gain and beam focusing.
+The central region and the spiral inflector must be designed carefully and together because they have strong interplay.
+This procedure has already undergone one past iteration in the summers of 2013 and 2014 at Best Cyclotron Systems Inc.
+In the following sections, we will discuss improvements to this design which will be incorporated into the preliminary design for the full IsoDAR machine.
+3.4.1 Spiral Inflector Overview & Design Considerations
+The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Figure 3.14).
+The electrostatic potential between these electrodes is able to bend the beam 90 from the axial line to the median plane of the cyclotron.
+The helical trajectory of the beam is determined both by the shape of the electrodes, the electrode potentials, and by the magnetic field of the cyclotron as the beam is bent into the median plane.
+In order to isolate the inflector region from external fields as much as possible, grounded entrance and exit apertures and a grounded copper housing are also included.
+An ideal spiral inflector would do nothing more than rotate the beam by 90, introducing no additional beam divergence or losses and seamlessly merging with initial accelerating in the central region.
+Further, in order to achieve this degree of bending in the small space of the central region, $\mathcal{O}(kV)$ potentials are required, so it is important to minimize or even eliminate any $\mathrm{H}_{2}^{+}$ impact on the electrodes to prevent sparking.
+It is also important to consider the impact of a realistic magnetic field when designing our spiral inflector.
+This is particularly important for the IsoDAR inflector because of its relatively large size.
+Unlike other spiral inflectors that have been constructed, we will require a larger gap and plate size due to the non-negligible space charge and higher rigidity of $\mathrm{H}_{2}^{+}$.
+This larger volume means that variations in the magnetic field along the trajectory are non-negligible.
+Accordingly, all following work uses a full 3D magnetic field map generated using OPERA.
+3.4.2 Designing the Inflector
+Ultimately, an additional iteration of production and empirical testing of the final IsoDAR inflector will be performed.
+To optimize the inflector beforehand, we use a numerical optimization procedure to inform the best theoretical parameters.
+For an inflector in a uniform magnetic field, analytic solutions to particle tracking exist for the central trajectory.
+However when 3D magnetic variations and/or off center trajectories are considered a numerical approach is required.
+An initial estimate based on an analytical solution for a central track serves as a numerical starting point, after which the mesh generating software gmsh [GMSHarticle] is used to produce surface representations which inform subsequent electrostatic analysis.
+Python’s BEMPP API [bempp] is then used to perform boundary analysis.
+In short, this procedure enables a geometric model of our inflector to yield a full 3D electrostatic. Particles can then be tracked with high precision using the Boris method [boris].
+A major strength of this procedure and choice of software is speed, with a full simulation iteration taking only $\mathcal{O}$(hours).
+Prior work [weigelspiral] comparing this process to commercial software such as COMSOL demonstrates that the differences in field calculations are negligible.
+3.4.3 Geometric Refinements
+In most spiral inflectors, the two electrodes would be rectangular prisms if flattened out.
+They remain parallel to each other with locally flat surfaces while spiraling.
+Hereafter we will call this a “Traditional” spiral inflector.
+A traditional inflector can effectively transfer particles, but for off center paths these inflectors can introduce beam divergences in energy and position which were not present in the initial beam.
+We address this by adding several refinements to prior inflector iterations.
+  • Beam focusing between the RFQ and spiral inflector
+  • Introducing V-shape
+  • Introducing plate angling
+  • Modifying the terminal end to a wedge
+The addition of intermittant beam focusing ensures that the $\mathrm{H}_{2}^{+}$ entering the spiral inflector are as near to the central trajectory as possible.
+This is accomplished by the addition of electrostatic quadrupoles in the gap between the RFQ and inflector entrance (see also Fig.
+1.7).
+Using the output of the RFQ simulation [rfqsim] as input to the quadrupoles, geometry and potentials were optimized such that the tightest possible input beam is produced.
+The remaining three modifications: V-shape, plate angling, and the terminal wedge were motivated by work done by Barnard et al.
+[barnard_inflector].
+These all have the effect of introducing quadrupole moments to the field of the inflector itself, which can provide mid-track focusing to minimize beam dispersion in flight.
+The V-shape introduces a concave v-shape to the anode and a convex v-shape to the cathode when viewed along the long axis of the electrode (Fig.3.15(a)).
+Plate angling reorients the plate angle relativce to each other along the course of the inflector.
+In other words the plates are parallel only in the center with a positive angle between the plates initially linearly transitioning to an equal and opposite negative angle by the end (Fig.3.15(b)).
+The terminal gamma wedge cut modifies the end of the anode into a triangular rather than flat profile at the end and shortens the cathode.
+This assists with the fringe fields at the end which can produce vertical spreading.
+(Fig.3.15(c)).
+(a) V-Shape Illustration
+(b) Angling Illustration
+(c) Gamma Wedge Illustration
+Figure 3.15: Illustrations of different focusing concepts. From Ref. [barnard_inflector].
+These modifications and their associated focusing fields interplay, but can be independently varied alongside other parameters like electrode potential, electrode separation, etc.
+We are actively exploring different techniques to handle this high-dimensional optimization problem.
+Monte Carlo sampling is typically well suited to this sort of optimization, but studies in which only one parameter is varied at a time have illustrated that optimization extrema can be relatively narrow, necessitating a significant, if not intractable, amount of computation time.
+As we are at the edge of practical optimization using Monte Carlo, this leads us to consider other potential alternatives.
+Surrogate models using a fully connected neural network capable of learning and interpolating underlying features have been previously explored[EdelenMachingLearning] [KoserMachineLearning] [villarrealMachineLearning].
+Initial iterations have proved promising and have informed the parameters for our current best inflector performance.
+As this work matures, it will be detailed in upcoming publications.
+3.4.4 Characterization of Spiral Inflector
+The result of adding optimized focusing quadrupoles on the beam distribution entering the spiral inflector is illustrated in Fig. 3.16.
+Figure 3.16: Illustration of beam tightness at the entrance of the spiral inflector before and after the addition of focusing quadrupoles between the RFQ and inflector
+The optimized spiral inflector, with $\mathrm{H}_{2}^{+}$ tagged through, is illustrated in Fig.3.17.
+Figure 3.17: Illustrations of the optimized spiral inflector.
+Note the v-shape to the electrodes as well as the non-parallel angle seen in the bottom view, and the angled wedge visible in the top view.
+The efficiency of this modified inflector is significantly improved.
+Efficiency must be thought of from two fronts.
+On the one hand we wish to transmit as much current from the RFQ into the cyclotron as possible.Particles which are not successfully transmitted must terminate somewhere.
+Termination on grounded apertures is acceptable while termination on the HV electrodes will result in sparking and can have significantly impact on running.
+It is therefore also vital to specifically consider the fraction of ions that terminate on an electrode.
+We thus consider two similar cases.
+Both have identical optimized spiral inflectors.
+The difference is whether or not a grounded collimator is placed at the entrance to further narrow the entering beam beyond what the quadrupoles have provided.
+This collimator is a circular aperture which, for this study, was chosen to be as large as possible while guaranteeing that less than one percent of $\mathrm{H}_{2}^{+}$ impact the electrodes.
+The resulting efficiencies and $\mathrm{H}_{2}^{+}$-electrode collisions are given in Table 3.3.
+Table 3.3: Comparison of efficiency and ion-electrode impact for the optimized vs traditional spiral inflector.
+Efficiency Electrode Impact %
+Traditional Inflector 74% 17%
+Optimized Inflector 97% 1%
+Traditional Inflector (Coll.) 39% $<1$%
+Optimized Inflector (Coll.) 93% $<1$%
+We also need to consider the impact of the new inflector on the quality of the beam which is successfully transmitted.
+These changes are characterized in Table 3.4.
+Detailed histograms showing the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ at the entrance and exit of the inflector are shown in Fig.
+3.18.
+Table 3.4: Comparison of transmitted beam parameters at the inflector exit.
+$\sigma_{x}$ [mm] $\sigma_{y}$ [mm] $\sigma_{z}$ [mm] $\sigma_{r}$ [mm] $\sigma_{KE}$ [keV]
+Traditional Inflector 3.0 3.9 10.1 2.3 4.0
+Optimized Inflector 4.9 2.7 6.4 2.4 3.1
+Figure 3.18: Illustration of the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ prior to the inflector entrance vs at the inflector exit.
+3.4.5 Orbit Matching and Beam Centering
+Figure 3.19: Beam tracking result from the AIMA central region study.
+Left beam centering and acceleration on top of a view of the electrodes.
+Colors correspond to prticle phase.
+Right: 1 cm turn separation in the fourth turn with a single collimator.
+From [aima:central].
+Once the beam emerges from the spiral inflector, it must be properly matched in terms of orbit radius, radial momentum, and phase.
+A specially designed shape of the inner ends of the RF cavities performs this duty.
+The angles and widths of the accelerating gaps of each cavity are modified in an optimization procedure during the first two turns to place the beam onto a static equilibrium orbit appropriate for the respective energy.
+The center of a full integrated orbit after these two turns should then align well with the cyclotron center.
+The spiral inflector end can be tilted (often this parameter is called $k^{\prime}$) to aid in the matching and thus becomes part of the optimization procedure.
+A preliminary design study of the IsoDAR central region was performed by the company AIMA in France [aima:central].
+The AIMA central region model is shown in Fig.
+3.19.
+This central region was designed with a slightly different magnetic pole shape and with its own matched spiral inflector, using the ouput beam from an early RFQ beam tracking and single electrostatic quadrupole transverse refocusing element.
+It demonstrated that the RFQ-generated beam could be injected and vertically focused.
+By using a single collimator (removing $\approx 42$% of the beam), an edge-to-edge turn separation of 1 cm was achieved in the fourth turn (approximately 1 MeV/amu).
+The RMS beam size and energy spread was significantly worse than what we see in our optimized inflector.
+Furthermore, the AIMA study did not consider space charge, which is fundamental to establishing vortex motion (see below).
+Somewhat counterintuitively, including space-charge will improve the results further; as will our new optimized spiral inflector.
+In the next step, we will modify the central region and optimized spiral inflector to match well and track particles through the finalized geometry with OPAL using the PIC method to demonstrate a stable vortex.
+For this preliminary design, we use a Gaussian beam with size and emittance informed by the AIMA study for further simulations of acceleration to 60 MeV/amu.
+Figure 3.20: Reduction of beam halo by collimators
+3.4.6 Low Energy Collimation
+Following injection, there is significant beam spread in the first few turns of the cyclotron that we need to manage.
+We do this by placing a series of collimators around the beam core in the early turns.
+We must do all of our collimation in the first 7 turns of the primary acceleration, corresponding approximately to a 2 MeV/amu limit, to avoid activating our collimators.
+The placement of the collimators is shown in Figure  3.21 The reduction in RMS spread is shown in Fig  3.41.
+In total, we have 15 collimators that remove 36.1% of the beam.
+The improvement in beam shape is shown in  3.20.
+With time, the halo slowly redevelops but to a lesser extent than it would without collimation.
+Figure 3.21: Placement of 15 collimators in first 7 turns
+It is necessary to make the beam as circular as possible following the low energy collimators so that the beam is well-matched to “vortex motion.” Vortex motion stabilizes the beam and significantly reduces beam growth over time.
+It arises from the interaction between the self-field of the beam and the external magnetic field of the cyclotron.
+Vortex motion is discussed in depth in Ref [baumgarten:vortex2].
+3.4.7 Risks and Mitigation
+Risk: High voltage on the spiral inflector electrodes.
+The high voltage on the spiral inflector electrodes can lead to arcing from one electrode to the other or to ground, creating unstable injection conditions.
+Mitigation: During the BCS tests we were able to run the spiral inflector electrodes as high as ±13.5 kV without beam using conventional polishing and cleaning methods for the electrodes.
+Electro-polishing and conditioning in an inert gas environment will improve this even more.
+Beam striking the electrodes could potentially induce arcs through electron emission.
+This can be avoided by presenting a well-defined beam to the spiral inflector to assure close to 100% transmission.
+We have and will continue to emphasize minimizing or completely eliminating ion impacts in simulo with these tools.
+Finally, we will set up machine protection algorithms that can break a spark and restart acceleration in a controlled manner to minimize downtimes due to sparking.
+3.5 Vacuum Chamber
+The vacuum enclosure comprises five parts: the upper and lower internal disk and 3 vacuum chambers.
+Fig 3.22 shows the vacuum enclosure in detail and the dimensions are in Table 3.5.
+The vacuum sealing is done using 2 metallic O-rings per interface (e.g., from Omniseal [PrecisionPolymerMetal]) except for the radial interface where Polymer O-rings could be used.
+The interface for the accessories will be done using a sleeve design and radial O-ring, described in Fig.
+3.22.
+Figure 3.22: Cyclotron vacuum chamber details
+Name Mass (Metric Ton) Dimensions (mm)
+Upper and lower internal disk 36 Ø4620*463
+Upper vacuum chamber 2.3 Ø4452*483
+Lower vacuum chamber - median plane 2.3 Ø4452*483
+Lower vacuum chamber - internal disk 1.8 Ø4386*380
+Table 3.5: Vacuum part properties
+3.5.1 Risks and Mitigation
+Risk: Achieving the required vacuum level in the presence of the beam.
+The calculated required vacuum is $5\cdot 10^{-8}$ mbar.
+Before the beam is turned on, the pumping plan provided will reach this with little difficulty, assuming there are no vacuum leaks.
+However, introduction of beam, with gas-desorption from surfaces hit by stray beam particles, could raise the pressure above the acceptable level.
+Mitigation: Beam losses from other sources will be minimized.
+In addition, options for further surface cleaning and treatment, as well as the possible addition of more pumping capacity, should be developed.
+3.6 Coil configuration
+The CDR anticipated a slightly shorter coil than we now anticipate using.
+Specifically, the coil height has been increased from 200 to 240 mm.
+Its radial width remains at 250 mm.
+The total current used in the calculations is 180,850 amp-turns in each of the half coils.
+The apparent current density is thus 3.01 A/mm2, i.e.
+$5\%$ lower than the 3.17 A/mm2 foreseen in the CDR.
+For now, no extraction channels are anticipated in the return yoke.
+If they are added, an increase in main coil current will be required
+Figure 3.23: Left: Schematic view of a cross-section of the main coil, with six pancakes per half coil.
+Right: Possible configuration of the winding, using a square conductor with a round cooling channel in its center.
+In Fig.
+3.23 a schematic view of the coil is given.
+Also, a possible configuration of winding and cooling is calculated.
+Although this configuration does not exactly sum to a cross-section of 240 by 250 mm2, it shows that a cooling configuration is feasible.
+Working in pancakes gives the possibility of manufacturing the pancakes outside the cave and stacking them inside.
+The increase in height from 200 to 240 mm is the equivalent of adding an extra pancake.
+3.7 Cyclotron RF System
+The RF system for the IsoDAR cyclotron produces the primary acceleration of the ion beam and can be divided into the following subsystems:
+  • The RF cavities which accelerate the beam and provide large inter-turn separation.
+The cavities are wedge-shaped and are located in the magnet valleys.
+There are 4 such cavities.
+The inner part is often called a dee whereas the outer part is called a dummy-dee.
+  • The RF tuning system that will compensate for the thermal drift of the cavities and maintain a stable resonating frequency.
+  • The RF amplifiers that provide the necessary power to create the required RF electric field inside the cavity.
+  • The RF couplers that inject the RF power from the amplifiers into the cavities
+  • The Low Level RF system (LLRF) that controls and regulates the RF amplitude on the Dee and drives the tuning system.
+Table 3.6: Details of the RF design.
+RF System Component Design Value
+Resonance Frequency 32.8 MHz
+Dee voltage in the central region $65$ kV
+Dee voltage at extraction radius $230$ kV
+Dee radial extension 2 m
+Acceleration gap angle 42
+Cavity height 1812 mm
+Number of Dee stems 4 per Dee (2 up, 2 down)
+Number of Dee’s 4
+Acceleration harmonic 4th
+Power dissipated per cavity 113 kW
+Cavity Q-factor 9620
+We consider these in the Sections below. The numerical parameters of the RF system are listed in Table 3.6.
+3.8 RF Cavities (Dees, Stems and Liners)
+The RF cavities of the cyclotron have been designed using CST microwave studio [cst:microwave_studio] in order to define the shape of the Dee and the Dee stems.
+We optimized our designs to reach the right resonance frequency, satisfy the Dee voltage Law, and minimize power dissipation.
+Ultimately, this improves the quality factor (Q-factor).
+Figure 3.24: Section view of the RF cavity.
+Figure 3.25: Outer view of the RF cavity.
+The Dees and cavity walls are all made of oxygen-free high-conductivity (OFHC) copper, which has good electrical conductivity and good solderability.
+The Dees are fabricated out of a 20 mm thick copper plate and are made rigid by a supporting arm.
+The four round stems support the Dee and provide the right inductance in order to produce the required resonance frequency.
+Their size and location produce the desired Dee voltage vs.
+radius profile.
+Each Dee will be equipped with vacuum pumping panels to efficiently pump the median plane.
+One Dee is equipped with one electrostatic deflector for extraction.
+This fact has not been taken into account in detail in the RF simulations to date, but some room for housing the deflector has been provisioned at the back of the Dee.
+Figures 3.24 and 3.25 present the main elements composing the RF cavities.
+3.8.1 Integration of preliminary central region in RF Cavities
+In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA [aima:central] into the RF cavity modelling, several technical choices were made to ensure the effective implementation of this innovative configuration:
+  • The central region design was trimmed at 150 mm from centre.
+  • Both the Dee and counter dee original central region was removed by trimming the cavities at 250 mm from centre.
+  • Sections were cut straight, and a linear transition (loft) technique was used to achieve the connections between the trimmed components.
+The figure 3.26 shows the final integration of the central region.
+The lofted transitions can be better seen from top at figure 3.27.
+They leads to an acceptable linear vertical gap opening from 26 mm at a radius of 150 mm, to 50 mm at a radius of 250 mm.
+The section view of the model, presented at figure 3.28, shows the radial evolution of the vertical gap.
+In addition, adequate clearance provisions were incorporated, specifically to fit the inflector and optics coil.
+Symmetry with regard to the median plane symmetry was maintained.
+Power bridges between dees in the central regions were removed since we recommend to power the 4 cavities independently.
+Figure 3.26: dee at central region close view
+Figure 3.27: dee lofted transition top view
+Figure 3.28: vertical gap evolution close view
+3.8.2 RFQ Integration in RF Cavities
+Some clearance provision has been set for the integration of the RFQ, as seen in figure 3.29
+  • Diameter 280 mm
+  • Down to 35 cm of median plane
+While the RFQ will only be in the upper location, we have decided to keep median plane symmetry, for computation, manufacturing, and performances purposes.
+We have observed that the RFQ clearance has a low impact on cavity frequency and performances, and the output of the RFQ could be lowered more without impacting the cavity behavior.
+Figure 3.29: Clearance for RFQ
+3.8.3 RFQ getter pumps in RF Cavities
+Incorporating getter pumps into the Dees allows us to improve our vacuum quality.
+By integrating three housing for SAES UHV1400 getter pumps [ModulesUHV] per dee, we aim to ensure highly efficient and reliable vacuum performance within the Cyclotron system.
+The use of six of those getter pumps per cavity facilitates the effective removal of residual gases, since achieving and maintaining ultra-high vacuum conditions essential for optimal particle acceleration.
+The figure 3.30 shows acceptable locations for three getter pump housings.
+Furthermore, our preliminary assessments suggest that it may be feasible to scale up the configuration to include up to two sets of five getter pumps per cavity.
+The very low electrical field of the locations of the housings will allow the getter pumps to function safely.
+Figure 3.30: Provisional getter housings
+This early integration of the housings of the getters in the dees is mandatory to assess their impact on the cavity frequency, and on surfaces currents.
+However, to finalize this integration, comprehensive vacuum computations will be needed.
+These computations will enable us to determine the ideal amount of getter pumps, ensuring a properly engineered vacuum system that underpins the overall success and efficiency of the Cyclotron.
+3.8.4 RF Cavities design optimization
+3.8.4.1 Performance requirement target
+We have optimized the cavity geometry to ensure the following requirements:
+  • RF frequency of accelerating mode: 32.8 MHz
+  • Reference accelerating voltages: 65 kV at 80 mm (capture), 230 kV at 2000 mm (extraction)
+The initial target value for the outer radius voltage was initially set at 250 kV, but according to our expertise, such high voltage will be very difficult to stand in cyclotron accelerator gaps of this size.
+It has been decided to reduce the target to 230 kV which remains a challenging value.
+3.8.4.2 Parametric optimization of the design
+We have intensively used parameters while building the CST model, to allow fast design exploration.
+We have retained four parameters to play with, for optimizing the cavity behavior.
+The parameters, shown at figure 3.31, are defined as a deviation from the early cavity design values.
+  • A: 1st pillar radial position
+  • B: 1st pillar diameter
+  • C: Deflector wall distance
+  • D: Deflector height
+Figure 3.31: design optimization parameters
+We did some initial exploration of the parameter space, but we did the final optimization using the built-in CST optimizer.
+Table 3.7: RF cavities optimization goals definition and achieved values.
+Goal description Target Final Value
+Ratio between voltage at 80 mm and 2000mm 0.28 0.283
+Mode1 frequency [MHz] 32.8 32.805
+We selected the trust region algorithm and have defined scaled goals that would ensure convergence toward the required frequency and voltages.
+The trust region algorithm converged in a few tens of iterations.
+The definition of the optimization goals, and their target and reached values are contained in table 3.7.
+The table 3.8 shows the optimal values reached by the optimizer.
+Table 3.8: RF cavities optimized design parameters.
+Design parameter Value [mm]
+First pillar radialposition (A) 444.6
+First pillar diameter (B) 117.7
+Distance between dee and cavity back wall (C) 42
+Total cavity height (D) 1412
+3.8.5 RF cavity computation results
+The RF cavity computation employing the eigen mode solver on the optimized geometry yielded the following performance and characteristics of the cyclotron system:
+  • Frequency: The RF cavity operates at a frequency of 32.805 MHz, ensuring compatibility with the designed acceleration, and the RFQ integration,
+  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.283, indicating that computed voltages at center would be 65.2 kV for a chosen voltage of 230 kV at the outer radius (2m),
+  • CST computed quality factor (Q): 10586,
+  • Estimated quality factor (Q): 7281,
+  • Stored Energy (for max RF of 230 kV at 2 m): 4.06 Joules
+  • CST Computed Losses: The computed losses per cavity using CST simulations are 79 kW,
+  • Estimated Power Dissipation: By experience, senior RF designers at IBA state that the CST eigenmode simulations underestimate the actual losses by up to 30%.
+If we increase the CST computed losses (79 kW) by this percentage, the actual estimated power dissipation per cavity becomes 113 kW.
+  • Estimated Power Requirement: To effectively drive the cyclotron, each cavity will require an estimated power input of at least 263 kW, which include power dissipation and beam loading.
+  • Maximal Surface Current: The rms surface current reaches a maximum of 17052.6 A/m in some reduced locations.
+Which leads to a maximum surface power of 43.5 W/cm².
+This maximal surface power is localized on the dee edges.
+Careful attention will be required for thermal management in these regions.
+For the rest of the surfaces, the thermal dissipation is well below 10 W/cm².
+See figure 3.36 and 3.37 While matching the requirement, these computation results provide valuable information for next steps of the Cyclotron’s RF cavity design including consideration in managing thermal effects.
+Some illustrations of the electrical and magnetic fields computed by CST are given for reference at figures 3.32, 3.33, 3.34 and 3.35.
+Figure 3.32: E field Vertical cross section
+Figure 3.33: E-field Horizontal cross section
+Figure 3.34: H-field horizontal cross section
+Figure 3.35: H-field vertical cross section
+Figure 3.36: Surface currents
+3.8.5.1 Detailed power analysis
+We will now detail the post treatment of the CST results that lead to the results presented in Sec.
+3.8.5.
+The CST eigenmode solver results are computed for a given reference energy in the system (about 1J).
+They must then be scaled accordingly, to reach the expected voltage at a reference location.
+In our case, the reference location, is the orbit at 2m, and the desired voltage is 230 kV.
+Figure 3.37: Surface current generating more than 10 W/cm2
+The ratio of the desired voltage to the reference voltage at 2m is 2.013.
+We can scale the currents and voltages using this factor.
+We can also scale the energies, power and losses using this factor squared (formulae 3.1, 3.2 and 3.3).
+We thus have:
+$E_{stored}=E_{stored,ref}\frac{V_{target}^{2}}{V_{ref}^{2}}$ (3.1)
+$P_{loss}=P_{loss,ref}{V_{target}^{2}\over V_{ref}^{2}}$ (3.2)
+$I{{}_{s}}=I_{s,ref}{V_{target}\over V_{ref}}$ (3.3)
+Table 3.9: Optimized cavities, functional characteristics.
+CST reference Scaled value
+Stored energy [J] 1 4.06
+Voltage at on the 2m orbit kV] 114.2 230
+Voltage at center [kV] 32.4 65.2
+Maximal voltage at outer edge of the dee [kV] 124.6 250.9
+Computed losses [kW] 19.48 78.9
+Maximal surface current [A/m] 8470 17052.6
+[…]
+$R_{s}={1\over\sigma.\delta}$ (3.4)
+$\delta=\sqrt{2\over{\sigma.\mu.\omega}}$ (3.5)
+$P_{s}=R_{s}I_{s}^{2}$ (3.6)
+Table 3.10: Cavity properties for computing surface dissipation
+Property Value
+Copper magnetic permeability $\mu$ $0.99991\times 4\pi\times 10^{-7}$ H/m
+Copper conductivity $\sigma$ $5.8\times 10^{7}$ S/m
+Angular Frequency $\omega$ $2\pi\times 32.805\times 10^{6}$ rad/s
+Skin Depth $\delta$ $11.5$ µm
+3.8.5.2 Radial voltage dependency
+The voltages at different radii have been calculated by integrating the electrical field along circular arcs crossing the acceleration gaps, and reference curves crossing central gap and the dee outer edge gap.
+The transit time factor has not been considered.
+The computed voltages have then been scaled to match a target voltage of 230 kV at 2 m.
+The resulting absolute and normalized voltage radial dependency are given in figure 3.38 and table 3.11
+Figure 3.38: Dee voltage radial dependency
+Table 3.11: RF voltage radial dependency.
+Radius [mm] Voltage [kV]
+Capture (80) 65.17
+100 64.07
+200 63.95
+300 64.60
+400 63.76
+500 66.37
+600 71.30
+700 77.60
+800 85.50
+900 95.67
+1000 107.09
+1100 119.72
+1200 133.04
+1300 147.22
+1400 160.77
+1500 174.31
+1600 187.84
+1700 199.73
+1800 211.94
+1900 222.49
+Extraction (2000) 230.00
+back of dee 250.91
+3.8.6 RF Tuning System
+The tuning system is a movable plate that corrects the resonance frequency drift mainly caused by RF heating.
+The system here is based on four separated tuning plates (one per cavity) and located in the median plane.
+These plates are located at the back of the Dee and are controlled via a ball screw that is driven by a motor.
+This system is similar to that of many commercial cyclotrons.
+3.8.7 RF Amplifiers
+[…]
+We plan to use vacuum tubes (triodes or tetrodes) to produce the RF power.
+We have considered using LDMOS amplifiers given the recent improvements in power density, reliability, and cost.
+However, we believe that vacuum tubes will perform better in the harsh environment and will be able to better accommodate the rapid variations in load due to the beam temporal patterns.
+3.8.8 RF Lines and Couplers
+The RF power is fed to the cavities through EIA 6 1/8” solid copper coax line. There are basically two ways to couple RF to the cavities:
+  • Capacitively in the median plane
+  • Inductively in the bottom (or top) of the cavity liner.
+Both of these systems are capable of providing the necessary 300 kW power to each cavity.
+Capacitive coupling is less prone to multipacting because the coupling antenna could be easily DC biased.
+But, our expertise in high power coupler design advocates for an inductive solution.
+The design of inductive coupler is easier to water-cool, and there is a wide area where H-field is high enough for efficient coupling behind the second pillar of the cavity.
+We have initiated a preliminary design computation of such an inductance coupler, feeding the RF power from a EIA 6”1/8 port.
+The suggested design is a loop made out of a 20x60 mm rectangular section bar, offering room for water-cooling.
+See figure 3.39.
+The adaptation tuning will be performed by rotation of the loop around the vertical axis.
+The design has been optimized with a default angle of 45° to the cavity axis, in order to allow a maximal tuning range.
+The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 235 mm, which provides a good matching (figure 3.40) and a $S_{1,1}$ computed scattering parameter inferior to -60 dB.
+Figure 3.39: RF loop coupler design
+Figure 3.40: Coupler port impedance
+3.8.9 LLRF Electronics
+The LLRF (Low Level RF) controls the RF amplitude and phase of the signal sent to the amplifiers.
+It also controls the tuning mechanism in order to maintain the resonant frequencies in the cavities.
+The amplitude and phase control is based on feed-back signals coming from the cavities and amplifiers.
+In order to avoid multipactoring, the LLRF manages also a pulsing mode that provides for an easy start-up and reduces the power dissipated in the amplifiers.
+One LLRF system per cavity is planned.
+An additional electronic element is needed to keep a stable phase difference between the different cavities.
+3.8.10 Risks and Mitigation
+Risk: Design and Manufacturing Process.
+The Q factor and eigenfrequency depend on the manufacturing process of the RF Cavity, e.g., surface quality and material purity.
+Furthermore, there is a risk of multipactoring during run-time as electrons can be accelerated back an forth in the RF field.
+Mitigation: A 30% margin was built into the RF power estimates, based on operational experience by IBA.
+An appropriate tuning range is built into the cavity design.
+A full scale RF cavity will be built and tested before the start of the construction of the other three cavities.
+Options of cavity coating exist in case of multipactoring, e.g., Aquadag.
+3.9 Cyclotron Particle Simulations
+It is necessary to perform an in-depth particle simulation of the primary acceleration of the beam to ensure that the beam reaches energy, maintains quality, and is possible to extract.
+We do this using OPAL  [adelmann_opal_2019] which is a particle accelerator simulation suite developed at the Paul Scherrer Institute that accounts for space-charge effects.
+It is based on particle-in-cell (PIC) calculations that track a ’bunch’ of particles through time-varying magnetic and electric fields.
+The temporal integration within OPAL is carried out using a fourth-order Runge-Kutta method.There is a mode within OPAL, OPAL-cycl, which is designed to simulate cyclotrons.
+This makes it well-suited for the primary acceleration phase of the IsoDAR cyclotron.
+Conveniently, OPAL-cycl has built-in objects called PROBEs, that can provide beam diagnostics, and and COLLIMATORs, which can provide halo reduction.
+3.9.1 Fieldmaps
+A critical aspect of these simulations is that they accurately simulate the electric and magnetic fields within the cyclotron.
+We performed these simulations using a 1D RF model derived from the voltages provided in Table 3.11.
+The radially increasing voltage leads to a roughly constant turn separation, which is important for ease of extraction.
+We conducted these simulations using a 2D median plane magnetic field that is derived from a physically realistic cyclotron model produced by IBA.
+Of note, this magnetic field map does not include a $\nu_{r}=1$ resonance, which would increase the inter-turn separation in the last few turns at the price of reduced beam quality.
+This magnetic field is significantly more realistic than our previous work because the cyclotron model allots space for the injection components.
+3.9.2 Probes
+Probes are a built-in OPAL object that records whenever a particle passes through the region defined by the probe.
+These particles are simply recorded, and the trajectories of the particles are not affected by their intersection with the probe.
+This is helpful for determining the “extractability” of the beam.
+The probe collects information about the position and the momenta of the passing particles.
+By placing a probe at an azimuth of 25 degrees, the angle at which our electrostatic channel starts, we are able to determine the overlap between sequential turns of the beam path.
+This is necessary because we want to minimize this overlap for ease of extraction.
+The overlap between beams provides a lower bound for the power that will be deposited on the electrostatic septum, so we aim to minimize this value.
+The data from the probe is displayed in Fig 3.42.
+3.9.3 Simulation Results
+Using the field maps as discussed earlier and the set of 15 collimators as described in the previous sections.
+We are able to smoothly accelerate the beam up to 60 MeV/amu as shown in Fig 3.43.
+Additionally, there is minimal growth in beam size in the collimated beam as shown in Fig 3.41.
+Additionally, a probe at an azimuth of 25 degrees sets a lower limit for extraction at 120W as seen in Fig 3.42.
+Figure 3.41: The collimated beam had a much reduced radial RMS spread.
+Figure 3.42: There are two potential extraction points
+Figure 3.43: There is a smooth acceleration of the beam
+3.9.4 Electrostatic Extraction Channels
+The first step in our extraction is the construction of a radially directed electric field that peels the last turn away from the second to last turn.
+This is done by placing a series of thin (0.2 mm) grounded electrodes between the last two turns and a larger puller electrode that is at high negative potential outside of the last turn.
+This generates a well-shielded electric field in the region between the puller and septum.
+A rendering of the electrostatic septum is shown in Fig 3.44.There are two such electrostatic channels that increase separation between the last two turns.
+Figure 3.44: CAD rendering of electrostatic septum. From Ref. [winklehner_new_2022]
+We determine the exact location of each element of the extraction channel iteratively.
+The first step is to run a Python script on a pre-existing OPAL simulation of the last three turns that determines the initial positions of the septum and the puller.
+We choose to place the septum evenly between the last two turns.
+A .txt file is generated that allows for a model of the septum to be built in Autodesk Inventor.
+With this geometry, we use COMSOL to determine the electric field and add it to the pre-existing RF field of the cyclotron in OPAL.
+We model the physical dimensions of the septum as COLLIMATOR objects in OPAL.
+This process is repeated twice at which point further iteration does not result in any improvements in ESC performance.
+Following this we were able to reduce our losses on the first septum to 205 W which is 2.5% greater than our 200 W goal and we lose 21 W on the second septum.
+In the future we plan to reintroduce a resonance into the magnetic field which will increase the inter-turn separation at extraction which will greatly reduce this number.
+The end result of the septum optimization process is shown in Fig 3.45.
+Figure 3.45: This is the tracking of 100 particles as they pass through the electrostatic channels (shown in green). It also shows the location of our magnetic extraction channel in grey.
+3.9.5 Magnetic Extraction Channels
+After the electrostatic extraction channels, the beam will rapidly de-focus as it passes through the radially decreasing magnetic field.
+In order to avoid this issue, we must include focusing elements that can counteract the effects of the fringe fields of the cyclotron.
+Due to space constraints, it is not possible to have an active element such as a quadrupole within the cyclotron.
+Instead, we will use the standard strategy of machining and placing iron bars along the extracted beam path.
+These bars allow us to locally reshape the magnetic field and are very compact.
+We model our magnetic extraction off of the MSU cyclotron [MSU_Cyclotron] which generates a strong radially increasing magnetic field that serves to refocus the beam after the ESC.
+By using the gradients that are produced at MSU as a reference point, we conservatively set a gradient of 1.5 kG/cm which is much less than the 5 kG/cm that they were able to achieve.
+We directly modified our magnetic field to contain these elements.
+It is visible in Fig 3.46 that a fairly week magnetic channel is sufficient to maintain beam quality.
+We do not anticipate difficulties with these channels because the electrostatic channels generate the majority of the turn separation.
+Figure 3.46: A 1.5 kG/cm gradient after the electrostatic channel dramatically reduces radial beam spread
+3.9.6 Risks and Mitigation
+Risk: Inadequate turn separation.
+If beam bunches are not radially separated by a sufficient amount, unacceptable beam loss will occur on the thin septum rendering its lifetime to be inadequately short.
+Mitigation: Understanding beam dynamics sufficiently well that this does not happen, and having adequate beam diagnostics to be able to tune the beam for maximum turn separation.
+Further, we could place a narrow carbon foil in front of teh septum, dissociating the few $\mathrm{H}_{2}^{+}$ ions that would otherwise hit the septum.
+The resulting protons have lower magnetic rigidity and bend away from the septum before hitting it.
+They follow a specific trajectory that can guide them outside the cyclotron to be dumped in a controlled way.
+Risk: Placement of the narrow stripping foil.
+It is not obvious where the optimum location is for the thin stripper.
+It must be such that stripped protons do not strike the septum, but also that halo particles do not miss the stripper and strike the septum.
+Mitigation: Adequate beam dynamics simulations, including tracing of particle orbits.
 Chapter 4 Installation in the Yemilab Setting
-The most likely deployment scenario is at the Yemilab underground laboratory, 1 km below the surface adjacent to the Handuk iron mine in the Republic of Korea.
+The most likely deployment location for IsoDAR is at the Yemilab underground laboratory, 1 km below the surface adjacent to the Handuk iron mine in the Republic of Korea.
 Fig.
-4.1 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 meters at a 15% slope from a surface access point into the mine, or by an elevator conveyance in a new 600 meter vertical shaft.
-Unfortunately, the elevator is suitable only for personnel access, all heavy and bulky equipment must be brought down through the ramp.
+4.1 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 meters at a 15% slope from a surface access point into the mine, or by an elevator which descends down a 600 meter vertical shaft.
+Unfortunately, the elevator is suitable only for personnel access.
+All heavy and bulky equipment must be brought down via the ramp.
 Fig.
 4.2 details the area where the IsoDAR experiment would be installed, with the target only a few meters from the planned 3 kiloton LSC (Liquid Scintillator Counter).
-Site-specific questions will be addressed in Volume III of this PDR and the following section aims at a general description of deployment strategies in an underground setting, pertaining only to the cyclotron.
-In this document we use Yemilab as the example, noting that most of the issues discussed will arise in an equivalent manner at other installation sites.
+The following section provides a general description of deployment strategies in an underground setting, pertaining only to the cyclotron.
+In this document we use Yemilab as the example, noting that most of the issues discussed will arise in an equivalent manner at any alternative installation sites.
+A similar approach will be taken for the target system in Volume II of this PDR.
 Figure 4.1: (a) Schematic of Handuk/Yemilab area, (b) Low resolution laser scan of Handuk mine and Yemilab access ramps, green area denotes high resolution scans of IsoDAR deployment area, shown in more detail in the next figure.
+From Ref.
+[kimYemilabNewUnderground2024].
 Figure 4.2: Layout of Yemilab, showing (in green) the area where IsoDAR will be installed. The cyclotron will be assembled in the dedicated cyclotron vault as shown.
@@
 Figures 4.8 and 4.9), the resolution of the laser scans of the access ramp areas are not of sufficient resolution to provide enough detailed information about the location or size of the possible constrictions.
-Improved scans must be performed as part of final installation planning, and will be included in volume III which will be site-specific to Yemilab.
+Improved scans must be performed as part of final installation planning in the Technical Design Report phase (TDR).
 As seen in the previous chapter, the iron pieces that make up the cyclotron magnet, although heavy, are not in themselves large, so the pinch points are not likely to be a problem for bringing the magnet steel to the installation site.
@@
 4.3 for bringing the magnet and vacuum liner pieces from the surface.
-These electrically-powered vehicles are flexible in their size, the sets of 4 tires are independently steerable, so the vehicle is extremely maneuverable, and can easily accommodate the heaviest piece of magnet steel [45].
-Figure 4.3: A typical SPMT vehicle, extremely versatile and capable of extremely heavy loads over irregular terrain
+These electrically-powered vehicles are flexible in their size, the sets of 4 tires are independently steerable, so the vehicle is extremely maneuverable, and can easily accommodate the heaviest piece of magnet steel [SPMT].
+Figure 4.3: A typical SPMT vehicle, a versatile way to transport heavy loads over irregular terrain.
 The only challenge will be the magnet coils.
@@
 In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections.
-This has already been studied in Chapter 3 of the CDR document [46, 2].
+This has already been studied in Chapter 3 of the CDR document [alonso_isodaryemilab_2021, alonso_isodaryemilab_2022].
 The heavy winding machine and fixtures would need to be transported, but could be broken down into parts that will fit through the ramps, then assembled on site.
@@
 There is precedence for sectioned coils.
-For the large TRIUMF cyclotron [47], coils are separated into six different segments as the outer diameter of the coils is almost 20 meters, making the transport of fully assembled coils essentially impossible over conventional roadways.
+For the large TRIUMF cyclotron [TRIUMF], coils are separated into six different segments as the outer diameter of the coils is almost 20 meters, making the transport of fully assembled coils essentially impossible over conventional roadways.
 The inconvenience for a segmented coil, is that the ends of every turn of every section must be spliced, substantially increasing the cost and decreasing the efficiency of operation, as the coil must have as few turns as possible with very high current flowing through each turn.
-Figure 4.4: Isometric view of the coil package, on a specially-designed cart, transporting the coil through the access ramp to the Yemilab campus.
 Figure 4.4: Isometric view of the coil package, on a specially-designed cart, transporting the coil through the access ramp to the Yemilab campus.
@@
 Conversations with a crane manufacturer, Industrias Electromecánicas GH, S.A.
-[48] gave us assurances that a 40-ton capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig.
+[GHcranes] gave us assurances that a 40-ton capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig.
 4.6.
 It could be configured to fit the dimensions of the cavern, and broken into modular parts that could be transported to the site.
-Figure 4.6: Examples of GH gantry cranes, capacities: foreground 20 tons, background 260 tons
-Figure 4.7: Cyclotron and crane on cavern floor.
+Figure 4.6: Examples of GH gantry cranes, capacities: foreground 20 tons, background 260 tons. From Ref. [GHcranes].
 Figure 4.7: Cyclotron and crane on cavern floor.
@@
 In addition, the large crane can be reconfigured and reassembled in the target hall, and used for assembly of the MEBT, target and shielding, and for subsequent maintenance of the target systems.
-Figure 4.9: Cycloton open and crane with cavern mesh - cut-view.
+Figure 4.9: Cyclotron open and crane with cavern mesh - cut-view.
+4.2.1 Risks and Mitigations
+Risk: Coil Size.
+The cyclotron coil will likely not fit through all of the sections of the main access ramp at Yemilab in horizontal orientation.
+A special mounting fixture capable of holding the coil packages at an angle should be able to bring these coils through diagonally.
+However, there is some uncertainty about possible choke points in the main access ramp to the Yemilab facility.
+Mitigation: In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections, as discussed above.
+Both come with drawbacks or additional cost, but are manageable.
 Chapter 5 Conclusion
 5.1 Summary
-In this version 0.9 of the document, we presented the preliminary design of the IsoDAR compact cyclotron with RFQ injection, including ion source and low energy beam transport.
-The presented design study shows the feasibility of producing a high current (10 mA of protons through stripping of a 5 mA $\mathrm{H}_{2}^{+}$ beam) in a compact cyclotron.
+In this version 2.0 of the document, we presented the preliminary design of the IsoDAR compact cyclotron with RFQ direct injection, including ion source, low energy beam transport, electrostatic extraction channels and some thoughts on typical magnetic extraction channels.
+This version presents significant updates over the previously published CDR and also version 0.9 of this PDR.
+Most notably, we included first information about the upgraded MIST-2 ion source, a much optimized spiral inflector, a realistic cyclotron magnet and RF cavity design and particle-in-cell simulations thereof.
+Over version 0.9 of this document, we added detailed subsections on the simulations and subsections on risks associated with the different components as well as mitigation strategies.
+The presented design study shows the feasibility of producing a high CW $\mathrm{H}_{2}^{+}$ beam current (10 mA of protons through stripping of a 5 mA $\mathrm{H}_{2}^{+}$ beam) in a compact cyclotron.
 It further shows a convincing strategy to segment the cyclotron yoke for transportation into the underground environment to be assembled in situ at the final location.
-RF calculations show manageable power and cooling, the required radial voltage distribution in the accelerating gaps, and adequate room for vacuum getter pumps.
-In a future update to this PDR document, we plan to show a full start-to-end beam dynamics study, based on the presented individual components’ beam simulations.
-We will also add a risk assessment, and additional details on the extraction system, based on the cited publications and ongoing design work.
-Finally, we will include a vacuum calculation based on MolFlow.
+Below, we summarize our designs and findings for the different subsystems (ion source and LEBT, RFQ, cyclotron, underground installation).
+5.1.1 Ion Source and LEBT Summary
+Based on literature and in-depth simulations using well-established codes, we have determined that a filament-driven multicusp ion source presents an attractive solution to producing 10-12 mA of DC $\mathrm{H}_{2}^{+}$ beam current with high purity and low emittance.
+The filament lifetime is the main drawback.
+We presented our design and measurements with such an ion source using our own MIST-1 prototype and first tests of the upgraded MIST-2 components.
+Our measurements agree extremely well with simulations and, although we have only reached about half of the design current with the MIST-1 prototype, we are confident that we have addressed all the “lessons learned” from it in our MIST-2 design and foresee no showstoppers in achieving the needed beam current with it.
+We estimate the filament lifetime to be about a week of continuous running, which would let us achieve the required 90% up-time.
+As a fallback, flat-field ECR ion sources are a tried-and-tested alternative, albeit with the (manageable) caveats of co-producing significant amounts of protons and having a higher emittance.
+The LEBT is a straightforward, highly compact design that includes safety mechanisms, beam positioning, and beam diagnostics.
+Our simulations include robustness tests and failure-mode simulations showing its capability to handle these.
+The beam is well matched into the RFQ.
+Parts of the LEBT have been manufactured with the remainder to be delivered at the same time as the RFQ.
+Experimental verification is thus forthcoming.
+5.1.2 RFQ Summary
+Here and through references to published papers, we presented the technical design of a split-coaxial RFQ of 28 cm diameter, operating at a frequency of 32.8 MHz (matched to the cyclotron frequency).
+The design includes thermal and mechanical vibration studies and full 3D PIC beam dynamics calculations including all fringe fields, using the LEBT output as input beam.
+The RFQ is currently being manufactured by Bevatech, GmbH in Germany.
+An experimental campaign has been started to put together ion source, LEBT and RFQ, and ultimately a 1.5 MeV/amu test cyclotron (the HCHC-1.5) for the ultimate demonstration of $\mathrm{H}_{2}^{+}$ acceleration, RFQ direct injection, and establishment of vortex motion.
+5.1.3 Cyclotron Design Summary
+We presented a realistic cyclotron mechanical and electrical design consisting of the main magnet, the RF Cavities, and vacuum chamber.
+The magnetic and electric fields calculated from these models by FEM and BEM were used to perform highly realistic PIC simulations using the OPAL code.
+Simulations merging the RFQ output with the spiral inflector in the central region have now also been performed; completing a simulation chain running seamlessly from initial plasma generation in the ion source into the central region of the cyclotron where acceleration will begin.
+Simulations of the acceleration beginning with the end of the central region and concluding with extraction at the target energy of 60 MeV/amu have been performed using an input beam that was informed by a standalone central region design.
+Connecting the latest spiral inflector simulations with a modified version of this central region inside the new magnet model will yield a full start-to-end simulation of the IsoDAR experiment.
+This section also discussed several improvements that have been explored to increase the efficiency, safety, and beam quality that emerges from the inflector into the cyclotron.
+Calculations performed with the full 3D model of the RF cavity showed manageable power and cooling.
+The required radial voltage distribution in the accelerating gaps, and adequate room for vacuum getter pumps.
+5.1.4 Installation Summary
+We described the layout of the Yemilab facility and how to transport the cyclotron steel components using SPMTs.
+In the cyclotron cavern, assembly will be aided by a gantry crane that can be disassembled after cyclotron installation to make room for the peripheral devices.
+We identified the main risk as the cyclotron coil size being on the edge of what can be transported through the main access ramp at Yemilab.
+A careful investigation of possible choke points is needed for the TDR.
+Possible mitigation is to wind the coil in place or to adopt a split coil design as we described in the CDR.
+For the next level of detail, we are working hand in hand with colleagues from Yemilab to identify further issues, devise an installation and removal plan, as well as overall integration plans with the facility (power, HVAC, controls, etc.).
+5.2 Outlook
+A future Technical Design Report will have to include a number of additional details, among them a full start-to-end simulation using as few beam dynamics codes as possible, benchmarked against a second set of codes; robustness analysis; a vacuum calculation (most likely using the CERN-developed code MolFlow); specifics of assembly and removal as well as shielding of the cyclotron cavern; RF system details; a commissioning strategy; data acquisition for beam diagnostics; and the cyclotron control system.
+We have investigated a number of these topics and published results thereon, but including them here would have gone beyond the level of a PDR.
+Others are current and ongoing studies and publications are forthcoming (well before the final TDR, much like we did here).
+Most notably, we are in the process of putting together the RFQ-Direct Injection Project (RFQ-DIP), combining the MIST-2 ion source (currently commissioning), the RFQ (currently being machined), and the HCHC-1.5 cyclotron (parts being ordered), as demonstrator of all novel concepts in the HCHC-60 cyclotron design.
 5.3 Acknowledgements
@@
 Preliminary work leading up to this was supported by NSF grants PHY-1912764 and PHY-1626069.
-Moon was supported by DOE grant DE-SC0024138.
-Winklehner was supported by the Heising-Simons Foundation.
+Work for V1.0 was supported by DOE grants DE-SC0024138 and DE-SC0024914 and NSF grant PHY-2411745.
+Winklehner was also supported by the Heising-Simons Foundation.
+Winkler thankfully acknowledges the support from the MIT UROP office.
+[…]

### v4

@@
-ToDo (#1): #2
-Abstract: This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source in two volumes which are mostly site-independent and describe the cyclotron driver, providing a 10 mA proton beam, and the medium energy beam transport line and target, respectively.
-The IsoDAR driver and target will produce a mole of electron-antineutrinos over the course of five years.
-Paired with a kton-scale liquid scintillator detector, it will enable an impressive particle physics program including searches for new symmetries, new interactions and new particles.
+
+Abstract
+This Preliminary Design Report (PDR) describes the IsoDAR electron-antineutrino source in two volumes which are mostly site-independent and describe the cyclotron driver providing a 60 MeV, 10 mA proton beam (this Volume); and the medium energy beam transport line (MEBT) and target (Volume II).
+The IsoDAR driver and target will produce about $1.15\cdot 10^{23}$ electron-antineutrinos over five years.
+Paired with a kton-scale liquid scintillator detector, it will enable a broad particle physics program including searches for new symmetries, new interactions and new particles.
 Here in Volume I, we describe the driver, which includes the ion source, low energy beam transport, and cyclotron.
-The latter features radiofrequency quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of so-called vortex motion.
-Section Editors (in alphabetical order): M.
-Abs1, J.R Alonso2, J.M.
-Conrad2, S.J.B.
-Engebretson2, E.
-Forton1, A.
-Herrod1, D.
-Joassin1, J.
-Moon2, S.
-de Neuter1, E.
-Van der Kraaij1, G.
-Wéry1, D.
-Winklehner2, E.
-Winkler2
-for the IsoDAR collaboration
-1Ion Beam Applications, 3 chemin du cyclotron Louvain-la-Neuve, BELGIUM
-2Massachusetts Institute of Technology, 77 Massachusetts Ave. Cambridge, MA, USA
-Corresponding Authors: Joshua Spitz (spitzj@umich.edu) and Daniel Winklehner (winklehn@mit.edu)
-V2.0
-Chapter 1 Introduction
-Figure 1.1: Schematic of the IsoDAR experiment installed at the Yemilab facility in South Korea.
-From right to left: The compact cyclotron with RFQ injector, the medium energy beam transport, the target, and the LSC neutrino detector.
-From Ref.
-[calvo_analyzing_2023].
-IsoDAR will be the first high-intensity, proton-driven source to be located alongside an underground multi-kiloton-scale scintillator detector.
-The first version of IsoDAR will be constructed at the Yemilab facility [alonso_isodaryemilab_2022] within the Handuk mine, located in South Korea where the “Liquid Scintillator Counter” (LSC) [^1] The LSC was renamed $\nu$Eye in March 2025, but for consistency throughout this document, we will refer to the detector as the LSC.
-is planned to be installed [seo_physics_2023].
-The novel source, consisting of a cyclotron, beam transport line, and a target/sleeve is shown in Fig.
-1.1.
-In five calendar years (four live years), the source will produce 1.67 million $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ (inverse beta decay or ”IBD”) events and 7000 $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$ (elastic scattering or ES) events in the LSC that can be used for searches for new physics.
-More than a mole of neutrons and mono-energetic photons are also produced in the target, and these may act as progenitors for new particles that enter the LSC and leave signatures from interactions and decays.
-IsoDAR can address a broad range of new physics manifested as new symmetries, new interactions and new particles that is inaccessible to traditional experiments.
-For further discussion see Refs.
-[alonso_neutrino_2022, waites_axionlike_2023].
-This document is Volume I in a two-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering solutions necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D.
+The latter features Radio-Frequency Quadrupole (RFQ) direct axial injection and represents the first accelerator purpose-built to make use of so-called vortex motion.
+keywords:
+Neutrinos, High-Intensity, Cyclotron, BSM Physics
+1 Introduction
+Figure 1: Main: The IsoDAR accelerator and target system adjacent to the LSC detector at Yemilab.
+Top: A to-scale rendering of the cyclotron overlaid on a photo of the cavern.
+The project comprises four major systems, three of which are shown: 1) the cyclotron; 2) the 60 MeV beam transport; and 3) the target and sleeve.
+The fourth is a data acquisition and monitoring system for all components.
+[…]
+This document is Volume I in a two-volume Preliminary Design Report (PDR) for IsoDAR that provides the engineering details necessary to establish credible solutions to the key issues and cost-drivers identified during the past decade of conceptual design and R&D.
 The contents of the PDR will cover the Technical Facility: the cyclotron, the beam transport and the target/sleeve.
-This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [abs_isodarkamland_2015].
-This step of producing a PDR substantially reduces risk related to IsoDAR approval: the remaining detailed engineering that will follow will build on these results and be straightforward to estimate in scope and cost.
-Volume I describes the IsoDAR proton driver.
-This consists of a “front end” that delivers ions to a cyclotron for acceleration and extraction and the cyclotron itself.
-This document presupposes a basic understanding of cyclotrons; for an introduction to cyclotrons and a cost/benefit analysis of this choice of machine for IsoDAR, see the discussion in the CDR.
+This document builds on the structure and text of the Conceptual Design Report (CDR) of the IsoDAR Technical Facility [3].
+This step of producing a PDR substantially reduces the risk related to IsoDAR approval.
+The remaining detailed engineering that will follow will build on these results and will be straightforward to estimate in terms of scope and cost.
+This document presupposes a basic understanding of cyclotrons. For an introduction to cyclotrons and a cost/benefit analysis of this choice of machine for IsoDAR, see the discussion in the CDR.
 The 60 MeV/amu IsoDAR Proton Driver is exceptional in that it is designed and engineered to produce an order of magnitude higher current proton beams at extraction than commercial machines of similar energy.
-Achieving this high power necessitated three key breakthroughs.
-First, the cyclotron accelerates $\mathrm{H}_{2}^{+}$, mitigating “space charge effects” (electromagnetic interactions between beam particles).
+Three novel concepts set this accelerator apart from commercial ones: First, the cyclotron accelerates $\mathrm{H}_{2}^{+}$, mitigating “space charge effects” (electromagnetic interactions between beam particles).
 Second, the beam is axially injected through an RFQ, acting as a high-efficiency buncher.
-Third, the design harnesses vortex motion, a complex effect previously observed at PSI Injector II [stetson:vortex, baumgarten:vortex1], inducing stable spiraling motion in the high-current accelerated beam, significantly reducing beam growth.
-For more concerning the accelerator physics that shapes the IsoDAR design, see Ref.
-[winklehner_order--magnitude_2022, winklehner_high_2018].
-These three physics requirements for reaching high current drive many of the engineering choices presented in this PDR.
-The other design driver is the need to construct the cyclotron underground.
-The cavern for the IsoDAR Proton Driver has already been constructed at Yemilab, and so the dimensions and constraints are understood, leading to fully realistic engineering plans.
+Third, the design harnesses “vortex motion”, a complex effect previously observed at the Paul Scherrer Institute at the PSI Injector II cyclotron [4, 5].
+Vortex motion induces spiraling within the individual beam bunches, providing compression and stability, which leads to substantially lower losses during extraction.
+For more information concerning the accelerator physics that shapes the IsoDAR design, see Refs.
+[6, 7].
+These three concepts drive many of the engineering choices presented in this PDR.
+The other design driver is the need to assemble the cyclotron underground.
+The cavern for the IsoDAR source has already been constructed at Yemilab, and so the dimensions and constraints are understood, leading to realistic engineering plans.
 This PDR is structured as follows:
-  • Chapter 1: (This chapter) proceeds with an overview of the physics searches that will be possible with IsoDAR, followed by an overview of the requirements for the proton driver.
-  • Chapter 2: Describes engineering related to the IsoDAR front end.
-  • Chapter 3: Describes engineering related to the general IsoDAR cyclotron design.
-  • Chapter 4: Describes specifics of the design related to installation underground.
-1.1 The IsoDAR Physics Case
-Figure 1.2: Energy dependence of fluxes.
-Left: $\bar{\nu}_{e}$, unit normalized; Middle: photons per proton on target (POT) from excited nuclei [alonso_neutrino_2022].
-Right: kinetic energy of neutrons at production in target and sleeve, normalized per proton.
-IsoDAR offers a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the proposed LSC at Yemilab [seo_physics_2023].
-The exceptional reach is owed to the high flux rates of $\bar{\nu}_{e}$, photons, and neutrons from the $\approx 2\cdot 10^{24}$ protons on target over 5 years (cf.
-Fig.
-1.2), and placing the source very close to the detector.
-Here we briefly discuss the highlights of the IsoDAR BSM physics program enabled by a 2.26 kton detector like the LSC in three categories: “New Particles” - experimentally and theoretically motivated; “New Symmetries” - CP-violation and the QCD Axion; And “New Interactions” - Precision Electroweak Physics.
-See Refs.
-[alonso_neutrino_2022, Hostert:2022ntu, waites_axionlike_2023] for further details.
-1.1.1 Fluxes in IsoDAR
-We consider three fluxes of particles produced in the IsoDAR target for BSM physics searches: the $\bar{\nu}_{e}$ flux, the neutron flux, and the $\gamma$ flux from the excitation of nuclei as neutrons are moderated.
-These fluxes are shown in Fig.
-1.2.
-The $\bar{\nu}_{e}$ flux stems from a single beta-decay of 8Li, which is produced by neutron capture on 7Li and is well-understood.
-It peaks at 6 MeV, which puts the majority of the events above backgrounds from the environment.
-These are clear advantages over reactor-neutrino fluxes.
-Interactions in the detector at these neutrino energies are exclusively inverse-beta-decay (IBD) and elastic scattering (ES).
-Cross sections for these reactions are known with uncertainties at $<1$% at these energies.
-Neutrons in the IsoDAR target are primarily produced by spallation interactions of the beam protons with the beryllium target hemispheres.
-Additional secondary neutrons stem from interactions of the primary neutrons with beryllium in the target and sleeve.
-We model the processes in the target with Geant4 using an accurate representation of the latest target and sleeve design (see Refs.
-[bungau_optimizing_2019]).
-We show the resulting neutron energy spectrum in Fig 1.2 (right).
-Most neutrons capture on the 7Li in the sleeve.
-To ensure $<5$ neutrons per year enter the LSC veto region, the target is further shielded using boron-loaded concrete and steel [bungau_shielding_2020].
-The third flux we consider is that of photons as seen in Fig.1.2 (middle).
-Here, one can observe a line structure that stems from the transitions of excited nuclei and and underlying smooth continuum due to bremsstrahlung.
-Again, the spectrum was generated using Geant4 using the full IsoDAR Target geometry [waites_axionlike_2023].
-1.1.2 New Particles: Experimentally and Theoretically Motivated
-Figure 1.3: An example of a potential 3+2 signal in IsoDAR, orange–true signature, black–IsoDAR reconstruction with uncertainty shown.
+  • Chapter 2: Proceeds with an overview of the physics searches that will be possible with IsoDAR.
+  • Chapter 3: Presents a brief overview of the requirements for the proton driver.
+  • Chapter 4: Describes engineering related to the IsoDAR front end.
+  • Chapter 5: Describes engineering related to the general IsoDAR cyclotron design.
+  • Chapter 6: Describes specifics of the design related to installation underground.
+2 IsoDAR Physics That Informs Design Goals
+The physics case for IsoDAR has driven the design decisions that are reported in this Preliminary Design Report.
+In this section, we provide a relatively high-level overview of the physics goals and connect them to the design needs, providing references for the reader to pursue more details.
+Today, particle physicists build massive underground detectors to capture low-energy interactions with unparalleled precision, while accelerator physicists develop powerful machines for new physics above ground; but these instruments have never been joined in close proximity.
+IsoDAR aims to change that paradigm.
+It will be the first high-intensity accelerator located underground, next to a multi-kiloton-scale detector, opening entirely new windows for particle physics.
+The designs presented in this document are largely site-independent.
+Still, we will use details of the planned initial run at Yemilab in Korea to provide a concrete description of the physics impact.
+Figure 1 shows a 3D rendering of IsoDAR’s planned installation at Yemilab.
+The volumes of this PDR focus on the IsoDAR source, which consists of the elements to the right of the LSC (cylinder at left).
+More information on the LSC is available in Refs.
+[2, 8, 9].
+Important quantities within this layout that impact the physics are listed in Table 1.
+Although this first volume concentrates only on the accelerator, the physics case relies on all systems, and this chapter will also be relevant to Volume 2 of the PDR.
+Table 1: Specifications for the IsoDAR@Yemilab design that directly impact the physics opportunities. See text for explanations of the live-time, fiducial volume cut and energy minimum cut.
+Accelerator Run Parameters
+Runtime 5 years
+Duty factor 80%
+Livetime 4 years
+Protons on IsoDAR target/year $1.97\cdot 10^{24}$
+Expected total antineutrino flux in 4 yr livetime
+$\bar{\nu}_{e}$ $1.15\cdot 10^{23}$
 […]
-Figure 1.4: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with 8Li-produced $\bar{\nu}_{e}$ background (gray).
+LSC design parameters
+LSC fiducial mass 2.26 kton
+Energy resolution at 3 MeV 3.7%
+Baseline distance from source 17 m center-to-center (9.5-25.6 m)
+Depth of LSC and source 2700 m.w.e.
+SM rates with livetime, energy and fiducial volume assumptions
+IBD ($\bar{\nu}_{e}+p\rightarrow e^{+}+n$) 1.67 million
+ES ($\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$) 7000
+2.1 Introduction to Design for BSM Searches in IsoDAR
+One typically searches for Beyond Standard Model (BSM) physics through its effect on Standard Model (SM) particles in three ways.
+First, through the production of a new particle via coupling to an SM particle that subsequently interacts or decays.
+Second, through the observation of new symmetries that introduce unexpected behavior of SM particles.
+Third, modifications to SM interactions due to new forces that couple to SM particles.
+From this, one sees that a crucial key to the discovery of new physics lies in the SM flux available to an experiment.
+The higher the quality and quantity of the SM flux, the deeper the reach in the search for new physics.
+The IsoDAR system focusses 60 MeV protons onto a beryllium target that is surrounded by a 99.99% pure 7Li sleeve that, in turn, is embedded in substantial shielding.
+Given this design, IsoDAR produces three unique SM fluxes for BSM searches: electron antineutrinos, monoenergetic photons and neutrons.
+We discuss the qualities and quantities of these fluxes below, and they are shown in Fig.
+2.
+These three SM fluxes fall into two categories of BSM physics that can be explored.
+The first category consists of SM particles that can exit the IsoDAR source and cross the shielding, ultimately producing signatures in the detector — electron antineutrinos.
+The second category is the set of copiously produced SM particles — the photons and neutrons that interact electromagnetically and strongly, respectively.
+A new BSM particle can couple to these contained fluxes and, if it has a very low interaction cross-section, will traverse the shielding and enter the detector.
+To be observed, the particle must subsequently mix with, interact with, or decay to SM particles that are then observable in the LSC.
+The IsoDAR design has three crucial advantages for BSM searches using these particles:
+  1. 1.
+    The rate of SM particle production is very high.
+  2. 2.
+    The solid angle to the LSC detector, 17 m from the source, is very large.
+  3. 3.
+    The cosmogenic and environmental backgrounds are very low.
+We consider the importance of each of these points below.
+Figure 2: Energy dependence of fluxes.
+Left: $\bar{\nu}_{e}$, normalized to unity; Middle: photons per Proton On Target (POT) from excited nuclei.
+Right: kinetic energy of neutrons at production in the target and sleeve, normalized per proton.
+2.2 Designing IsoDAR’s Electron Antineutrino Flux
+IsoDAR was originally designed for BSM searches using the electron-antineutrino flux.
+Although the physics program has since expanded to include state-of-the-art searches using both photon and neutron fluxes, the collaboration has prioritized the quality of the electron-antineutrino flux as a primary design goal.
+The quality issues that have been maximized by the design are: understanding the spatial dependence of the flux, understanding the energy dependence of the flux, and purity of antineutrino content (i.e., minimizing backgrounds).
+The quantity of the electron-antineutrino flux is also a design priority, where rates ahve been balanced against cost and practicality of installing and running at an underground site.
+The outcome of the design is a world-leading rate of electron-antineutrino interactions, which we quote for a live time of 4 years.
+IsoDAR will collect 1.67 million $\bar{\nu}_{e}+p\rightarrow e^{+}+n$ (inverse beta decay or “IBD”) events and 7000 $\bar{\nu}_{e}+e^{-}\rightarrow\bar{\nu}_{e}+e^{-}$ (elastic scattering or “ES”) events in five calendar years—unprecedented sample sizes to probe new interactions [10].
+These are the only two interactions open within the SM for $\bar{\nu}_{e}$ scattering at $\sim 6$ MeV.
+The IBD signature requires a coincidence between the initial positron light-flash and the subsequent neutron capture, which greatly reduces backgrounds compared to the single-flash ES signature.
+To further minimize background, both the positron (for IBD) and electron (for ES), signals are required to be reconstructed to >3>3 MeV and to be within the 2.26 kt (1.15 kt) fiducial volume for IBD (ES) events.
+The resulting backgrounds to the IBD signal are found to be negligible, while the assumed backgrounds for physics from the ES signal, with all cuts, are shown in Fig.
+3.
+Figure 3: The ES signal and background event rates expected in 5 years of running IsoDAR@Yemilab. No energy smearing has been applied to the distributions. From Ref. [10].
+Some of the neutrino physics goals require precise knowledge of the spatial distribution of the beam.
+To this end, an isotope decay-at-rest (hence the name IsoDAR) flux was selected.
+This flux has the advantage of being produced in $4\pi$ steradians, and the relative flux within a given solid angle is straightforward to determine.
+An isotope that $\beta$-decays with high endpoint (or Q-value) is desirable since the interaction rate in the detector increases with energy.
+After studying many isotopes, the collaboration has chosen 8Li, which has the well-predicted $\beta$ decay spectrum [11], shown in Fig.
+2, that peaks at approximately 6 MeV and has an endpoint of 16 MeV, minimizing systematic uncertainties on the energy dependence of the flux.
+By choosing a flux with high endpoint, we can apply a visible energy cut above 3 MeV, removing most environmental background events while retaining 10% of the total flux.
+The tradeoff is that a high Q-value corresponds to a short beta decay lifetime—839 ms for 8Li—requiring continuous replenishment via an accelerator-based system that produces neutrons that capture on 7Li.
+This leads to several design choices for the target, the surrounding “sleeve” and the accelerator that we discuss below.
+The design requires a target that generates high quantities of neutrons at energies that must then be moderated efficiently to allow capture.
+After evaluating alternatives and considering radiation handling issues, we selected 9Be as the target, since its final neutron is bound by only 1.7 eV.
+A tungsten target would release more neutrons per proton on target, but would also generate unwanted neutrinos in the flux.
+An important design choice was to cool the target using a 17-liter closed-loop system of D2O, as opposed to H2O, which has a high neutron capture cross section.
+The neutrons exit the target and enter a surrounding region with 7Li, which we refer to as the “sleeve.” This sleeve serves three purposes: to multiply neutrons, to moderate fast neutrons, and to provide the 7Li target for capture.
+Natural lithium is a mixture of 6Li and 7Li.
+Due to the very high 6Li neutron capture cross section, we require isotopic purification to 99.99% 7Li.
+Accessing 480 kg of 7Li at this purity has required coordination with industry.
+A cost-effective method to further enhance neutron production is to add beryllium to the sleeve.
+As the fast neutrons moderate, they inelastically scatter off the beryllium, releasing additional neutrons.
+We found the 8Li production rate is maximized with a 25% 7Li / 75% 9Be mixture.
+The goal is to maximize neutron absorption in 7Li while minimizing losses in all other elements of the target and sleeve and minimize background neutrino flux rates, hence the design focuses on low-Z materials with low capture cross-sections, including, for example, a carbon neutron reflector surrounding the sleeve.
+With the design presented here, the background neutrino production rate above the 3 MeV analysis cut is negligible compared to the signal anti-neutrino rate.
+[…]
+In the case of the ES single-flash signal, beam-off running allows measurement of environmental backgrounds above 3 MeV.
+This data can be collected during the 20% time period when the accelerator is down for maintenance.
+In the future, we may also investigate introducing beam chopping, which allows for nearly simultaneous beam-on and beam-off running.
+2.2.1 The Contained Monoenergetic Photon Flux
+As the neutrons moderate in the sleeve, nuclei are excited through collisions.
+The excited nuclei decay, producing monoenergetic photons.
+These are absorbed within the target and sleeve material; however, prior to absorption, the photons can potentially mix with BSM particles that can subsequently enter the detector and decay or interact inside.
+The total energy of the resulting BSM particle is also monoenergetic.
+This energy signature allows for discovery through searches for sharp peaks in the energy measured within the detector if the BSM particle decays to electrons.
+Broad peaks are produced if the BSM particle interacts to produce scattered electrons.
+These signatures have a single electromagnetic flash and, therefore, must be differentiated from those environmental backgrounds with energies higher than the 3 MeV analysis cut that are seen in Fig.
+3.
+However, beam-off running allows measurement of the environmental backgrounds and, if peaks are relatively narrow, the background estimate can be obtained by extrapolation from above and below the peak energies.
+The set of monoenergetic peaks produced by photons in the source region is seen in Fig.
+2.
+The highest flux is from isotopes produced by neutrons that are captured in the iron shielding that surrounds the source.
+We discuss design considerations for the shielding under the neutron flux, below.
+The second-highest peak, resulting from the breakup of deuterons in the cooling water, at 2.2 MeV, is below the 3 MeV analysis cut.
+Thus, at present, we do not consider possible BSM physics searches from this peak, but may do so in the future.
+Below these, a range of peaks extending above 10 MeV can provide progenitor photons for BSM particles.
+In summary, although the design has not been optimized for this physics (it is optimized for the electron antineutrino flux), this flux is sufficiently high to open doors to new physics.
+2.2.2 The Contained Neutron Flux
+The neutrons may also be progenitors of BSM particles that can penetrate the shielding around the source region.
+Neutron–dark sector couplings are motivated by the neutron beam/bottle experiment lifetime discrepancy [13, 14, 15, 16] .
+This allows them to disappear from the source area and reemerge in the LSC’s central and far side through $n\rightarrow n^{\prime}\rightarrow n$ transitions ($n\ldots$ neutron, $n^{\prime}\ldots$ dark sector neutron).
+IsoDAR has much higher sensitivity than surface experiments to this signature due to the large LSC detector and reduced backgrounds [17].
+The signature for the neutron reappearance is a single flash from neutron capture in the center of the LSC detector.
+The 2.2 MeV capture on H is below the 3 MeV cut for analysis, but the 4.95 MeV capture on C, which occurs 1% of the time, will be observable within the analysis cut.
+The most obvious background that must be controlled for this analysis is neutron punch-through from the source.
+IsoDAR has performed extensive GEANT-4 studies of shielding designs for Yemilab.
+These are presented in Ref.
+[18], but not discussed in detail in this preliminary design report, which discusses the elements of the design that are not Yemilab specific.
+(Yemilab specific designs will be collated into a separate design report.) The present design reduces neutron levels above 3 MeV reaching the outside of the Yemilab detector to just a few neutrons per year.
+For neutrons of any energy to penetrate from the outer wall of the detector to the fiducial volume, they must pass through 1.5 m of oil in the veto region, and 1.0 m of oil in the buffer region without capturing.
+If further protection from penetrating neutrons is desired, cuts can be placed in the fiducial region.
+In addition, we note that beam-off running allows studies of the environmental backgrounds.
+2.3 Connecting the Fluxes to the IsoDAR Physics Program
+The purpose of the design is to allow IsoDAR to offer a rich BSM physics program if paired with a kton-scale liquid scintillator detector like the LSC at Yemilab [2].
+Here we briefly discuss some of the unique capabilities of the IsoDAR BSM physics program, dividing the discussion into three categories: new particles, new symmetries, and new interactions.
+Details of these example analyses are beyond the scope of this design report, but can be found in the references.
+2.3.1 New Particles
+Figure 4: An example of a potential 3+2 signal in IsoDAR, orange–true signature, black–IsoDAR reconstruction with uncertainty shown.
+The true oscillations are dampened due to limited statistics and energy smearing.
+The oscillation parameters chosen for this example are allowed by today’s experiments [19].
+Example 1: Additional neutrinos—
+Many new physics models introduce additional neutrinos beyond the three Standard Model (SM) flavors that have feeble or even zero SM couplings, though they may interact through new BSM couplings [20].
+For light new neutrinos (with masses less than 10 MeV) it is feasible to design sensitive neutrino oscillation experiments.
+The probability that a neutrino interacts through the SM will vary with $L/E$, where $L$ is the distance traveled, and $E$ is the neutrino energy.
+The signature in the case of IsoDAR is a disappearance pattern of the $\bar{\nu}_{e}$ flux as a function of $L/E$ that will be distinct depending on the model.
+For example, Fig.
+4 shows the complex flux disappearance pattern that arises for the case where two additional light, non-interacting neutrinos are added to the three SM neutrinos, hence called “3+2,” for an example set of oscillation parameters allowed by today’s experiments [19].
+Table 1 gives the assumptions for the IsoDAR measurements in Fig.
+4.
+The fine measurement capability in $L/E$ will allow the observed data to precisely constrain the oscillation pattern [i.e., separating (3+1) versus (3+2)], greatly narrowing the theoretical models that can produce an effect.
+[…]
+Figure 5: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with 8Li-produced $\bar{\nu}_{e}$ background (gray).
 Right: $X$ mass and $X$-to-photon production ratio 90% CL sensitivity.
 From Ref.
-[alonso_neutrino_2022].
-Another example of IsoDAR’s capability to look for new particles is the “$X$,” which is a low mass mediator that would couple to the photons produced in the target [alonso_neutrino_2022].
-The $X$ would be a light new particle that decays to $\bar{\nu}\nu$ and would leave a distinct signature above background in the IBD spectrum.
-Fig.1.4 (left) displays resulting signatures for various masses (see color bar) assuming 0.5% $X$-to-photon production.
+[10].
+Example 2: Production of a new particle coupling to photons—
+An entirely new design enables searches for previously unknown particles.
+For an IsoDAR example, consider a new low mass mediator, which we will call “XX”, that mixes photons produced in the target and decays to $\bar{\nu}\nu$, where the $\bar{\nu}$ is observable through the IBD interaction [10].
+This analysis used the contained monoenergetic photon flux.
+The monoenergetic nature of the flux, hence the total energy of the $X$, leads to distinct peaks in the energy deposited by the $\bar{\nu}_{e}$ interactions.
+Fig.
+5 (left) displays resulting signatures for various masses (see color bar) assuming 0.5% $X$-to-photon production.
 The IBD background from $\bar{\nu}_{e}$ stemming from 8Li decays (gray) constitutes the primary background and has a distinct shape.
-The sensitivity is depicted in Fig.1.4 (right), where the blue region is excluded in the case of the SM.
-No existing limits are shown because this search is first-of-its-kind.
-A 17 MeV $X$ particle is consistent with the Atomki anomaly.
-1.1.3 New Symmetries: Strong CP Violation and the QCD Axion
-Figure 1.5: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [waites_axionlike_2023].
-Gold dots–axion allowed; Teal–excluded regions; Gray–parameters with astrophysical implications.
-The neutron beam/bottle experiment lifetime discrepancy [Nico:2004ie, Yue:2013qrc, Serebrov:2017bzo, UCNt:2021pcg] motivates neutron–dark sector couplings.
-Through $n\rightarrow n^{\prime}\rightarrow n$ transitions, we may see neutrons disappearing from the target and reappear in the LSC.
-The underground environment provides a low background and the LSC’s size further adds to the sensitivity [Hostert:2022ntu].
-Alternatively, if the $n^{\prime}$ are Majorana, then $n\rightarrow n^{\prime}\rightarrow\bar{n}$ may occur, depositing $\sim$2 GeV of energy in the LSC upon annihilation.
-This is a $B$ violating effect with implications for the early universe [Berezhiani:2020vbe].
-The photons and electrons in the target may also couple to axions and IsoDAR can probe uncharted territory with astrophysical and cosmological implications [waites_axionlike_2023].
+The sensitivity is depicted in Fig.
+5 (right), where the blue region is excluded in the case of the SM.
+No existing limits are shown because this search is a first-of-its-kind.
+2.3.2 New Symmetries
+Figure 6: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [21]. Gold dots–axion allowed; Teal–excluded regions; Gray–parameters with astrophysical implications.
+Example 1: A Dark Sector—
+Dark sector models were initially motivated by the existence of dark matter, which cannot be accommodated within the SM.
+The model postulates a “dark sector” of particles and forces that interact with the SM through “portals.” Although motivated by the dark matter puzzle, dark sector models have expanded to address other theoretical and experimental puzzles.
+IsoDAR can search for evidence of the dark sector through mixing of the SM neutron, $n$, to a dark-sector neutron, $n^{\prime}$.
+Along with general connections to dark matter, by way of a dark sector, this effect is motivated by the discrepancy in neutron lifetime measurement between neutron beam and neutron bottle experiments [13, 14, 15, 16].
+This analysis utilizes the contained neutron flux.
+Through $n\rightarrow n^{\prime}\rightarrow n$ transitions, neutrons can disappear from the target and reappear in the LSC.
+Along with the essential aspects of the source design discussed above, the underground environment that provides a low background and the LSC’s size, allowing tight fiducial cuts add to the sensitivity [17].
+Alternatively, if the $n^{\prime}$ are Majorana, then $n\rightarrow n^{\prime}\rightarrow\bar{n}$ may occur ($\bar{n}\ldotsanti$-neutron), depositing $\sim$2 GeV of energy in the LSC upon annihilation.
+This process is a $B$ violating effect with implications for the early universe [22].
+This contribution to dark-sector studies is unique within a wide range of proposed dark-sector experiments, as discussed in Ref.
+[23].
+Example 2: The QCD Axion—
+The QCD axion arises in theories that introduce a new symmetry to explain why CP violation is suppressed in QCD [24, 25].
+The introduction of the symmetry-breaking mechanism to make the new model parameter sufficiently small to agree with experimental data leads to the introduction of a new particle, the axion, $a$.
+There has been a long history of searches for the QCD axion at accelerators [25], and only a small open window of parameter space remains.
+IsoDAR can probe this territory [21].
+This search employs the contained monoenergetic photon flux, generating axions by their coupling to photons.
+Axions can also couple to electrons, and the signature in the LSC is from interactions and decays involving both electrons and photons.
 In Fig.
-1.5 we show QCD axions coupling to electrons.
-1.1.4 New Interactions: Tests via Precision Electroweak Physics
-IsoDAR will record about 7000 elastic scatter events over its 4-year accumulated up-time.
-This constitutes the world’s largest sample at low momentum transfer.
-LEP and LHC data predict rate and energy dependency very precisely within the standard model [ParticleDataGroup:2020ssz].
-With the normalization determined by the $\approx 1.7$M IBD events in the LSC, we will be able to measure $\sin^{2}\theta_{W}$ with a precision of 1.9% (see Fig.1.6 left).
-As a neutrino-nucleon experiment, the observed deviation of NuTeV may stem from unaccounted-for nuclear physics effects; however, numerous possibilities have been investigated, but none fully explain the deviation.
-Alternatively, the deviation might signal new physics in the neutrino sector.
-If so, this BSM signature will also manifest in the IsoDAR ES measurement.
-Fig.1.6, right, presents two expectations for IsoDAR, where the red assumes background reduction from Cherenkov ring identification in LSC, while the green indicates the expectation without this additional information.
-Statistics limit the outcome, and further running can proceed with IsoDAR and DUNE as it comes online if a NuTeV-like deviation is observed.
-Figure 1.6: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation.
-Right: expectation for IsoDAR in NSI parameter space.
-From Ref.
-[alonso_neutrino_2022].
+6 we show QCD axions coupling to electrons as an example of the capability.
+2.3.3 New Interactions
+Example 1: The Weak Mixing Angle —
+Despite many open questions within the SM, the theory is remarkably internally self-consistent within the electroweak sector [26].
+As a result, LEP and LHC data can be used to predict the rate and energy dependence of neutrino scattering precisely.
+Deviations might indicate new physics.
+The NuTeV experiment observed a deviation of the weak mixing angle ($\sin^{2}\theta_{W}$) parameter from the SM prediction [27].
+If this is new physics, it must be associated with neutrino properties, as polarized electron experiments have not observed the effect.
+However, as a neutrino-nucleon experiment, the NuTeV anomaly may stem from unaccounted-for nuclear physics effects; however, numerous possibilities have been investigated, but none fully explain the deviation and so the deviation remains unexplained [26].
+IsoDAR offers a new approach to this question through measurement using neutrino-electron elastic scattering—a purely leptonic process that evades the concerns related to nuclear physics in NuTeV.
+IsoDAR will record $\sim$7000 ES events over its 4-year accumulated up-time.
+This constitutes the world’s largest sample, allowing the most precise measurement of $\sin^{2}\theta_{W}$ at low momentum transfer, $Q$.
+Fig.7, right, presents two expectations for IsoDAR, where the red assumes background reduction from Cherenkov ring identification in LSC, while the green indicates the expectation without this additional information—both are presented since this design capability of LSC is yet to be determined.
+Total statistics limit the outcome, and further running can proceed with IsoDAR and DUNE as it comes online if a NuTeV-like deviation is observed.
+This result makes use of the $\bar{\nu}_{e}$ flux.
+An important design point is that the contemporaneous $\approx 1.7$M IBD events, which have a very well understood cross section, as discussed above, can be used to determine the flux normalization, allowing measurement of $\sin^{2}\theta_{W}$ with a precision of 1.9% (see Fig.7 left).
+Figure 7: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation. Right: expectation for IsoDAR in NSI parameter space. From Ref. [10].
+Example 2: Non-Standard Interactions of Neutrinos —
 The accurately predicted energy dependence of the ES sample enables a precise phenomenological search for non-standard neutrino interactions (NSIs).
 This powerful search for lepton-lepton couplings complements the lepton-quark coupling searches by the active, global coherent neutrino scattering program.
-The SM right- and left-handed couplings, $g_{R}$ and $g_{L}$, are altered by NSI corrections for electron flavor antineutrino interactions with electrons (hence ”ee”) to $\tilde{g}_{R}=g_{R}+\varepsilon_{ee}^{eR}$ and $\tilde{g}_{L}=g_{L}+\varepsilon_{ee}^{eL}$.
+The SM right- and left-handed couplings, $g_{R}$ and $g_{L}$, are altered by NSI corrections ($\epsilon_{ee}^{eR}$ and $\epsilon_{ee}^{eL}$) for electron flavor antineutrino interactions with electrons (hence “ee”) to $\tilde{g}_{R}=g_{R}+\epsilon_{ee}^{eR}$ and $\tilde{g}_{L}=g_{L}+\epsilon_{ee}^{eL}$.
 The SM expectation for IsoDAR@Yemilab in the parameter space of these deviations is seen in Fig.
-1.6 (right), where red assumes background reduction using Cherenkov light, and green assumes only scintillation light reconstruction.
-This can be compared to the world result [guzzo] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments and, hence, is rotated with respect to IsoDAR, which has a pure $\bar{\nu}_{e}$ flux.
-1.2 Requirements for the Proton Driver
+7 (right), where red assumes background reduction using Cherenkov light, and green assumes only scintillation light reconstruction.
+This can be compared to the world result [28] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments.
+This analysis uses the $\bar{\nu}_{e}$ flux.
+Because the IsoDAR result is very pure antineutrino, the expectation seen in Fig.
+7 (right) is rotated with respect to the past results, where the fluxes had a mix of $\nu_{e}$ and $\bar{\nu}_{e}$ content.
+The result complements searches for NSIs from coherent scattering experiments, which aim to discover new couplings between neutrinos and nucleons.
+For a review of NSI searches, see Ref.
+[20].
+3 Requirements for the Proton Driver
 As shown in Fig.
-1.1, the Proton Driver consists of a 60 MeV/amu cyclotron accelerating and extracting 5 electrical milliamperes of $\mathrm{H}_{2}^{+}$, a transport line that strips the $\mathrm{H}_{2}^{+}$ ions close to the extraction point, analyzes the stripped beam to monitor the health of the stripping foil, and transports the 10 milliamperes of protons to the target.
+1, the Proton Driver consists of a 60 MeV/amu cyclotron accelerating and extracting 5 mA of $\mathrm{H}_{2}^{+}$, a transport line that strips the $\mathrm{H}_{2}^{+}$ ions close to the extraction point, analyzes the stripped beam to monitor the condition of the stripping foil, and transports the 10 mA of protons to the target.
 The target consists of three nested hemispherical beryllium shells, the outermost being 20 cm in diameter, cooled with heavy water flowing around the shell structure.
-The proton beam is shaped, with quadrupoles and wobbler magnets to produce a distribution that allows for optimal heat transfer to the cooling water.
+The proton beam is shaped, with quadrupole magnets and wobbler magnets to produce a distribution that allows for optimal heat transfer to the cooling water.
 Neutrons produced in the target system are moderated and flood into a roughly-spherical sleeve, about 1 meter in diameter, containing a mixture of beryllium chips and highly-enriched 7Li.
 The beta-decay of the 8Li resulting from the capture of neutrons produces the neutrinos of interest in the experiment.
-The principal requirements for the accelerator system are to reliably produce a 10 mA beam of protons in the most efficient manner possible with the lowest possible loss of particles in the acceleration process.
-Particularly above the Coulomb threshold.
-The experience of the cyclotron system at PSI [Stammbach] is that keeping the total power of beam loss inside the cyclotron vault to less than 200 W still allows access for component maintenance.
-Considering that the total beam power is 600 kW, one sees that this is a very stringent requirement.
-As mentioned above, Volume I of this PDR addresses the physics and engineering aspects of the 60 MeV/amu cyclotron.
-The final chapter addresses the complex issue of transport and assembly of large and heavy cyclotron components in the underground environment at Yemilab.
-A forthcoming Volume II will address the details of the beam transport system and the target, along with shielding requirements for the environment of the target, including the prevention of fast neutrons from reaching the fiducial volume of the liquid scintillator detector.
+The principal requirements for the accelerator system are to reliably produce a 10 mA, 60 MeV beam of protons in the most efficient manner possible with the lowest possible loss of particles in the acceleration process, particularly above the Coulomb threshold.
+The experience of the cyclotron system at PSI [29] is that keeping the total power of beam loss during cyclotron extraction to less than 200 W still allows access for component maintenance.
+Considering that the total beam power is 600 kW, this is a very stringent requirement.
+To be decisive within five years, we require a CW beam with 80% availability.
+Secondary requirements like compactness stem from the underground installation.
+Section 6 addresses the complex issue of transport and assembly of large and heavy cyclotron components in the underground environment at Yemilab.
 3.1 Overview of the Accelerator Layout
-Figure 1.7: Overview of the components of the HCHC-60 Cyclotron System. Only one RF cavity is shown to not clutter the image.
+Figure 8: Overview of the components of the HCHC-60 cyclotron system.
+Only one RF cavity is shown to not clutter the image.
+“Quads+SI” refers to focusing quadrupoles leading into the spiral inflector (SI).
+The beam is produced in the ion source at the top, shaped and guided into the RFQ by the LEBT, which is partially embedded in the cyclotron, bent onto the horizontal plane by the spiral inflector, and then acceleratred in the cyclotron.
 To guide the reader through this PDR and, indeed, the path of an ion from genesis in the ion source to the full 60 MeV/amu and extraction, we provide an overview in Fig.
-1.7.
+8.
 The main components of the system, which will be described in much detail in the following chapters, are:
-  • The ion source, providing 10 mA of $\mathrm{H}_{2}^{+}$ at an emittance $<0.1\pi$-mm-mrad.
-  • The Low Energy Beam Transport (LEBT), shaping and guiding the beam to the RFQ and providing diagnostics, and chopping for commissioning and personnel- and machine protection.
+  • The ion source, providing 10 mA of $\mathrm{H}_{2}^{+}$ at an emittance $0.1\pi-$$m$-mrad.
+  • The Low Energy Beam Transport (LEBT), shaping and guiding the beam to the RFQ and providing diagnostics, and chopping for commissioning and personnel and machine protection.
   • The Radiofrequency Quadrupole (RFQ), embedded in the cyclotron yoke, to bunch and pre-accelerate the beam to 35 keV/amu.
@@
   • (not shown) The extraction system consisting of electrostatic and magnetic channels guiding the beam out of the cyclotron with the goal of extracting 5 mA of $\mathrm{H}_{2}^{+}$.
-Chapter 2 Design of the Front End
+3.2 Status of the HCHC-60 Development
+The HCHC-60 contains several subsystems that have never been used or combined in this way.
+To demonstrate the most critical aspects, we have begun an experimental campaign, building and testing some of these subsystems.
+Some already exist and are being tested, others are at the preliminary design level.
+Hence the overall status of the project is the PDR stage.
+The following gives a brief overview of the status of each subsystem.
+  • Ion Source. We have designed and built two iterations of a multicusp ion source (MIST-1 and MIST-2). MIST-2 is currently being experimentally tested.
+  • RFQ.
+The RFQ technical design is finished and it is currently being fabricated by Bevatech GmbH in Germany.
+We expect delivery in spring 2026 and will commission the RFQ together with the ion source.
+  • HCHC-1.5 test cyclotron.
+A smaller version of the HCHC-60, this 1.5 MeV/amu test cyclotron is fully funded and currently in the Request For Proposal (RFP) stage.
+We expect vendor selection by the end of summer 2025.
+  • HCHC-60.
+The full 60 MeV/amu cyclotron has reached the PDR stage.
+We describe the design and simulations in detail in the later sections of this document.
+We have not yet begun procurement of any hardware for the full cyclotron.
+The ion source and RFQ could be reused directly in the final HCHC-60 cyclotron.
+The central region of the HCHC-1.5, while designed to mimic as much as possible the HCHC-60 central region, will have to be adapted to the magnetic field of the HCHC-60.
+4 Design of the Front End
 4.1 The Ion Source
@@
 We have built a first prototype, the MIST-1 (Multicusp Ion Source at MIT - Version 1), the technical details and first measured results of which are reported in Refs.
-[axani_high_2015, winklehner_first_2018, winklehner_high-current_2021-1, winklehner_new_2022, weigel_epics_2023].
+[30, 31, 32, 33, 34].
 Here we summarize these results and present ongoing upgrade work (MIST-2).
-The upgrade to MIST-2 improves on the MIST-1 by implementing several lessons learned during the systematic measurements and runtime experiences described below.
+The upgrade to MIST-2 improves on MIST-1 by implementing refinements from lessons learned during the systematic measurements and runtime experiences described below.
+The design of the upgraded MIST-2 has been finalized and is in the final stages of fabrication and assembly.
 A publication on the MIST-2 design and performance is forthcoming.
 4.1.1 Operating Principle
-In a multicusp ion source [ehlers:multicusp1], a plasma is confined by a set of permanent magnets that are arranged around the plasma chamber with alternating polarity, thereby creating “cusps” in the magnetic field.
+In a multicusp ion source [35], a plasma is confined by a set of permanent magnets that are arranged around the plasma chamber with alternating polarity, thereby creating “cusps” in the magnetic field.
 On average, the electrons and ions produced in the center of the chamber see an increasing magnetic field as they travel outward, which eventually turns them back.
@@
 Electron-impact ionization with neutral molecules and atoms in the source chamber produces secondary electrons and ions.
-Neutral gas is fed into the source to facilitate the generation of a high-density plasma.
-The type of gas determines the ion species produced in the source.
-Here we use H2 gas to form a plasma – and beam – comprising protons, $\mathrm{H}_{2}^{+}$, and H+${}_{3}^{+}$.
-Table 2.1: A summary of the parameters for the MIST-1 ion source.
+[…]
+Table 2: A summary of the parameters for the MIST-1 ion source.
 Parameter Nominal Value
@@
 4.1.2 Mechanical Design
-Figure 2.1: Cross section of MIST-1 (left) with extraction system and LEBT. Also shown is an example set of ion trajectories simulated with IBSimu and WARP. From Ref. [waites_high_2022].
+Figure 9: Cross section of MIST-1 with extraction system, LEBT, and diagnostics.
+The diagnostics comprise an AC Current Transformer (ACCT) and a set of four capacitive button pickups right.
+Both are placed between the solenoid magnet and the RFQ at the end of the LEBT.
+Also shown is an example set of ion trajectories simulated with IBSimu and WARP.
+Adapted from Ref.
+[36].
 For visual reference as we discuss the details of the ion source, 3D CAD renderings of MIST-1 (with subsequent LEBT) and MIST-2 are provided in Figs.
-2.1 and 2.2, respectively.
-The main ion source parameters are also summarized in Table 2.1.
+9 and 10, respectively.
+The main ion source parameters are also summarized in Table 2.
 The MIST-type ion sources comprise a plasma chamber surrounded by a “water-jacket,” a back plate, and a front plate.
@@
 The radial confinement magnets are mounted inside the water-jacket and are cooled directly.
-At 6.5 cm, the plasma chamber is short compared to other ion sources.
+At 6.5 cm, the plasma chamber is short compared to typical ion sources.
 This is a direct consequence of the short mean free path of $\mathrm{H}_{2}^{+}$ in the plasma/neutral gas of the ion source discussed above.
@@
 The filament itself is made from tungsten and can have different thickness as well as shapes (this is a parameter currently being optimized).
-The front plate holds the front confinement magnets as well as the plasma aperture, a 75%W-25%Cu (aka “Elkonite”) plate that is embedded in the front plate.
+The front plate holds the front confinement magnets as well as the plasma aperture, a W-Cu (75% - 25%) alloy (aka “Elkonite”) plate that is embedded in the front plate.
 We have several plasma apertures with different-size openings ranging from 4 mm to 8 mm to test beam quality and total extractable current for various aperture diameters.
 As seen in Figs.
-2.1 and 2.2, following the plasma aperture is the “extraction system” which shapes and guides the beam as it is initially formed.
+9 and 10, following the plasma aperture is the “extraction system” which shapes and guides the beam as it is initially formed.
 The MIST-1 extraction system comprises five electrodes which shape and accelerate the ion beam.
-These include a plasma electrode (+15 kV), a puller electrode (-3 to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV).
-Figure 2.2: MIST-2 CAD rendering calling out the most important features.
-2.1.2.1 MIST-2 vs. MIST-1
+These include a plasma electrode (+15 kV), a puller electrode (-3 kV to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV).
+Figure 10: MIST-2 CAD rendering calling out the most important features.
+MIST-2 versus MIST-1
 The changes we made in the MIST-2 design are mainly to improve cooling, assembly and maintenance, and diagnostics. They are:
-  • Replace all stainless steel components with OFHC copper to improve heat transfer
-  • Increase the cross-section areas of all water cooling channels
+  • Replace all stainless steel components with oxygen-free high-conductivity (OFHC) copper to improve heat transfer
+  • Increase the cross-sectional areas of all water cooling channels
   • New high-power feedthroughs that are detachable from the back plate
@@
 These changes will improve long-term running stability, reduce maintenance downtime, and allow for a better understanding of the plasma and beam conditions.
-Our high-fidelity simulations using IBSimu [kalvasIBSIMUThreedimensionalSimulation2010a] and WARP [vayNovelMethodsParticleInCell2012] suggest that the beam dynamics downstream of the six-way cross will not change between the MIST-1 and MIST-2 configurations.
+Our high-fidelity simulations using IBSimu [37] and WARP [38] suggest that the beam dynamics downstream of the six-way cross will not change between the MIST-1 and MIST-2 configurations.
 All particle simulations and designs still hold for LEBT, RFQ, and cyclotron.
@@
 Hydrogen gas is introduced via an inlet in the backplate of the source as indicated in Fig.
-2.2.
-The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 SCCM full range.
+10.
+The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 standard cubic centimeter per minute (SCCM) full range.
 This system interfaces with the EPICS PC control system which is documented in detail in Ref.
-[weigel_epics_2023].
+[34].
 The filament is heated by an TDK Lambda (8.4 V, 300 A max) power supply.
@@
 A schematic of the wiring of the ion source is illustrated in Fig.
-2.3.
-The ion source (platform) is lifted to high voltage using a Matsusada AU 15 mA, 20 kV power supply, all lens elements in the extraction system are powered by Matsusada AU 7.5 mA, 20 kV power supplies of either negative or positive polarity.
+11.
+The ion source (platform) is lifted to high voltage using a Matsusada AU, 15 mA, 20 kV power supply, all lens elements in the extraction system are powered by Matsusada AU, 7.5 mA, 20 kV power supplies of either negative or positive polarity.
 The devices on the HV platform (control computer, filament discharge, filament heating power supply, etc.) are inside a Faraday cage on platform potential to reduce electromagnetic interference (EMF).
 Surrounding the platform and ion source is another Faraday cage at ground potential for personnel protection and EMF reduction.
-Figure 2.3: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Green lines are data connections. Black lines are power cables.
+Figure 11: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Green lines are data connections. Black lines are power cables.
 4.1.4 Beam Dynamics
-We simulate beam formation and shaping in the extraction system with the IBSimu package [kalvasIBSIMUThreedimensionalSimulation2010a], which performs cylindrically symmetric 2D (aka “RZ”) and 3D simulations including space charge through an iterative process.
+We simulated beam formation and shaping in the extraction system with the IBSimu package [37], which performs cylindrically symmetric 2D (aka “RZ”) and 3D simulations including space charge through an iterative process.
 IBSimu also contains a 1D plasma sheath model which is a good approximation for the plasma-beam interface at the source aperture.
-Our extraction simulations of the MIST-1 ion source followed by a WARP [vayNovelMethodsParticleInCell2012] simulation of the LEBT yielded excellent agreement with measurements [winklehner_first_2018, winklehner_new_2022].
+Our extraction simulations of the MIST-1 ion source followed by a WARP [38] simulation of the LEBT yielded excellent agreement with measurements [31, 33].
 Example trajectories of the new MIST-1 extraction system and RFQ matching LEBT are shown in Fig.
-2.1.
+9.
 An extraction system simulation of the MIST-2 system is shown in Fig.
-2.4.
+12.
 The simulated normalized RMS emittance for the new MIST-2 extraction system using an 8 mm diameter plasma aperture is 0.077 $\pi$-mm-mrad.
-Simulated normalized RMS emittances for the MIST-1 extraction system were between 0.05 and 0.1 $\pi$-mm-mrad depending on source aperture and extracted current and agreed well with measurements [winklehner_first_2018].
-Figure 2.4: IBSimu Simulation of MIST-2 extraction. Left: trajectories, Right: x-x’ phase space. Both MIST-1 and MIST-2 exhibit a very low emittance (see text).
+Simulated normalized RMS emittances for the MIST-1 extraction system were between 0.05 and 0.1 $\pi$-mm-mrad depending on source aperture and extracted current and agreed well with measurements [31].
+Figure 12: IBSimu Simulation of MIST-2 extraction.
+Left: particle trajectories in the z-x plane, Right: x-x’ phase space 25.5 cm after the extraction aperture.
+Both MIST-1 and MIST-2 exhibit a very low emittance (see text).
 4.1.5 Performance
-2.1.5.1 Goals & Measurement Methods
+Goals & Measurement Methods
 In this section we summarize the latest measured results as well as discussing the results of several optimizations that have been thus far performed as well as several that remain to be performed.
 The ion source target parameters that will be required for IsoDAR to achieve the planned decisive results for a sterile neutrino search include the following benchmarks:
-  • High current: $\approx$10 mA of H+${}_{2}^{+}$
-  • High purity: H+${}_{2}^{+}$ fraction $>$ 80%
-  • Low emittance: $<$ 0.1 $\pi$-mm-mrad (RMS, normalized)
+[…]
+[…]
+  • Low emittance: << 0.1 $\pi$-mm-mrad (RMS, normalized)
   • Ability to run continuously in DC mode
@@
 Allison-type electrostatic emittance scanners measure each species’ horizontal and vertical emittances.
-2.1.5.2 Optimization of Ion Source Parameters
+Optimization of Ion Source Parameters
 The ion source is designed such that several parameters can be varied in order to optimize the magnitude, purity, and emittance tightness of the current from the ion source.
@@
   • Permanent magnet type
-H2 input flow optimization.
-Three series of optimizations on the input gas flow rate have been performed.
-The three series differ in how the filament discharge voltage is managed.
-In the first series, the PID loop which stabilizes discharge was disabled allowing the discharge current to rise to the source steady-state value and then stabilized at that value by reengaging the PID.
-In the second series, the discharge current was kept fixed at 4 A.
-In the third series, the discharge current was reduced to keep the total extracted current fixed at 1.25 mA.
-Under each of these three conditions, the gas input rate was varied from 0.25 - 1.25 SCCM.
-This study demonstrates a general trend of greater H+${}_{2}^{+}$ production at higher discharge currents and, as theoretically expected, a higher fraction of H+${}_{3}^{+}$ with higher flow rates.
+[…]
 Discharge voltage optimization.
@@
 Discharge current optimization.
-One series of current optimizations was explored in which the flow rate and discharge were kept constant while discharge current was varied from 2 to 8 A.
+One series of current optimizations was explored in which the flow rate and discharge were kept constant while discharge current was varied from 2 A to 8 A.
 No strong dependence on current was observed.
-2.1.5.3 Latest Optimized Results
-To date, only an incomplete optimization of the ion source has been performed.
-As such, parameters which simultaneously meet all the target goals have not yet been fully identified.
-The ion source’s peak performance to date varies, therefore, depending on which of the performance metrics is optimized.
-Optimization has been performed for maximum total current density, maximum H+${}_{2}^{+}$ purity, and a hybrid option optimizing for the maximum H+${}_{2}^{+}$ current density with the constraint that H+${}_{2}^{+}$ be dominant.
-Highest total current density: 41 mA/cm2 with 31% H+${}_{2}^{+}$ purity.
-Highest H+${}_{2}^{+}$ purity: 76% purity with 11.4 mA/cm2 total current density.
-Highest H+${}_{2}^{+}$ current density with H+${}_{2}^{+}$ dominant: 10.4 mA/cm2 with 62% purity [^2] Note that this is the current density of $\mathrm{H}_{2}^{+}$ alone..
-With current optimizations and a 4mm aperture, the ion source is capable of delivering approximately 1.1 mA of H+${}_{2}^{+}$ with 76% purity.
-With an 8 mm aperture this corresponds to 4.4 mA.
-This is a world-leading result and is only about a factor of 2 short of the target.
-The main bottleneck for an increased beam current was the extraction system, which was designed for initial testing and commissioning.
+Latest Optimized Results
+[…]
+[…]
+[…]
+[…]
+[…]
+The main bottleneck for an increased beam current historically has been the extraction system, not the actual initial plasma generation.
+The extraction system characterized in this work was designed for initial testing and commissioning.
 We have recently replaced it with the version presented in this chapter.
-Our high-fidelity simulations suggest that the new extraction system will be able to handle the nominal beam current.
-Measurements are planned for summer 2024.
-2.1.5.4 Magnet Study
+Our high-fidelity simulations suggest that the new extraction system will be able to transport and focus the nominal beam current the source can generate, removing this bottleneck.
+Magnet Study
 The choice of permanent magnet material and strength is one of the optimization parameters of this type of ion source.
@@
 However, long-term running tests suggest that the magnet temperature is well controlled, which would allow the use of Nd2Fe14B magnets, if they yielded a better beam.
-We aim to maximize the $\mathrm{H}_{2}^{+}$ fraction of the MIST-1 ion source.
-Within the plasma, atomic processes include ionization, dissociation, and recombination, leading to protons, $\mathrm{H}_{2}^{+}$, and H+${}_{3}^{+}$ in the beam.
-Magnets with a higher field strength yield higher plasma density and longer confinement of the ions in the plasma, which can alter the rates of these processes.
-Thus, we are interested in how different magnet types alter the ion fraction and total current.
-We investigated the current and ion fraction of the beam with no magnets, Sm2Co17 magnets (grade 28), and Nd2Fe14B (grade N50H).
-The configuration without magnets provides no confinement, and the Nd2Fe14B magnets provide the strongest confinement.
-We performed our measurements using the diagnostic beam line presented in Ref.
-[winklehner_first_2018], containing a dipole magnet and a Faraday cup with 8 mm entrance slit for mass analysis.
-We investigated each magnet type at a range of H2 mass flows (0.25 SCCM to 1.5 SCCM) and discharge voltages (60 V to 140 V).
-In our preliminary analysis, we observed a trade-off between the total beam current and the $\mathrm{H}_{2}^{+}$ fraction: with lower confinement, the $\mathrm{H}_{2}^{+}$ fraction was high, but the current was low.
-This result was expected; lower confinement and lower plasma density leads to less time and fewer reactants for the ions to undergo processes that result in $\mathrm{H}_{2}^{+}$ dissociation or the pick-up of another proton (producing H+${}_{3}^{+}$).
-However, the lower plasma density also leads to a lower total beam current.
+[…]
+[…]
 For MIST-2, the baseline design is now to use Nd2Fe14B magnets because of the higher current they provide. A paper on the detailed analysis of the magnet study is forthcoming.
-2.1.5.5 Results with the MIST-2 back plate
-Figure 2.5: A typical spectrum with the new MIST-2 back plate.
+Results with the MIST-2 back plate
+Figure 13: A typical spectrum with the new MIST-2 back plate.
 The total beam current was 3.43 mA, the $\mathrm{H}_{2}^{+}$ fraction $\approx 60$%.
-For a discussion about the small contaminant peak, cf.
-text.
-We already performed the first test of the MIST-2 components described in Section 2.1.2.1 by installing the new back plate (which is backward compatible) on MIST-1.
+See main text for a discussion on the small contaminant peak.
+We already performed the first test of the MIST-2 components described in Section 4.1.2 by installing the new back plate (which is backward compatible) on MIST-1.
 We were able to run the source up to a stable beam current of 3.43 mA.
@@
 A sample spectrum from this series of studies is shown in Fig.
-2.5.
+13.
 Here, a small cluster of beam contaminants can still be seen, which we attribute to insufficient baking and out-gassing of source components, including Viton O-rings.
-Past experience has shown that vacuum baking the O-rings before installation (done here), overall good vacuum conditions, and baking the source and beam line before operation (not done in this instance), decreases the contaminants to a negligible level after a few hours of operation [winklehner_first_2018].
+Past experience has shown that vacuum baking the O-rings before installation (done here), overall good vacuum conditions, and baking the source and beam line before operation (not done in this instance), decreases the contaminants to a negligible level after a few hours of operation [31].
 4.1.6 Risks and Mitigation
@@
 Mitigation: Our current upgrade path should lead us comfortably to a total extracted current of 12 mA, particularly considering that MIST-1 has demonstrated an exceptionally low emittance, which opens up the possibility of going to even larger extraction apertures (10 mm or 12 mm).
-However, we also have tested a 2.45 GHz flat-field ECR ion source (the VIS [alonso_isodar_2015]) in the past, from which we measured 15 mA of $\mathrm{H}_{2}^{+}$.
+However, we also have tested a 2.45 GHz flat-field ECR ion source (the VIS [39]) in the past, from which we measured 15 mA of $\mathrm{H}_{2}^{+}$.
 The low ratio of 50% $\mathrm{H}_{2}^{+}$ (30 mA total current) deterred us, but as a mitigation, a VIS-type ion source paired with a dipole magnet is a viable alternative.
@@
     Provide safety mechanisms for personnel and machine protection.
-Details of the LEBT design and performance were given in Refs. [waites_matching_2021, waites_low_2022, waites_high_2022] here we summarize and update these.
+Details of the LEBT design and performance were given in Refs.
+[40, 41, 36] here we summarize and update these.
+It should be noted that the full LEBT design for the HCHC-XX as described here is at the level of a PDR.
+The only exception is the system of copper electrodes shown in Fig.
+9, which we already tested and reported on in the previous section.
 4.2.1 Design Overview
@@
 The beam now passes through a gap surrounded by a series of six electrodes arranged axially within a six-way cross (see Fig.
-2.1).
+9).
 These six electrodes can have their voltage optimized in order to best match the parameters for acceptance into the rest of the LEBT.
@@
 The conical shape minimizes beam loss on exit.
-Figure 2.6: Left: Photo of Alpha Magnetics, Inc.
-steerers [2562NotreDame].
+Figure 14: Left: Photo of Alpha Magnetics, Inc.
+magnetic steerers [42].
 Right: Result of LEBT steering simulations.
@@
 From Ref.
-[waites_high_2022].
+[36].
 Following the six-way cross lies a diagnostic segment.
@@
 The diagnostic section also includes a solenoid and two magnetic steerers which can be used for additional beam tuning.
-A flange in place toward the end allows space for an AC Current Transformer (ACCT) [bergozACCT] which characterizes the beam immediately prior to entering the RFQ.
+A flange in place toward the end allows space for an AC Current Transformer (ACCT) [43] which characterizes the beam immediately prior to entering the RFQ.
 The diagnostic segment is illustrated in Fig.
-2.1.
+9.
 4.2.2 Beam Steering & Chopping
@@
 The effect of the steerers is demonstrated in Fig.
-2.6.
+14.
 In order to control the amount of power delivered from the ion source, a beam chopper is built into the cross of the diagnostic segment to allow for a duty factor in an otherwise DC beam.
@@
 See Fig.
-2.8.
+16.
 This method permits controlling the average beam power, without changing the charger per bunch resulting from the RFQ.
@@
 The system was modeled by assuming a water cooling line at room temperature.
-To be conservative, the simulation was run using four times the actual anticipated beam power, i.e.
-720 W.
+To be conservative, the simulation was run using four times the actual anticipated beam power, i.e., 720 W.
 The beam dump never rose above 360 K in our simulation, significantly below the melting point of copper.
 The COMSOL simulation is shown in Fig.
-2.8
-Figure 2.7: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid.
+16
+Figure 15: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid.
 If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring.
 From Ref.
-[waites_high_2022].
-Figure 2.8: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior.From Ref. [waites_high_2022].
+[36].
+Figure 16: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior. From Ref. [36].
 4.2.3 Beam Dynamics
@@
 However, as is well-known, in the absence of external electrical fields, slow secondary charged particles from the interaction of the beam with the residual gas can be accumulated inside the beam envelope, effectively neutralizing the beam charge.
-This process is called “space charge compensation” [winklehnerSpaceChargeCompensation2015].
+This process is called “space charge compensation” [44].
 In all LEBT simulations, we estimate space charge compensation according to the model presented in Ref.
-[winklehnerSpaceChargeCompensation2015].
+[44].
 It is usually above 80%, alleviating the detrimental effect of space charge in some parts of the LEBT.
 However, we also test our designs with no compensation and find that, while beam losses are slightly higher, we can still obtain the required beam inside the cyclotron, albeit requiring about 20% higher beam from the source.
-For the IBSimu simulations a 12 mA beam composed of 80% H+${}_{2}^{+}$, 10% H+, and 10% H+${}_{3}^{+}$ was used.
-This provides ample current for IsoDAR assuming 50% global transmission efficiency while using current from the ion source that has already been empirically demonstrated.
-As illustrated in Fig.
-2.9 and in Table 2.2 the LEBT can match the required input parameters for the RFQ.
-Moreover, as illustrated in Fig.
-2.10 this match is achieved with emittance that surpasses that used for baseline RFQ simulations.
-Figure 2.9: A temperature plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT. [waites_high_2022]
-Figure 2.10: Phase space from the LEBT simulation (left) vs phase space for ideal beam input to the RFQ (right). From Ref. [waites_high_2022].
-Table 2.2: A comparison of LEBT Twiss parameters and baseline RFQ parameters.
+[…]
+Figure 17: A contour plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT. From Ref. [36]
+Figure 18: Phase space from the LEBT simulation (left) versus the phase space for the ideal beam input to the RFQ (right). From Ref. [36].
+Table 3: A comparison of LEBT Twiss parameters and baseline RFQ parameters.
 Parameter LEBT Output Baseline
-[…]
+Normalized RMS Emittance 0.175 $\pi$-$m$-mrad 0.3 $\pi$-$m$-mrad
 $\alpha$ 1.3 1.0
 $\beta$ 0.13 mm/mrad 0.17 mm/mrad
-2.2.4 Risks and Mitigations
+4.2.4 Risks and Mitigation
 Risk: Unexpected Power Loss.
@@
 4.3 The Radiofrequency Quadrupole
-Figure 2.11: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle. From Ref. [waites_high_2022].
-The $\mathrm{H}_{2}^{+}$ beam coming from the LEBT is injected into a Radio Frequency Quadrupole (RFQ) that is partially embedded in the cyclotron yoke.
+Figure 19: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle. From Ref. [36].
+The $\mathrm{H}_{2}^{+}$ beam coming from the LEBT will be injected into a Radio Frequency Quadrupole (RFQ) that is partially embedded in the cyclotron yoke.
 A schematic of the RFQ and cyclotron can be seen in Fig.
-3.1 where only about a third of the RFQ is visible above the top surface of the cyclotron.
+21 where only about a third of the RFQ is visible above the top surface of the cyclotron.
 We show a cut view of the RFQ in Fig.
-2.11.
+19.
 Of note are the small diameter of the RFQ (27.6 cm) while operating at a low resonant frequency of 32.8 MHz.
 This is achieved by utilizing the so-called “split-coaxial” mode where the vanes are attached longitudinally at the entrance and exit flanges instead of radially at the mantle.
-The main operating parameters of the RFQ are listed in Tab.
-2.3.
+The main operating parameters of the RFQ are listed in Table 4.
 The RFQ accelerates only moderately from 7.5 keV/amu to 35 keV/amu and acts mainly as a very efficient buncher.
-This leads to a conservatively estimated power consumption of $<6$ kW.
-Table 2.3: A summary of the design parameters for the RFQ.
+This leads to a conservatively estimated power consumption of <6<6 kW.
+Table 4: A summary of the design parameters for the RFQ.
 Parameter Nominal Value
@@
 Frequency 32.8 MHz
-Tuning range ±$\pm 270$ kHz
+Tuning range $\pm 270$ kHz
 Q simulated 2800
@@
 Duty factor 100% (CW)
-Power $<6$ kW
+Power <6<6 kW
 Cooling DI Water channels in tank and vanes
@@
 The technical design was presented in Refs.
-[ratzingerTechnicalDesignReport2020, winklehner_high-current_2021, holtermann_technical_2021] and the optimization of the RF loop coupler in Ref.
-[sangroula_design_2021].
+[45, 46, 47] and the optimization of the RF loop coupler in Ref.
+[48].
 The thermal properties and cooling considerations were presented in Ref.
-[koser_thermal_2021].
+[49].
 An optimization scheme using machine learning was presented in Ref.
-[koser_input_2022].
-Figure 2.12: Modes of the mechanical deformation of the RFQ vanes. The frequencies are 18 Hz and 25 Hz. From Ref. [holtermann_technical_2021].
+[50].
+Figure 20: Modes of the mechanical deformation of the RFQ vanes calculated with COMSOL. The frequencies are 18 Hz and 25 Hz. From Ref. [47].
 In terms of mechanical stability, our analyses show two quarter wave modes which are associated with the choice of a split-coaxial resonator.
 The addition of bridges between opposite vanes (see Fig.
-2.12) significantly reduces detuning due to the Lorentz force with mechanical frequencies around 18 Hz and 25 Hz.
+20) significantly reduces detuning due to the Lorentz force with mechanical frequencies around 18 Hz and 25 Hz.
 4.3.2 Beam Dynamics
@@
 This was performed by the subcontracted company Bevatech GmbH in Germany, who is currently also manufacturing the RFQ.
-The well-established tracking code TraceWin was used [tracewin].
+The well-established tracking code TraceWin was used [51].
 Fine tuning of the phase of the exit gap (i.e.
-tuning the length of a drift preceding the exit gap), leads to final output beam parameters listed in Tab.
-2.4.
-Table 2.4: RFQ output beam parameters.
+tuning the length of a drift preceding the exit gap), leads to final output beam parameters listed in Table 5.
+Table 5: RFQ output beam parameters.
 Parameter Nominal Value
-$\varepsilon_{x}$ (1-RMS, norm.) 0.54 mm-mrad
-$\varepsilon_{y}$ (1-RMS, norm.) 0.52 mm-mrad
-$\varepsilon_{z}$ (1-RMS, norm.) 0.114 MeV-deg
-Phase width (1-RMS) 40
+$\epsilon_{x}$ (RMS, norm.) 0.54 mm-mrad
+$\epsilon_{y}$ (RMS, norm.) 0.52 mm-mrad
+$\epsilon_{z}$ (RMS, norm.) 0.114 MeV-deg
+Phase width (RMS) 40
 4.3.3 Risks and Mitigation
@@
 More aggressive collimation is possible as well, at the expense of requiring higher current from the source.
-Chapter 3 Design of the Cyclotron
-Figure 3.1: Isometric view of the cyclotron.
+5 Design of the Cyclotron
+Figure 21: Isometric view of the cyclotron.
 In this section, we describe the details of the cyclotron at the level of a preliminary design report.
 This section includes details of the design of the cyclotron, the spiral inflector that transmits ions from the RFQ into the cyclotron, and the extraction system.
-This section also characterizes the latest in simulo performance of these systems.
+This section also characterizes the latest simulated performance of these systems.
 5.1 Magnet Yoke
@@
 As we finalize the details of beam extraction small modifications to the magnets, especially at the fringes, may accordingly be modified.
-The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron (seen in Fig 3.2), which is slightly larger, at 7 meters outer diameter.
-Figure 3.2: C400 on the milling machine, C230 in front of it. Courtesy of IBA.
+The engineering design of the IsoDAR magnet benefits from the experience gained by IBA in the fabrication of its C400 cyclotron , which is slightly larger, at 7 m outer diameter [52].
 5.1.1 Design and Fabrication of the Cyclotron Magnet
-As with the C400, the IsoDAR cyclotron is divided into smaller parts to facilitate transport to the site.
+As with the C400, the IsoDAR cyclotron will be divided into smaller parts to facilitate transport to the site.
 The magnet is split radially and will be assembled on site.
@@
 An exploded visualization of the upper and lower parts is shown in Figs.
-3.4 and 3.4.
+23 and 23.
 The material used for the magnet parts is cyclotron steel (the exact composition shall be determined after a detailed physics study), the vacuum chambers will be made of stainless steel.
-Figure 3.3: Cyclotron upper steel exploded.
-Figure 3.4: Cyclotron lower steel exploded.
-An cut-section view of the cyclotron is shown in Fig.
-3.5.
+Figure 22: HCHC-60 cyclotron upper steel exploded.
+Figure 23: HCHC-60 cyclotron lower steel exploded.
+A cut-section view of the cyclotron is shown in Fig.
+24.
 The external and vacuum disks are connected to each other as well as to the flux return using specific bolts (such as the Superbolt from Nord-Lock).
@@
 A detailed view of the cyclotron components is shown in Fig.
-3.6 and additional details about the yoke parts can be found in Table 3.1
-Figure 3.5: Cyclotron cut view
-Figure 3.6: Cyclotron detailed cut view
-Name Mass (Metric Ton) Dimensions (mm)
-Upper and lower external disk 30 Ø4700*250
-Upper and lower internal disk 36 Ø4620*463
-Upper and lower flux return (1 and 2) 36 3100*1950*1015
-Upper and lower median plane flux return (1 and 2) 29 3100*1950*705
-Upper and lower pole 7 2013*1290*600
-Table 3.1: Yoke part properties
-3.2 Cyclotron magnetic performance
-In this section we describe the latest modifications applied to the cyclotron regarding its magnetic properties.
-We will compare the results with the model from the Conceptual Design Report (CDR) from [abs_isodarkamland_2015] to ensure its magnetic performance remains within specifications.
+25 and additional details about the yoke parts can be found in Table 6
+Figure 24: Cyclotron cut view
+Figure 25: HCHC-60 cyclotron, detailed cut view calling out the major components of the steel, vacuum chamber, and coils.
+Table 6: Steel yoke, individual parts masses and dimensions.
+Name Mass (t) Dimensions (mm)
+Upper and lower external disk 30 $4700\times 250$
+Upper and lower internal disk 36 $4620\times 463$
+Upper and lower flux return (1 and 2) 36 $3100\times 1950\times 1015$
+Upper and lower median plane flux return (1 and 2) 29 $3100\times 1950\times 705$
+Upper and lower pole 7 $2013\times 1290\times 600$
+5.2 Simulated cyclotron magnetic performance
+In this section we describe the latest modifications applied to the cyclotron design regarding its magnetic properties.
+We will compare these results with the model from the Conceptual Design Report (CDR) from [3] to ensure its theoretical magnetic performance remains within specifications.
 With the latest design of the vacuum chamber, see Section 5.6, the pole radius was slightly reduced.
 To keep the magnetic field isochronous up to extraction at 120 MeV, the outer radius required small modifications, see left image in Fig.
-3.7.
+26.
 First, the curvature was modified to follow the curvature of the last orbit.
@@
 This reduction of the pole gap should have no effect on the transmission and can be tuned to improve the dB/dr for the last orbits.
-Figure 3.7: Left: Detailed view of the outer radius of the pole; only half of one pole is shown.
+Figure 26: Left: Detailed view of the outer radius of the pole; only half of one pole is shown.
 In red is the coil, in light green is the pole made of SNVA6342 steel.
@@
 The steel used for the cyclotron yoke and poles is the SNVA6342, also used, for example, in the C400 cyclotron (see Ref.
-[bib:c400_iron]).
-Its BH-curve is shown in Fig.3.7.
+[53]).
+Its BH-curve is shown in Fig.26.
 Here, we also see the better performance of the permendur alloy.
@@
 In Fig.
-3.8 the latest results are given.
-The left image shows the shimming applied, the right image shows the offset in magnetic field, as required for a given RF-frequency.
+27, the latest results are given.
+The left image shows the shimming applied, the right image shows the offset in magnetic field, as required for a given RF frequency.
 For comparison, the offset (expressed as dB in Gauss) for the model of the CDR is also shown.
-Figure 3.8: Left: Pole edge of isochronized model.
-Right: The dB [Gauss] as a function of radius [meters], for the CDR model (in black) and the newest model (in red).
-Note that a positive dB means there is too little field at specific radius, a negative dB implies there is too much field.
-The RF-frequency for the CDR model is at 32.77 MHz, for the newest model it is 32.82 MHz.
-The plot shows that the magnetic field is stable and strong enough up to a radius of 2 meters, which is the equivalent of 121 MeV.
-In Fig 3.9 the phase slip can be seen to stay within the ±$\pm 10^{\circ}$.
+[…]
+The plot shows that the magnetic field is stable and strong enough up to a radius of 2 m, which is the equivalent of 121 MeV.
+In Fig.
+28, the phase slip can be seen to stay within the $\pm 10^{\circ}$.
 The horizontal and vertical tunes are also given in the same figure.
 Here we see that there is a crossing of the tunes at high radius, invoking resonances which could impact the extraction.
-Figure 3.9: Left: phase slip as a function of radius.
+Figure 28: Left: phase slip as a function of radius.
 Right: vertical and horizontal tunes as a function of energy (dashed and solid lines respectively).
@@
 5.3 Interface with the RFQ
-Figure 3.10: RFQ overview.
-The RFQ will sit on top of the cyclotron along its central axis, as shown in Fig 3.10.
+The RFQ will sit on top of the cyclotron along its central axis, as shown in Fig.
+21.
 This places the RFQ outside of the cyclotron’s vacuum.
 Therefore, a vacuum sleeve will be used to ensure the integrity of the cyclotron’s vacuum.
-This sleeve will be fixed on top of the upper external disk and the vacuum sealing will be done radially on the upper internal disk (as seen in figure number 3.12 ).
-The O-rings willhave a cross sectional diameter of 8mm and an inner diameter of 305mm.
-In order to allow for adjustment, the RFQ sleeve allows a rotation of 5mrad around the lower O-ring contact patch (discussed with the manufacturer of the RFQ).
+This sleeve will be fixed on top of the upper external disk, and the vacuum sealing will be done radially on the upper internal disk (see Fig.
+29 ).
+The O-rings will have a cross-sectional diameter of 8 mm and an inner diameter of 305 mm.
+To allow for adjustment, the RFQ sleeve permits a rotation of 5 mrad around the lower O-ring contact patch (as discussed with the RFQ manufacturer).
 The RFQ sleeve is as thin as possible to avoid any loss in magnetic field in the central region (hence the O-rings in the yoke parts).
-Figure 3.11: RFQ zoom.
-Figure 3.12: RFQ cut section 1.
+Figure 29: CAD rendering of the RFQ inside the cyclotron yoke. The vacuum sleeve is called out in magenta.
+Figure 30: CAD rendering of the RFQ inside the cyclotron yoke. The vacuum sleeve is called out in magenta and the O-ring forming the vacuum interface between RFQ and sleeve is zoomed in.
 As shown in Fig.
-3.13, vacuum sealing will occur at the base of the RFQ using a thick O-ring (diameter 10 mm) which will self-seal under the RFQ’s weight ($\approx 3000$ N).
-The dimensions of the vacuum sleeve are listed in Table 3.2.
-In order to allow fine-tuning the cyclotron alignment, a set of adjustment feet, shown in Fig.
-3.11, are fixed to the cyclotron yoke.
+30, vacuum sealing will occur at the base of the RFQ using a thick O-ring (diameter 10 mm) which will self-seal under the RFQ’s weight ($\approx 3000$ N).
+The dimensions of the vacuum sleeve are listed in Table 7.
+To allow fine-tuning the cyclotron alignment, a set of adjustment feet, shown in Fig.
+31, are fixed to the cyclotron yoke.
 By using four groups of set screws and one vertical adjustment nut, the following alignment modifications can be made.
-  • Rotation: ±$\pm 5$
-  • Tilt: ±$\pm 5$ mrad
-  • Lateral movement: ±$\pm 1$ mm
+  • Rotation: $\pm 5$
+  • Tilt: $\pm 5$ mrad
+  • Lateral movement: $\pm 1$ $m$
 In addition to this fine-tuned alignment, a coarse alignment step is envisioned during the initial assembly and commissioning procedure.
-Figure 3.13: RFQ cut section 2.
-Table 3.2: Vacuum Sleeve Dimensions.
+Figure 31: CAD rendering of the top of the RFQ above the cyclotron top surface. The adjustment feet are called out in yellow.
+Table 7: Vacuum Sleeve Dimensions.
 Name Value
@@
 5.4 Central Region
-Figure 3.14: CAD rendering of the spiral inflector and particle trajectories as implemented previously for early tests at Best Cyclotron Systems, Inc.
-[alonso_isodar_2015].
+Figure 32: CAD rendering of the spiral inflector and particle trajectories as implemented previously for early tests at Best Cyclotron Systems, Inc.
+[39].
 The positive and negative electrostatic deflectors are shown in blue and green respectively.
@@
 From Ref.
-[winklehner_realistic_2017].
+[54].
 The injection system design is a crucial component that allows us to achieve a high beam current.
@@
 As in commercial cyclotrons, an axial injection system based on a Spiral Inflector (SI) will be used to bend the beam from the axial direction to the median plane.
-We have investigated several refinements to a traditional spiral inflector which will be discussed subsequently, but the concept, as implemented successfully in a previous iteration of this work and tested at Best Cyclotron Systems, Inc., is visualized in Figure 3.14 using trajectories calculated with OPAL [adelmann_opal_2019].
+We have investigated several refinements to a traditional spiral inflector which will be discussed subsequently, but the concept, as implemented successfully in a previous iteration of this work and tested at Best Cyclotron Systems, Inc.
+(BCSI), is visualized in Fig.
+32 using trajectories calculated with OPAL [55].
 The exit point of the spiral inflector is also the starting point of acceleration in the Central Region (CR) of the cyclotron.
@@
 The central region and the spiral inflector must be designed carefully and together because they have strong interplay.
-This procedure has already undergone one past iteration in the summers of 2013 and 2014 at Best Cyclotron Systems Inc.
+This procedure has already undergone one past iteration in the summers of 2013 and 2014 at BCSI.
 In the following sections, we will discuss improvements to this design which will be incorporated into the preliminary design for the full IsoDAR machine.
 5.4.1 Spiral Inflector Overview & Design Considerations
-The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Figure 3.14).
+The axial injection of the ionized beam into a cyclotron is realized using an electrostatic device called a spiral inflector, which consists of two curved electrode deflectors (see Fig.
+32).
 The electrostatic potential between these electrodes is able to bend the beam 90 from the axial line to the median plane of the cyclotron.
@@
 However when 3D magnetic variations and/or off center trajectories are considered a numerical approach is required.
-An initial estimate based on an analytical solution for a central track serves as a numerical starting point, after which the mesh generating software gmsh [GMSHarticle] is used to produce surface representations which inform subsequent electrostatic analysis.
-Python’s BEMPP API [bempp] is then used to perform boundary analysis.
-In short, this procedure enables a geometric model of our inflector to yield a full 3D electrostatic. Particles can then be tracked with high precision using the Boris method [boris].
+An initial estimate based on an analytical solution for a central track serves as a numerical starting point, after which the mesh generating software Gmsh [56] is used to produce surface representations which inform subsequent electrostatic analysis.
+Python’s BEMPP API [57] is then used to perform boundary analysis.
+In short, this procedure enables a geometric model of our inflector to yield a full 3D electrostatic. Particles can then be tracked with high precision using the Boris method [58].
 A major strength of this procedure and choice of software is speed, with a full simulation iteration taking only $\mathcal{O}$(hours).
-Prior work [weigelspiral] comparing this process to commercial software such as COMSOL demonstrates that the differences in field calculations are negligible.
+Prior work [59] comparing this process to commercial software such as COMSOL demonstrates that the differences in field calculations are negligible.
 5.4.3 Geometric Refinements
@@
 This is accomplished by the addition of electrostatic quadrupoles in the gap between the RFQ and inflector entrance (see also Fig.
-1.7).
-Using the output of the RFQ simulation [rfqsim] as input to the quadrupoles, geometry and potentials were optimized such that the tightest possible input beam is produced.
+8).
+Using the output of the RFQ simulation as input to the quadrupoles, geometry and potentials were optimized such that the tightest possible input beam is produced.
 The remaining three modifications: V-shape, plate angling, and the terminal wedge were motivated by work done by Barnard et al.
-[barnard_inflector].
+[60].
 These all have the effect of introducing quadrupole moments to the field of the inflector itself, which can provide mid-track focusing to minimize beam dispersion in flight.
-The V-shape introduces a concave v-shape to the anode and a convex v-shape to the cathode when viewed along the long axis of the electrode (Fig.3.15(a)).
-Plate angling reorients the plate angle relativce to each other along the course of the inflector.
-In other words the plates are parallel only in the center with a positive angle between the plates initially linearly transitioning to an equal and opposite negative angle by the end (Fig.3.15(b)).
+The V-shape introduces a concave v-shape to the anode and a convex v-shape to the cathode when viewed along the long axis of the electrode (see Fig.
+33(a)).
+Plate angling reorients the plate angle relative to each other along the course of the inflector.
+In other words the plates are parallel only in the center with a positive angle between the plates initially linearly transitioning to an equal and opposite negative angle by the end (see Fig.
+33(b)).
 The terminal gamma wedge cut modifies the end of the anode into a triangular rather than flat profile at the end and shortens the cathode.
 This assists with the fringe fields at the end which can produce vertical spreading.
-(Fig.3.15(c)).
+(see Fig.
+33(c)).
 (a) V-Shape Illustration
@@
 (c) Gamma Wedge Illustration
-Figure 3.15: Illustrations of different focusing concepts. From Ref. [barnard_inflector].
+Figure 33: Illustrations of different focusing concepts. From Ref. [60].
 These modifications and their associated focusing fields interplay, but can be independently varied alongside other parameters like electrode potential, electrode separation, etc.
@@
 As we are at the edge of practical optimization using Monte Carlo, this leads us to consider other potential alternatives.
-Surrogate models using a fully connected neural network capable of learning and interpolating underlying features have been previously explored[EdelenMachingLearning] [KoserMachineLearning] [villarrealMachineLearning].
+Surrogate models using a fully connected neural network capable of learning and interpolating underlying features have been previously explored [61, 62, 63].
 Initial iterations have proved promising and have informed the parameters for our current best inflector performance.
@@
 5.4.4 Characterization of Spiral Inflector
-The result of adding optimized focusing quadrupoles on the beam distribution entering the spiral inflector is illustrated in Fig. 3.16.
-Figure 3.16: Illustration of beam tightness at the entrance of the spiral inflector before and after the addition of focusing quadrupoles between the RFQ and inflector
-The optimized spiral inflector, with $\mathrm{H}_{2}^{+}$ tagged through, is illustrated in Fig.3.17.
-Figure 3.17: Illustrations of the optimized spiral inflector.
-Note the v-shape to the electrodes as well as the non-parallel angle seen in the bottom view, and the angled wedge visible in the top view.
+The result of adding optimized focusing quadrupoles on the beam distribution entering the spiral inflector is illustrated in Fig. 34.
+Figure 34: Illustration of beam tightness at the entrance of the spiral inflector before and after the addition of focusing quadrupoles between the RFQ and inflector.
+The optimized spiral inflector, with $\mathrm{H}_{2}^{+}$ tagged through, is illustrated in Fig. 35.
+Figure 35: CAD renderings of the optimized spiral inflector.
+Note the v-shape to the electrodes as well as the non-parallel angle seen in the bottom view with the apertures removed, and the angled wedge visible in the top view.
 The efficiency of this modified inflector is significantly improved.
@@
 This collimator is a circular aperture which, for this study, was chosen to be as large as possible while guaranteeing that less than one percent of $\mathrm{H}_{2}^{+}$ impact the electrodes.
-The resulting efficiencies and $\mathrm{H}_{2}^{+}$-electrode collisions are given in Table 3.3.
-Table 3.3: Comparison of efficiency and ion-electrode impact for the optimized vs traditional spiral inflector.
+The resulting efficiencies and $\mathrm{H}_{2}^{+}$-electrode collisions are given in Table 8.
+Table 8: Comparison of efficiency and ion-electrode impact for the optimized versus traditional spiral inflector.
 Efficiency Electrode Impact %
@@
 Optimized Inflector 97% 1%
-Traditional Inflector (Coll.) 39% $<1$%
-Optimized Inflector (Coll.) 93% $<1$%
+Traditional Inflector (Coll.) 39% <1<1%
+Optimized Inflector (Coll.) 93% <1<1%
 We also need to consider the impact of the new inflector on the quality of the beam which is successfully transmitted.
-These changes are characterized in Table 3.4.
+These changes are characterized in Table 9.
 Detailed histograms showing the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ at the entrance and exit of the inflector are shown in Fig.
-3.18.
-Table 3.4: Comparison of transmitted beam parameters at the inflector exit.
-$\sigma_{x}$ [mm] $\sigma_{y}$ [mm] $\sigma_{z}$ [mm] $\sigma_{r}$ [mm] $\sigma_{KE}$ [keV]
+36.
+Table 9: Comparison of transmitted beam parameters at the inflector exit.
+$\sigma_{x}$ ($m$) $\sigma_{y}$ ($m$) $\sigma_{z}$ ($m$) $\sigma_{r}$ ($m$) $\sigma_{KE}$ (keV)
 Traditional Inflector 3.0 3.9 10.1 2.3 4.0
 Optimized Inflector 4.9 2.7 6.4 2.4 3.1
-Figure 3.18: Illustration of the kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ prior to the inflector entrance vs at the inflector exit.
+Figure 36: Kinetic energy distribution of the $\mathrm{H}_{2}^{+}$ before the inflector entrance versus at the inflector exit.
 5.4.5 Orbit Matching and Beam Centering
-Figure 3.19: Beam tracking result from the AIMA central region study.
-Left beam centering and acceleration on top of a view of the electrodes.
+Figure 37: Beam tracking result from the AIMA central region study.
+Left: beam centering and acceleration on top of a view of the electrodes.
 Colors correspond to prticle phase.
 Right: 1 cm turn separation in the fourth turn with a single collimator.
-From [aima:central].
+From [64].
 Once the beam emerges from the spiral inflector, it must be properly matched in terms of orbit radius, radial momentum, and phase.
@@
 The spiral inflector end can be tilted (often this parameter is called $k^{\prime}$) to aid in the matching and thus becomes part of the optimization procedure.
-A preliminary design study of the IsoDAR central region was performed by the company AIMA in France [aima:central].
+A preliminary design study of the IsoDAR central region was performed by the company AIMA in France [64].
 The AIMA central region model is shown in Fig.
-3.19.
+37.
 This central region was designed with a slightly different magnetic pole shape and with its own matched spiral inflector, using the ouput beam from an early RFQ beam tracking and single electrostatic quadrupole transverse refocusing element.
@@
 For this preliminary design, we use a Gaussian beam with size and emittance informed by the AIMA study for further simulations of acceleration to 60 MeV/amu.
-Figure 3.20: Reduction of beam halo by collimators
 5.4.6 Low Energy Collimation
-Following injection, there is significant beam spread in the first few turns of the cyclotron that we need to manage.
+Figure 38: Placement of 16 collimators in the first seven turns. All collimators are $10\times 20$ $m$.
+Following the injection, there is a significant beam spread in the first few turns of the cyclotron that we need to manage.
 We do this by placing a series of collimators around the beam core in the early turns.
 We must do all of our collimation in the first 7 turns of the primary acceleration, corresponding approximately to a 2 MeV/amu limit, to avoid activating our collimators.
-The placement of the collimators is shown in Figure  3.21 The reduction in RMS spread is shown in Fig  3.41.
-In total, we have 15 collimators that remove 36.1% of the beam.
-The improvement in beam shape is shown in  3.20.
+The placement of the collimators is shown in Fig.
+38.
+The reduction in RMS is shown in Fig.
+59.
+In total, we have 16 collimators that remove 36.4% of the beam.
+The improvement in beam shape is shown in Fig.
+39.
 With time, the halo slowly redevelops but to a lesser extent than it would without collimation.
-Figure 3.21: Placement of 15 collimators in first 7 turns
+Figure 39: Snapshots of the bunch in radial-longitudinal coordinates (aka local frame of the bunch) demonstrating the reduction of beam halo by collimators.
 It is necessary to make the beam as circular as possible following the low energy collimators so that the beam is well-matched to “vortex motion.” Vortex motion stabilizes the beam and significantly reduces beam growth over time.
 It arises from the interaction between the self-field of the beam and the external magnetic field of the cyclotron.
-Vortex motion is discussed in depth in Ref [baumgarten:vortex2].
-3.4.7 Risks and Mitigation
+Vortex motion is discussed in depth in Ref.
+[5].
+5.5 Risks and Mitigation
 Risk: High voltage on the spiral inflector electrodes.
@@
 The vacuum enclosure comprises five parts: the upper and lower internal disk and 3 vacuum chambers.
-Fig 3.22 shows the vacuum enclosure in detail and the dimensions are in Table 3.5.
-The vacuum sealing is done using 2 metallic O-rings per interface (e.g., from Omniseal [PrecisionPolymerMetal]) except for the radial interface where Polymer O-rings could be used.
+Fig.
+40 shows the vacuum enclosure in detail and the dimensions are in Table 10.
+The vacuum sealing is done using 2 metallic O-rings per interface (e.g., from Omniseal [65]) except for the radial interface where Polymer O-rings could be used.
 The interface for the accessories will be done using a sleeve design and radial O-ring, described in Fig.
-3.22.
-Figure 3.22: Cyclotron vacuum chamber details
+40.
+Figure 40: Cross-section view of a CAD rendering of the HCHC-60 cyclotron, calling out vacuum chamber details.
+Table 10: Vacuum part properties
 Name Mass (Metric Ton) Dimensions (mm)
-Upper and lower internal disk 36 Ø4620*463
-Upper vacuum chamber 2.3 Ø4452*483
-Lower vacuum chamber - median plane 2.3 Ø4452*483
-Lower vacuum chamber - internal disk 1.8 Ø4386*380
-Table 3.5: Vacuum part properties
+Upper and lower internal disk 36 ⌀4620*463
+Upper vacuum chamber 2.3 ⌀4452*483
+Lower vacuum chamber - median plane 2.3 ⌀4452*483
+Lower vacuum chamber - internal disk 1.8 ⌀4386*380
 5.6.1 Risks and Mitigation
@@
 Its radial width remains at 250 mm.
-The total current used in the calculations is 180,850 amp-turns in each of the half coils.
+The total current used in the calculations is 180,850 amp-turns in each of the 180 half coils.
 The apparent current density is thus 3.01 A/mm2, i.e.
@@
 If they are added, an increase in main coil current will be required
-Figure 3.23: Left: Schematic view of a cross-section of the main coil, with six pancakes per half coil.
-Right: Possible configuration of the winding, using a square conductor with a round cooling channel in its center.
+Figure 41: Left: Schematic view of a cross-section of the main coil, with six pancakes per half coil. The parameters of these coils are given in Table 11.
 In Fig.
-3.23 a schematic view of the coil is given.
-Also, a possible configuration of winding and cooling is calculated.
+41 a schematic view of the coil is given.
+We show a possible configuration of winding and cooling in Table 11.
 Although this configuration does not exactly sum to a cross-section of 240 by 250 mm2, it shows that a cooling configuration is feasible.
@@
 The increase in height from 200 to 240 mm is the equivalent of adding an extra pancake.
+Table 11: Proposed coil configuration for a single pancake. The HCHC-60 will require 12 pancakes.
+Parameter Value
+# of axial turns 2
+# of radial turns 12
+Wire size $20.5\times 20.5$ mm2
+Cooling hole diameter 15 mm
+Pancake width 251 mm
+Copper weight 708 kg
+Current $1256\times 24$ Ampere-turns
+Power 45 kW
+Water flow 21.5 L/min
+Pressure drop 12.3 kg/cm2
 5.8 Cyclotron RF System
@@
   • The Low Level RF system (LLRF) that controls and regulates the RF amplitude on the dee and drives the tuning system.
-Table 3.6: Details of the RF design.
+Table 12: Important parameters of the RF cavities. These are per cavity. The HCHC-60 will require four RF cavities.
 RF System Component Design Value
@@
 Number of dee stems 4 per dee (2 up, 2 down)
-Number of Dee’s 4
+Number of dees 4
 Acceleration harmonic 4th
@@
 Cavity Q-factor 9620
-We consider these in the Sections below. The numerical parameters of the RF system are listed in Table 3.6.
+We consider these in the Sections below. The numerical parameters of the RF system are listed in Table 12.
 5.9 RF Cavities (Dees, Stems and Liners)
-The RF cavities of the cyclotron have been designed using CST microwave studio [cst:microwave_studio] in order to define the shape of the Dee and the Dee stems.
+The RF cavities of the cyclotron have been designed using CST microwave studio [66] in order to define the shape of the dee and the dee stems.
 We optimized our designs to reach the right resonance frequency, satisfy the dee voltage Law, and minimize power dissipation.
 Ultimately, this improves the quality factor (Q-factor).
-Figure 3.24: Section view of the RF cavity.
-Figure 3.25: Outer view of the RF cavity.
-The Dees and cavity walls are all made of oxygen-free high-conductivity (OFHC) copper, which has good electrical conductivity and good solderability.
+Figure 42: Section view of the HCHC-60 double-stem, double-gap RF cavity, calling out the most important internal features.
+Figure 43: CAD rendering of the HCHC-60 double-stem, double-gap RF cavity, calling out the most important external features.
+The dees and cavity walls are all made of OFHC copper, which has good electrical conductivity and good solderability.
 The dees are fabricated out of a 20 mm thick copper plate and are made rigid by a supporting arm.
 The four round stems support the dee and provide the right inductance in order to produce the required resonance frequency.
-Their size and location produce the desired Dee voltage vs.
-radius profile.
+Their size and location produce the desired Dee voltage versus radius profile.
 Each dee will be equipped with vacuum pumping panels to efficiently pump the median plane.
@@
 This fact has not been taken into account in detail in the RF simulations to date, but some room for housing the deflector has been provisioned at the back of the dee.
-Figures 3.24 and 3.25 present the main elements composing the RF cavities.
+Figures 42 and 43 present the main elements composing the RF cavities.
 5.9.1 Integration of preliminary central region in RF Cavities
-In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA [aima:central] into the RF cavity modelling, several technical choices were made to ensure the effective implementation of this innovative configuration:
-  • The central region design was trimmed at 150 mm from centre.
-  • Both the Dee and counter dee original central region was removed by trimming the cavities at 250 mm from centre.
+In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA [64] into the RF cavity modeling, several technical choices were made to ensure the effective implementation of this innovative configuration:
+  • The central region design was trimmed at 150 mm from center.
+  • Both the dee and dummy-dee original central region was removed by trimming the cavities at 250 mm from center.
   • Sections were cut straight, and a linear transition (loft) technique was used to achieve the connections between the trimmed components.
-The figure 3.26 shows the final integration of the central region.
-The lofted transitions can be better seen from top at figure 3.27.
+Figure 44 shows the final integration of the central region.
+The lofted transitions can be better seen from the top view in Fig.
+45.
 They leads to an acceptable linear vertical gap opening from 26 mm at a radius of 150 mm, to 50 mm at a radius of 250 mm.
-The section view of the model, presented at figure 3.28, shows the radial evolution of the vertical gap.
+The section view of the model, presented in Fig.
+46, shows the radial evolution of the vertical gap.
 In addition, adequate clearance provisions were incorporated, specifically to fit the inflector and optics coil.
@@
 Power bridges between dees in the central regions were removed since we recommend to power the 4 cavities independently.
-Figure 3.26: dee at central region close view
-Figure 3.27: dee lofted transition top view
-Figure 3.28: vertical gap evolution close view
+Figure 44: CAD rendering of the dee–close-up of the central region, showing the spiral inflector housing and field-shaping posts.
+Figure 45: CAD rendering of the dee–top view, showing the spiral inflector housing and field-shaping posts.
+Figure 46: CAD rendering of the dee–side view, showing the radial evolution of the gap.
 5.9.2 RFQ Integration in RF Cavities
-Some clearance provision has been set for the integration of the RFQ, as seen in figure 3.29
+Some clearance provision has been set for the integration of the RFQ, as seen in Fig. 47:
   • Diameter 280 mm
@@
 We have observed that the RFQ clearance has a low impact on cavity frequency and performances, and the output of the RFQ could be lowered more without impacting the cavity behavior.
-Figure 3.29: Clearance for RFQ
+Figure 47: CAD rendering of the dee, showing the clearance for RFQ.
 5.9.3 RFQ getter pumps in RF Cavities
 Incorporating getter pumps into the dees allows us to improve our vacuum quality.
-By integrating three housing for SAES UHV1400 getter pumps [ModulesUHV] per dee, we aim to ensure highly efficient and reliable vacuum performance within the Cyclotron system.
+By integrating three housing for SAES UHV1400 getter pumps [67] per dee, we aim to ensure highly efficient and reliable vacuum performance within the cyclotron system.
 The use of six of those getter pumps per cavity facilitates the effective removal of residual gases, since achieving and maintaining ultra-high vacuum conditions essential for optimal particle acceleration.
-The figure 3.30 shows acceptable locations for three getter pump housings.
+Figure 48 shows acceptable locations for three getter pump housings.
 Furthermore, our preliminary assessments suggest that it may be feasible to scale up the configuration to include up to two sets of five getter pumps per cavity.
 The very low electrical field of the locations of the housings will allow the getter pumps to function safely.
-Figure 3.30: Provisional getter housings
+Figure 48: CAD rendering of the deer, showing the provisional getter housings (rectangular recesses).
 This early integration of the housings of the getters in the dees is mandatory to assess their impact on the cavity frequency, and on surfaces currents.
@@
 5.9.4 RF Cavities design optimization
-3.8.4.1 Performance requirement target
+Performance requirement target
 We have optimized the cavity geometry to ensure the following requirements:
@@
   • Reference accelerating voltages: 65 kV at 80 mm (capture), 230 kV at 2000 mm (extraction)
-The initial target value for the outer radius voltage was initially set at 250 kV, but according to our expertise, such high voltage will be very difficult to stand in cyclotron accelerator gaps of this size.
-It has been decided to reduce the target to 230 kV which remains a challenging value.
-3.8.4.2 Parametric optimization of the design
+The initial target value for the outer radius voltage was initially set at 250 kV, but from extensive prior experience we expect such high voltage will be very difficult to maintain in cyclotron accelerator gaps of this size.
+It has been decided to reduce the target to 230 kV which remains a challenging, albeit manageable, value.
+Parametric optimization of the design
 We have intensively used parameters while building the CST model, to allow fast design exploration.
 We have retained four parameters to play with, for optimizing the cavity behavior.
-The parameters, shown at figure 3.31, are defined as a deviation from the early cavity design values.
+The parameters, shown in Fig.
+49, are defined as a deviation from the early cavity design values.
   • A: 1st pillar radial position
@@
   • D: Deflector height
-Figure 3.31: design optimization parameters
+Figure 49: CAD rendering of the dee, overlayed with an indication of the design optimization parameters.
 We did some initial exploration of the parameter space, but we did the final optimization using the built-in CST optimizer.
-Table 3.7: RF cavities optimization goals definition and achieved values.
+Table 13: RF cavity optimization goals definition and achieved values.
 Goal description Target Final Value
 Ratio between voltage at 80 mm and 2000 mm 0.28 0.283
-Mode1 frequency [MHz] 32.8 32.805
-We selected the trust region algorithm and have defined scaled goals that would ensure convergence toward the required frequency and voltages.
-The trust region algorithm converged in a few tens of iterations.
-The definition of the optimization goals, and their target and reached values are contained in table 3.7.
-The table 3.8 shows the optimal values reached by the optimizer.
-Table 3.8: RF cavities optimized design parameters.
-Design parameter Value [mm]
+Mode 1 frequency 32.8 MHz 32.805 MHz
+Table 14: RF cavities optimized design parameters.
+Design parameter Value (mm)
 First pillar radial position (A) 444.6
@@
 Total cavity height (D) 1412
+We selected the trust region algorithm and have defined scaled goals that would ensure convergence toward the required frequency and voltages.
+The trust region algorithm converged in a few tens of iterations.
+The definition of the optimization goals, and their target and reached values are contained in Table 13.
+Table 14 shows the optimal values reached by the optimizer.
 5.9.5 RF cavity computation results
 The RF cavity computation employing the eigen mode solver on the optimized geometry yielded the following performance and characteristics of the cyclotron system:
-  • Frequency: The RF cavity operates at a frequency of 32.805 MHz, ensuring compatibility with the designed acceleration, and the RFQ integration,
-  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.283, indicating that computed voltages at center would be 65.2 kV for a chosen voltage of 230 kV at the outer radius (2m),
-  • CST computed quality factor (Q): 10586,
-  • Estimated quality factor (Q): 7281,
-  • Stored Energy (for max RF of 230 kV at 2 m): 4.06 Joules
-  • CST Computed Losses: The computed losses per cavity using CST simulations are 79 kW,
+  • Frequency: The RF cavity operates at a frequency of 32.805 MHz, ensuring compatibility with the designed acceleration and the RFQ integration.
+  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.283, indicating that computed voltages at center would be 65.2 kV for a chosen voltage of 230 kV at the outer radius (2 m).
+  • CST computed quality factor (Q): 10586
+  • Estimated quality factor (Q): 7281
+  • Stored Energy (for max RF of 230 kV at 2 m): 4.06 J
+  • CST Computed Losses: The computed losses per cavity using CST simulations are 79 kW.
   • Estimated Power Dissipation: By experience, senior RF designers at IBA state that the CST eigenmode simulations underestimate the actual losses by up to 30%.
@@
   • Maximal Surface Current: The RMS surface current reaches a maximum of 17052.6 A/m in some reduced locations.
-Which leads to a maximum surface power of 43.5 W/cm².
+Which leads to a maximum surface power of 43.5 W/cm2.
 This maximal surface power is localized on the dee edges.
 Careful attention will be required for thermal management in these regions.
-For the rest of the surfaces, the thermal dissipation is well below 10 W/cm².
-See figure 3.36 and 3.37 While matching the requirement, these computation results provide valuable information for next steps of the Cyclotron’s RF cavity design including consideration in managing thermal effects.
-Some illustrations of the electrical and magnetic fields computed by CST are given for reference at figures 3.32, 3.33, 3.34 and 3.35.
-Figure 3.32: E field Vertical cross section
-Figure 3.33: E-field Horizontal cross section
-Figure 3.34: H-field horizontal cross section
-Figure 3.35: H-field vertical cross section
-Figure 3.36: Surface currents
-3.8.5.1 Detailed power analysis
-We will now detail the post treatment of the CST results that lead to the results presented in Sec.
-3.8.5.
+For the rest of the surfaces, the thermal dissipation is well below 10 W/cm2.
+See Figs.
+54 and 55.
+While matching the requirement, these computation results provide valuable information for next steps of the cyclotron’s RF cavity design including consideration in managing thermal effects.
+Some illustrations of the electrical and magnetic fields computed by CST are given for reference in Figs. 50, 51, 52 and 53.
+Figure 50: RF cavity: E-field vertical cross section, calculated with CST.
+Figure 51: RF cavity: E-field horizontal cross section, calculated with CST.
+Figure 52: RF cavity: H-field horizontal cross section, calculated with CST.
+Figure 53: RF cavity: H-field vertical cross section, calculated with CST.
+Figure 54: RF cavity: Surface currents, calculated with CST.
+Detailed power analysis
+We will now detail the post treatment of the CST results that lead to the results presented in Section 5.9.5.
 The CST eigenmode solver results are computed for a given reference energy in the system (about 1 J).
@@
 In our case, the reference location, is the orbit at 2 m, and the desired voltage is 230 kV.
-Figure 3.37: Surface current generating more than 10 W/cm2
+Figure 55: RF cavity: Surface current generating more than 10 W/cm2
 The ratio of the desired voltage to the reference voltage at 2 m is 2.013.
 We can scale the currents and voltages using this factor.
-We can also scale the energies, power and losses using this factor squared (formulae 3.1, 3.2 and 3.3).
+We can also scale the energies, power and losses using this factor squared (formulae 1, 2 and 3).
 We thus have:
-$E_{stored}=E_{stored,ref}\frac{V_{target}^{2}}{V_{ref}^{2}}$ (3.1)
-$P_{loss}=P_{loss,ref}{V_{target}^{2}\over V_{ref}^{2}}$ (3.2)
-$I{{}_{s}}=I_{s,ref}{V_{target}\over V_{ref}}$ (3.3)
-Table 3.9: Optimized cavities, functional characteristics.
+$E_{stored}=E_{stored,ref}\frac{V_{target}^{2}}{V_{ref}^{2}}$ (1)
+$P_{loss}=P_{loss,ref}{V_{target}^{2}\over V_{ref}^{2}}$ (2)
+$I{{}_{s}}=I_{s,ref}{V_{target}\over V_{ref}}$ (3)
+Here, $E_{stored}$ and $E_{stored,ref}$ are the actual stored anergy and the reference stored energy from the 1 J CST solution.
+$P_{loss}$ and $P_{loss,ref}$ are the actual power loss and the reference power loss from the 1 J CST solution.
+$V_{target}$ is the desired cavity accelerating voltage and $V_{ref}$ is the voltage obtained from the 1 J CST solution.
+Table 15: Optimized cavities, functional characteristics.
 CST reference Scaled value
-Stored energy [J] 1 4.06
-Voltage at on the 2m orbit kV] 114.2 230
-Voltage at center [kV] 32.4 65.2
-Maximal voltage at outer edge of the dee [kV] 124.6 250.9
-Computed losses [kW] 19.48 78.9
-Maximal surface current [A/m] 8470 17052.6
-[…]
-$R_{s}={1\over\sigma.\delta}$ (3.4)
-$\delta=\sqrt{2\over{\sigma.\mu.\omega}}$ (3.5)
-$P_{s}=R_{s}I_{s}^{2}$ (3.6)
-Table 3.10: Cavity properties for computing surface dissipation
+Stored energy (J) 1 4.06
+Voltage at on the 2m orbit (kV) 114.2 230
+Voltage at center (kV) 32.4 65.2
+Maximal voltage at outer edge of the dee (kV) 124.6 250.9
+Computed losses (kW) 19.48 78.9
+Maximal surface current (A/m) 8470 17052.6
+CST provides the maximal surface current $I_{s}$ in the cavity which isn’t a directly useful value, but we can compute from it the maximal surface power dissipation $P_{s}$, which will be useful for designing the cooling system sizing and layout.
+The physical properties needed for this determination are included in Table 16.
+We first need to compute the surface resistance $R_{s}$ whose formula, 4, uses the conductivity $\sigma$ and the RF skin depth $\delta$ (see formulae 4, 5 and 6).
+we end up with a maximal surface dissipation Power $43.5$ W/cm2.
+This value, despite being high, only concerns the very small surface of the dee edges, and should not be considered as an issue for the design of the cooling system.
+$R_{s}={1\over\sigma\cdot\delta}$ (4)
+$\delta=\sqrt{2\over{\sigma\cdot\mu\cdot\omega}}$ (5)
+$P_{s}=R_{s}I_{s}^{2}$ (6)
+Here, $\omega=2\pi f$ is the angular frequency and $\mu=\mu_{r}\cdot\mu_{0}$ the product of relative permeability and vacuum permeability.
+Table 16: Cavity properties for computing surface dissipation
 Property Value
@@
 Angular Frequency $\omega$ $2\pi\times 32.805\times 10^{6}$ rad/s
-Skin Depth $\delta$ $11.5$ µm
-3.8.5.2 Radial voltage dependency
+[…]
+Radial voltage dependency
 The voltages at different radii have been calculated by integrating the electrical field along circular arcs crossing the acceleration gaps, and reference curves crossing central gap and the dee outer edge gap.
@@
 The computed voltages have then been scaled to match a target voltage of 230 kV at 2 m.
-The resulting absolute and normalized voltage radial dependency are given in figure 3.38 and table 3.11
-Figure 3.38: Dee voltage radial dependency
-Table 3.11: RF voltage radial dependency.
-Radius [mm] Voltage [kV]
+The resulting absolute and normalized voltage radial dependency are given in Fig.
+56 and Table 17.
+Figure 56: Radial dependency of the potential difference (“accelerating voltage”) across the accelerating gaps.
+Table 17: RF voltage radial dependency.
+Radius (mm) Voltage (kV)
 Capture (80) 65.17
@@
 We plan to use vacuum tubes (triodes or tetrodes) to produce the RF power.
-We have considered using LDMOS amplifiers given the recent improvements in power density, reliability, and cost.
-However, we believe that vacuum tubes will perform better in the harsh environment and will be able to better accommodate the rapid variations in load due to the beam temporal patterns.
+We have considered using LDMOS (laterally diffused metal-oxide semiconductor) amplifiers given the recent improvements in power density, reliability, and cost.
+However, considering the harsh environment in question, vacuum tubes will perform more reliably and will also be better able to accommodate the rapid variations in load due to the beam temporal patterns.
 5.9.8 RF Lines and Couplers
@@
 Both of these systems are capable of providing the necessary 300 kW power to each cavity.
-Capacitive coupling is less prone to multipacting because the coupling antenna could be easily DC biased.
+Capacitive coupling is less prone to multipactoring because the coupling antenna could be easily DC biased.
 But, our expertise in high power coupler design advocates for an inductive solution.
 The design of inductive coupler is easier to water-cool, and there is a wide area where H-field is high enough for efficient coupling behind the second pillar of the cavity.
-We have initiated a preliminary design computation of such an inductance coupler, feeding the RF power from a EIA 6”1/8 port.
-The suggested design is a loop made out of a 20x60 mm rectangular section bar, offering room for water-cooling.
-See figure 3.39.
+We have initiated a preliminary design computation of such an inductance coupler, feeding the RF power from a EIA 6 1/8” port.
+The suggested design is a loop made out of a $20\times 60$ mm rectangular section bar, offering room for water-cooling.
+See Fig.
+57.
 The adaptation tuning will be performed by rotation of the loop around the vertical axis.
-The design has been optimized with a default angle of 45° to the cavity axis, in order to allow a maximal tuning range.
-The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 235 mm, which provides a good matching (figure 3.40) and a $S_{1,1}$ computed scattering parameter inferior to -60 dB.
-Figure 3.39: RF loop coupler design
-Figure 3.40: Coupler port impedance
+The design has been optimized with a default angle of 45 to the cavity axis, in order to allow a maximal tuning range.
+The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 235 mm, which provides a good matching (see Fig.
+58) and a $S_{11}$ computed scattering parameter smaller than -60 dB.
+Figure 57: CAD rendering of the RF cavity showing the position and size of the loop coupler.
+Figure 58: RF loop coupler reflected power ($S_{11}$) showing excellent coupling at -60 dB.
 5.9.9 LLRF Electronics
@@
 It is necessary to perform an in-depth particle simulation of the primary acceleration of the beam to ensure that the beam reaches energy, maintains quality, and is possible to extract.
-We do this using OPAL  [adelmann_opal_2019] which is a particle accelerator simulation suite developed at the Paul Scherrer Institute that accounts for space-charge effects.
-It is based on particle-in-cell (PIC) calculations that track a ’bunch’ of particles through time-varying magnetic and electric fields.
-The temporal integration within OPAL is carried out using a fourth-order Runge-Kutta method.There is a mode within OPAL, OPAL-cycl, which is designed to simulate cyclotrons.
+We do this using OPAL  [55] which is a particle accelerator simulation suite developed at the Paul Scherrer Institute that accounts for space-charge effects.
+It is based on particle-in-cell (PIC) calculations that track a “bunch” of particles through time-varying magnetic and electric fields.
+The temporal integration within OPAL is carried out using a fourth-order Runge-Kutta method.
+There is a mode within OPAL, OPAL-cycl, which is designed to simulate cyclotrons.
 This makes it well-suited for the primary acceleration phase of the IsoDAR cyclotron.
-Conveniently, OPAL-cycl has built-in objects called PROBEs, that can provide beam diagnostics, and and COLLIMATORs, which can provide halo reduction.
+Conveniently, OPAL-cycl has built-in objects called probes, that can provide beam diagnostics, and collimators, which can provide halo reduction.
 5.10.1 Fieldmaps
 A critical aspect of these simulations is that they accurately simulate the electric and magnetic fields within the cyclotron.
-We performed these simulations using a 1D RF model derived from the voltages provided in Table 3.11.
+We performed these simulations using the latest updates to our 1D RF model discussed previously.
+Specifically, the field is derived from the voltages provided in Table 17.
 The radially increasing voltage leads to a roughly constant turn separation, which is important for ease of extraction.
@@
 The overlap between beams provides a lower bound for the power that will be deposited on the electrostatic septum, so we aim to minimize this value.
-The data from the probe is displayed in Fig 3.42.
+The data from the probe is displayed in Fig.
+60.
 5.10.3 Simulation Results
 Using the field maps as discussed earlier and the set of 15 collimators as described in the previous sections.
-We are able to smoothly accelerate the beam up to 60 MeV/amu as shown in Fig 3.43.
-Additionally, there is minimal growth in beam size in the collimated beam as shown in Fig 3.41.
-Additionally, a probe at an azimuth of 25 degrees sets a lower limit for extraction at 120W as seen in Fig 3.42.
-Figure 3.41: The collimated beam had a much reduced radial RMS spread.
-Figure 3.42: There are two potential extraction points
-Figure 3.43: There is a smooth acceleration of the beam
+We are able to smoothly accelerate the beam up to 60 MeV/amu as shown in Fig.
+61.
+Additionally, there is minimal growth in beam size in the collimated beam as shown in Fig.
+59.
+Additionally, a probe at an azimuth of 25 sets a lower limit for extraction at 127 W as seen in Fig.
+60.
+Figure 59: Rms beam size in the radial direction as a function of beam energy. We observe a much reduced radial spread in the collimated beam.
+Figure 60: Beam power versus radius on a radial probe inserted at 25 azimuth, binned in 0.5 mm wide bins to account for the thickness of the electrostatic septum.
+The final two turns exhibit low enough inter-turn power to extract the beam.
+[…]
 5.10.4 Electrostatic Extraction Channels
@@
 This generates a well-shielded electric field in the region between the puller and septum.
-A rendering of the electrostatic septum is shown in Fig 3.44.There are two such electrostatic channels that increase separation between the last two turns.
-Figure 3.44: CAD rendering of electrostatic septum. From Ref. [winklehner_new_2022]
+A rendering of the electrostatic septum is shown in Fig.
+62.
+Two such electrostatic channels increase separation between the last two turns.
+Figure 62: CAD rendering of electrostatic septum. Inventor model (left). Visualization of electric field from COMSOL (right). From Ref. [6].
 We determine the exact location of each element of the extraction channel iteratively.
 The first step is to run a Python script on a pre-existing OPAL simulation of the last three turns that determines the initial positions of the septum and the puller.
-We choose to place the septum evenly between the last two turns.
-A .txt file is generated that allows for a model of the septum to be built in Autodesk Inventor.
+We chose to place the septum evenly between the last two turns.
+A text file is generated that allows for a model of the septum to be built in Autodesk Inventor.
 With this geometry, we use COMSOL to determine the electric field and add it to the pre-existing RF field of the cyclotron in OPAL.
@@
 This process is repeated twice at which point further iteration does not result in any improvements in ESC performance.
-Following this we were able to reduce our losses on the first septum to 205 W which is 2.5% greater than our 200 W goal and we lose 21 W on the second septum.
-In the future we plan to reintroduce a resonance into the magnetic field which will increase the inter-turn separation at extraction which will greatly reduce this number.
-The end result of the septum optimization process is shown in Fig 3.45.
-Figure 3.45: This is the tracking of 100 particles as they pass through the electrostatic channels (shown in green). It also shows the location of our magnetic extraction channel in grey.
+Following this we were able to reduce our losses on the first septum to 158 W which is below our 200 W goal and we lose 7 W on the second septum.
+In the future, we are considering reintroducing a resonance into the magnetic field which would increase the inter-turn separation at extraction and greatly reduce this number.
+The end result of the septum optimization process is shown in Fig.
+63.
+Figure 63: 100 randomly selected particle trajectories out of the simulation sample of 200,000 particles as they pass through the electrostatic channels (shown in green).
+Also shown is the location of the first magnetic extraction element in gray.
 5.10.5 Magnetic Extraction Channels
@@
 These bars allow us to locally reshape the magnetic field and are very compact.
-We model our magnetic extraction off of the MSU cyclotron [MSU_Cyclotron] which generates a strong radially increasing magnetic field that serves to refocus the beam after the ESC.
+We model our magnetic extraction off of the MSU cyclotron [68] which generates a strong radially increasing magnetic field that serves to refocus the beam after the ESC.
 By using the gradients that are produced at MSU as a reference point, we conservatively set a gradient of 1.5 kG/cm which is much less than the 5 kG/cm that they were able to achieve.
-We directly modified our magnetic field to contain these elements.
-It is visible in Fig 3.46 that a fairly week magnetic channel is sufficient to maintain beam quality.
-We do not anticipate difficulties with these channels because the electrostatic channels generate the majority of the turn separation.
-Figure 3.46: A 1.5 kG/cm gradient after the electrostatic channel dramatically reduces radial beam spread
+We directly modified our magnetic field to mimic these elements.
+It is visible in Fig.
+64 that a single relatively weak magnetic channel can improve beam quality.
+We do not anticipate difficulties implementing a series of such channels to guide and focus our extracted beam because we have significantly increased turn-to-turn separation following the electrostatic channels.
+Figure 64: A 1.5 kG/cm gradient after the electrostatic channel dramatically reduces radial beam spread
 5.10.6 Risks and Mitigation
@@
 Mitigation: Understanding beam dynamics sufficiently well that this does not happen, and having adequate beam diagnostics to be able to tune the beam for maximum turn separation.
-Further, we could place a narrow carbon foil in front of teh septum, dissociating the few $\mathrm{H}_{2}^{+}$ ions that would otherwise hit the septum.
+Further, we could place a narrow carbon foil in front of the septum, dissociating the few $\mathrm{H}_{2}^{+}$ ions that would otherwise hit the septum.
 The resulting protons have lower magnetic rigidity and bend away from the septum before hitting it.
@@
 Mitigation: Adequate beam dynamics simulations, including tracing of particle orbits.
-Chapter 4 Installation in the Yemilab Setting
+6 Installation in the Yemilab Setting
 The most likely deployment location for IsoDAR is at the Yemilab underground laboratory, 1 km below the surface adjacent to the Handuk iron mine in the Republic of Korea.
 Fig.
-4.1 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 meters at a 15% slope from a surface access point into the mine, or by an elevator which descends down a 600 meter vertical shaft.
+65 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 m at a 15% slope from a surface access point into the mine, or by an elevator which descends down a 600 m vertical shaft.
 Unfortunately, the elevator is suitable only for personnel access.
@@
 Fig.
-4.2 details the area where the IsoDAR experiment would be installed, with the target only a few meters from the planned 3 kiloton LSC (Liquid Scintillator Counter).
+66 details the area where the IsoDAR experiment would be installed, with the target only a few meters from the planned 3 kt LSC (Liquid Scintillator Counter).
 The following section provides a general description of deployment strategies in an underground setting, pertaining only to the cyclotron.
@@
 A similar approach will be taken for the target system in Volume II of this PDR.
-Figure 4.1: (a) Schematic of Handuk/Yemilab area, (b) Low resolution laser scan of Handuk mine and Yemilab access ramps, green area denotes high resolution scans of IsoDAR deployment area, shown in more detail in the next figure.
+Figure 65: (a) Schematic of Handuk/Yemilab area, (b) Low resolution laser scan of Handuk mine and Yemilab access ramps, green area denotes high resolution scans of IsoDAR deployment area, shown in more detail in the next figure.
 From Ref.
-[kimYemilabNewUnderground2024].
-Figure 4.2: Layout of Yemilab, showing (in green) the area where IsoDAR will be installed. The cyclotron will be assembled in the dedicated cyclotron vault as shown.
+[9].
+Figure 66: Layout of Yemilab, showing (in green) the area where IsoDAR will be installed. The cyclotron will be assembled in the dedicated cyclotron vault as shown.
 6.1 Transport of Components
@@
 The preferred method is to use trucks along the highways.
-Prior experience of IBA with rail transport has been that it is difficult to provide adequate shock protection, and heavy pieces can be damaged in transit.
+IBA has had prior experience with rail transport and has determined that it is difficult to provide adequate shock protection, and heavy pieces can be damaged in transit.
 As stated above, transport from the surface to the experimental site must be accomplished by way of the mine ramp depicted in Fig.
-4.1.
-Nominally, its cross section is 5 x 5 meters, however we are told that pinch points and sharp bends may occur and might provide challenging areas to transport large, heavy loads through.
-We have received assurance that large mining trucks, 3.5 meters wide, can traverse the entire ramp, as this is the way ore was transported to the surface for many years.
+65.
+Nominally, its cross section is $5\times 5$ m, however we are told that pinch points and sharp bends may occur and might provide challenging areas to transport large, heavy loads through.
+We have received assurance that large mining trucks, 3.5 m wide, can traverse the entire ramp, as this is the way ore was transported to the surface for many years.
 While high-resolution laser scans of the IsoDAR and LSC areas have been performed, (cf.
-Figures 4.8 and 4.9), the resolution of the laser scans of the access ramp areas are not of sufficient resolution to provide enough detailed information about the location or size of the possible constrictions.
+Figs.
+72 and 73), the resolution of the laser scans of the access ramp areas are not of sufficient resolution to provide enough detailed information about the location or size of the possible constrictions.
 Improved scans must be performed as part of final installation planning in the Technical Design Report phase (TDR).
@@
 The plan is to use an SPMT vehicle (Self-Propelled Modular Transporter) shown in Fig.
-4.3 for bringing the magnet and vacuum liner pieces from the surface.
-These electrically-powered vehicles are flexible in their size, the sets of 4 tires are independently steerable, so the vehicle is extremely maneuverable, and can easily accommodate the heaviest piece of magnet steel [SPMT].
-Figure 4.3: A typical SPMT vehicle, a versatile way to transport heavy loads over irregular terrain.
+67 for bringing the magnet pieces from the surface.
+These electrically-powered vehicles are flexible in their size, comprising sets of 4 tires, each set independently steerable, the vehicle is extremely maneuverable and can easily accommodate the heaviest piece of magnet steel [69].
+The vacuum liner is made mostly of stainless steel pieces, but though large and bulky, the pieces can be brought down and welded in the underground assembly cavern.
+Figure 67: A typical SPMT vehicle, a versatile way to transport heavy loads over irregular terrain.
 The only challenge will be the magnet coils.
-Each of the two coil packages measures 5 meters across and 0.3 meters high, the individual coils inside measure 4.960 x 0.252 meters.
-Figs.
-4.5 and 4.5 gives an approximate view of a coil package and the clearances in the access ramp to the Laboratory (marked as the ”Entrance Tunnel” in Fig.
-4.1 (a).
-This particular ramp is well-constructed and engineered, and has a measured width of 5 meters.
+Each of the two coil packages measures 5 meters across and 0.3 m high, the individual coils inside measure $4.960\times 0.252$ m.
+Figures 68 and 69 give an approximate view of a coil package and the clearances in the access ramp to the Laboratory (marked as the ”Entrance Tunnel” in Fig.
+65 (a).
+This particular ramp is well-constructed and engineered, and has a measured width of 5 m.
 A special mounting fixture capable of holding the coil packages at an angle should be able to bring these coils through this particular passageway.
@@
 In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections.
-This has already been studied in Chapter 3 of the CDR document [alonso_isodaryemilab_2021, alonso_isodaryemilab_2022].
+This has already been studied in chapter 3 of the CDR document [70, 1].
 The heavy winding machine and fixtures would need to be transported, but could be broken down into parts that will fit through the ramps, then assembled on site.
@@
 There is precedence for sectioned coils.
-For the large TRIUMF cyclotron [TRIUMF], coils are separated into six different segments as the outer diameter of the coils is almost 20 meters, making the transport of fully assembled coils essentially impossible over conventional roadways.
+For the large TRIUMF cyclotron [71], coils are separated into six different segments as the outer diameter of the coils is almost 20 m, making the transport of fully assembled coils essentially impossible over conventional roadways.
 The inconvenience for a segmented coil, is that the ends of every turn of every section must be spliced, substantially increasing the cost and decreasing the efficiency of operation, as the coil must have as few turns as possible with very high current flowing through each turn.
-Figure 4.4: Isometric view of the coil package, on a specially-designed cart, transporting the coil through the access ramp to the Yemilab campus.
-Figure 4.5: This ramp is engineered to a 5-meter width, and maximum height of 5 meters.
-Unfortunately we do not have sufficient information about the winding access ramp from the surface to the base of the shaft to assess at this time whether transport of the coils will be possible with this conveyance device.
+Figure 68: Isometric view of the coil package, on a specially-designed cart, transporting the coil through the access ramp to the Yemilab campus.
+Figure 69: Front view of the coil package in the access tunnel. The ramp is nominally engineered to a 5 m width, and maximum height of 5 m. Dimensions are in mm and degrees.
 6.2 Assembly underground
-The cyclotron will be assembled in a hall measuring 12 meters wide, 17 meters deep and 10 meters high.
-This hall was excavated to our specifications at the 90 degree bend in the access ramp to the LSC, as seen in Fig.
-4.2.
-A flat area, roughly 10 m x 5 m at the corner will be the staging area for components transported from the surface.
+The cyclotron will be assembled in a hall measuring 12 m wide, 17 m deep and 10 m high.
+This hall was excavated to our specifications at the 90 bend in the access ramp to the LSC, as seen in Fig.
+66.
+A flat area, roughly $10\times 5$ m at the corner will be the staging area for components transported from the surface.
 A bridge crane riding on tracks will be used to rig parts to the cyclotron assembly point.
 Conversations with a crane manufacturer, Industrias Electromecánicas GH, S.A.
-[GHcranes] gave us assurances that a 40-ton capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig.
-4.6.
+[72] gave us assurances that a 40 t capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig.
+70.
 It could be configured to fit the dimensions of the cavern, and broken into modular parts that could be transported to the site.
-Figure 4.6: Examples of GH gantry cranes, capacities: foreground 20 tons, background 260 tons. From Ref. [GHcranes].
-Figure 4.7: Cyclotron and crane on cavern floor.
-Figure 4.8: Cyclotron and crane inside the cavern.
-An overview of the cyclotron inside the cavern with the rigging equipment is shown in Figures 4.8 - 4.9.
+Figure 70: Examples of GH gantry cranes, capacities: foreground 20 t, background 260 tons. From Ref. [72].
+Figure 71: Cyclotron and crane on cavern floor.
+Figure 72: Cyclotron and crane inside the cavern.
+An overview of the cyclotron inside the cavern with the rigging equipment is shown in Figures 72 - 73.
 It should be noted that the gantry crane will be dismantled once the heavy parts are rigged.
@@
 In addition, the large crane can be reconfigured and reassembled in the target hall, and used for assembly of the MEBT, target and shielding, and for subsequent maintenance of the target systems.
-Figure 4.9: Cyclotron open and crane with cavern mesh - cut-view.
-4.2.1 Risks and Mitigations
+Note, the crane is not needed for all routine maintenance of the cyclotron, as the top half is raised on jacks, providing access to the central plane and the injection, RF and extraction components.
+Figure 73: Cyclotron open and crane with cavern mesh - cut-view.
+6.2.1 Risks and Mitigation
 Risk: Coil Size.
-The cyclotron coil will likely not fit through all of the sections of the main access ramp at Yemilab in horizontal orientation.
-A special mounting fixture capable of holding the coil packages at an angle should be able to bring these coils through diagonally.
-However, there is some uncertainty about possible choke points in the main access ramp to the Yemilab facility.
-Mitigation: In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections, as discussed above.
+The cyclotron coils may not fit through all of the constrictions of the mining ramp, even with the most flexible conveyance device, because of irregularities in the ramp shape and size.
+Mitigation: In the case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections, as discussed above.
 Both come with drawbacks or additional cost, but are manageable.
-Chapter 5 Conclusion
+7 Conclusion
 7.1 Summary
@@
 Most notably, we are in the process of putting together the RFQ-Direct Injection Project (RFQ-DIP), combining the MIST-2 ion source (currently commissioning), the RFQ (currently being machined), and the HCHC-1.5 cyclotron (parts being ordered), as demonstrator of all novel concepts in the HCHC-60 cyclotron design.
-5.3 Acknowledgements
+[…]
+Acknowledgements
 This work was primarily supported by NSF grant PHY-2012897.
@@
 Winkler thankfully acknowledges the support from the MIT UROP office.
-[…]
+Declarations
+  • Data availability: Data is available from the authors upon reasonable request.
+Appendix A Appendix A - List of Acronyms
+  • ACCT: AC Current Transformer
+  • APS: American Physical Society
+  • AIMA: AIMA Développement
+  • BCSI: Best Cyclotron Systems, Inc.
+  • BEMPP: Boundary Element Method Python Package
+  • BSM: Beyond Standard Model
+  • CDR: Conceptual Design Report
+  • CERN: European Organization for Nuclear Research
+  • COMSOL: Computational Software for Physics Simulations
+  • CST: Computer Simulation Technology
+  • CW: Continuous Wave
+  • DI: De-Ionized (Water)
+  • EPICS: Experimental Physics and Industrial Control System
+  • ESC: Electrostatic Channel
+  • FEM: Finite Element Method
+  • FFA: Fixed-Field Alternating Gradient Accelerator
+  • HCHC: High-Current Hydrogen Cyclotron
+  • HV: High Voltage
+  • IBA: Ion Beam Applications
+  • IBSIMU: Ion Beam Simulation Software
+  • IBP: International Benchmarking Panel
+  • IBD: Inverse Beta Decay
+  • IEEE: Institute of Electrical and Electronics Engineers
+  • INFN: Istituto Nazionale di Fisica Nucleare
+  • ISO: International Organization for Standardization
+  • LEBT: Low Energy Beam Transport
+  • LDMOS: Laterally Diffused Metal-Oxide Semiconductor
+  • LLRF: Low-Level Radio Frequency
+  • LS: Liquid Scintillator
+  • LSC: Liquid Scintillator Counter (also referred to as vEYE)
+  • MSU: Michigan State University
+  • MEBT: Medium Energy Beam Transport
+  • MIT: Massachusetts Institute of Technology
+  • NME: Numerical Methods in Engineering
+  • OPAL: Object-Oriented Parallel Accelerator Library
+  • PSI: Paul Scherrer Institute
+  • PID: Proportional Integral Derivative (Control Loop)
+  • PDR: Preliminary Design Report
+  • QCD: Quantum Chromodynamics
+  • RF: Radio Frequency
+  • RFQ: Radio-Frequency Quadrupole
+  • SCCM: Standard Cubic Centimeter per Minute
+  • SI: Spiral Inflector
+  • SM: Standard Model
+  • SPMT: Self-Propelled Modular Transporter
+  • UHV: Ultra-High Vacuum
+  • VIS: Versatile Ion Source (built at INFN-LNS in Catania)